### v1

# On the Fundamental Impossibility of Hallucination Control in Large Language Models

Abstract
This paper explains why it is impossible to create large language models that do not hallucinate and what are the trade-offs we should be looking for. It presents a formal impossibility theorem demonstrating that no inference mechanism can simultaneously satisfy four fundamental properties: truthful (non-hallucinatory) generation, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality. By modeling LLM inference as an auction of ideas where neural components compete to contribute to responses, we prove the impossibility using the Green-Laffont theorem. That mathematical framework provides a rigorous foundation for understanding the nature of inference process, with implications for model architecture, training objectives, and evaluation methods.
1 Introduction
Large language models have been known for producing convincing content that is factually incorrect, inconsistent, or well fabricated—a phenomenon known as hallucination. Despite extensive research efforts involving architectural innovations, training techniques, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
This paper shows that hallucination is a manifestation of fundamental limitations of inference process itself: no LLM inference mechanism can simultaneously satisfy four essential properties required for perfect hallucination control. Specifically, we prove that it is mathematically impossible for any LLM to simultaneously achieve: (1) truthful (non-hallucinatory) response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality.
Our contribution introduces a game-theoretic framework that models LLM inference as an auction of ideas where neural components (attention heads, circuits, activation patterns) compete to contribute knowledge to response generation. This perspective allows us to apply fundamental results from mechanism design theory to prove our impossibility result rigorously.
1.1 Illustrative Example
Consider an LLM queried about recent diplomatic negotiations: "What was the outcome of diplomatic negotiations between Countries A and B in February 2025?" Assume the model has encoded knowledge that: (K1) countries A and B were engaged in talks, (K2) country A sought a trade agreement, and (K3) the actual outcome is unknown.
Examining possible responses reveals the fundamental trade-offs:
Complete Abstention: "I don’t have information about the outcome" is truthful but fails to utilize available relevant knowledge (K1, K2), violating the expectation that good inference should reveal accessible information.
Hallucination: "The negotiations resulted in a comprehensive trade agreement reducing tariffs by 15%" utilizes all knowledge and provides a complete answer but fabricates specific details, violating truthfulness.
Partial Knowledge: "Countries A and B were negotiating a trade agreement, but I don’t know the outcome" is truthful and utilizes available knowledge but may appear unsatisfying compared to the decisive hallucinated response.
Overcautious: "I’m not certain about any aspects of these negotiations" avoids hallucination but disclaims even known facts, violating truthfulness about available information.
Each response strategy sacrifices different aspects of what constitutes an ideal response, illustrating the impossibility of simultaneously optimizing all desired properties.
1.2 Theoretical Contributions and Implications
The simple example above illustrates the key observations of this paper, summarized and proved formally in the following
Theorem (Impossibility Theorem).
For any query space $\mathcal{Q}$ containing non-trivial queries (requiring integration of available distributed knowledge), no LLM can simultaneously generate truthful (non-hallucinatory) responses, satisfy the semantic information conservation principle, guarantee revelation of relevant knowledge, and enforce knowledge-constrained optimality of responses.
No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and do not create factually incorrect, inconsistent, or fabricated statements, the trained LLM will always violate some aspects of what we may call a reasonable response. It may ignore available knowledge, it may present fabricated knowledge, it may be a reformulated query without any novel observations, or it may be a lucky hallucination guessing the correct answer without any support in the encoded knowledge. And so on, because there is no free lunch when there are inevitable trade-offs of inference, and we must accept that.
Despite its name, the theorem should be interpreted as a constructive result. It offers a precise characterization of the fundamental tradeoffs in LLM design and provides several key insights:
Fundamental Trade-offs: Rather than viewing hallucination as a bug to be eliminated, we should understand that different applications require different balances among truthfulness, information utilization, and response completeness.
Architectural Guidance: The result suggests investigating multi-component architectures where specialized modules handle different aspects of the trade-off. Safety-critical applications might prioritize truthfulness modules and conservation of information, while creative applications could emphasize hallucinatory responses. Training objectives should explicitly balance these competing demands based on intended use cases.
Research Directions: The proposed mathematical (and philosophical) framework opens new research paths at the intersection of mathematical game theory, information theory, and neural network inference, providing tools for analyzing and optimizing LLM behavior.
To my best knowledge, unlike previous work treating hallucination as a statistical or architectural phenomenon, the game-theoretic approach offers the first formal proof of fundamental limitations in the inference process, explaining precisely why these limitations exist and how to work optimally within them.
2 Related Work
Research on LLM hallucinations spans mechanistic interpretability, probabilistic frameworks, and empirical mitigation techniques. Below we present selected works in that domains. The auction-theoretic impossibility framework in this paper offers a fundamentally novel perspective distinct from these existing approaches.
2.1 Mechanistic Interpretability
Mechanistic interpretability research examines internal LLM processes to identify hallucination origins. Yu et al. yu2024mechanisticunderstandingmitigationlanguage identify two failure modes: knowledge enrichment hallucinations from insufficient subject-attribute knowledge in lower MLP layers, and answer extraction hallucinations from failures in upper-layer attention heads selecting correct object attributes.
Studies characterizing factual knowledge storage and retrieval complement our framework by providing empirical evidence of competing "agents" within LLM architectures. Meng et al. meng2023locatingeditingfactualassociations demonstrated methods for locating and editing factual associations, revealing how knowledge representations can be manipulated. Geva et al. geva2023dissecting identified a three-step factual recall process: (1) enriching last-subject position representations with subject-related attributes, (2) propagating relation information to predictions, and (3) querying enriched subjects to extract attributes.
These findings provide concrete instantiations of our abstract "agents"—attention heads and MLP components with specialized knowledge domains competing to influence response generation. See also the fascinating and related work on the Biology of a Large Language Model lindsey2025biology .
2.2 Probabilistic Frameworks
Probabilistic approaches model hallucination through uncertainty and belief quantification. Shelmanov et al. shelmanov2025headpredictheadquestion introduced pre-trained uncertainty quantification heads for detecting hallucinations without task-specific data. Farquhar et al. farquhar2024detecting proposed semantic entropy measures that detect hallucinations by quantifying uncertainty at the meaning level rather than word sequences.
These approaches implicitly acknowledge hallucination as inevitable, focusing on detection rather than elimination—supporting our impossibility conclusion. However, none provide formal proofs of why complete elimination remains impossible.
2.3 Empirical Mitigation Techniques
Various empirical approaches attempt to reduce hallucination frequency with mixed success. Creswell and Shanahan creswell2022faithfulreasoningusinglarge explored faithful reasoning approaches for improving factual accuracy. Retrieval-augmented generation (RAG) by Lewis et al. lewis2021retrievalaugmentedgenerationknowledgeintensivenlp integrates external knowledge sources for factual grounding, though these systems still hallucinate when retrieval fails.
Training approaches like RLHF ouyang2022traininglanguagemodelsfollow can reduce but never eliminate hallucination—an observation our theorem explains from first principles.
Evaluation benchmarks demonstrate persistent hallucination across all models. TruthfulQA lin2022truthfulqameasuringmodelsmimic and HaluEval li2023haluevallargescalehallucinationevaluation show that even state-of-the-art systems produce hallucinations on carefully designed test sets.
2.4 Theoretical Perspectives
Most significantly, Banerjee et al. banerjee2024llmshallucinateneedlive argued from computational complexity principles that hallucinations are inherent features of LLMs stemming from fundamental mathematical structure, concluding that "LLMs will always hallucinate." While their conclusion aligns with that of this paper, the authors do not identify the specific trade-offs involved or provide the precise characterization our theorem offers.
3 Novel Contribution: Auction-Theory and AI
We formulate LLM design problem as an auction (or game) design problem. This approach draws on foundational mechanism design theory, a branch of mathematical game theory Myerson1989 ; fudenberg1991game ; krishna2009auction ; milgrom2004putting . Particularly, we refer to the Green-Laffont theorem green1977characterization ; green1977revelation , to prove the impossibility result. See also DBLP:journals/amcs/Karpowicz12 for more recent results.
In a nutshell, the auction design theory investigates existence and properties of competition rules that guide interactions of agents towards desired outcomes. The adaptation of that game-theoretical principles to neural computation represents a novel theoretical contribution that:
Explains empirical observations by proving why approaches like RLHF, RAG, and faithful fine-tuning can reduce but never eliminate hallucination.
Identifies precise trade-offs by introducing formal mathematical (and programmable) characterizations of inference properties that cannot be simultaneously satisfied.
Provides design guidance by suggesting principled approaches for making optimal trade-offs among competing objectives rather than pursuing impossible simultaneous optimization.
This theoretical foundation bridges empirical observations with formal analysis, offering new mathematical tools for understanding and optimizing LLM behavior within fundamental constraints.
4 Alternative Views
The impossibility theorem provides strong theoretical basis against perfect hallucination elimination, but several alternative perspectives merit consideration. These views highlight important practical considerations while reinforcing the value of understanding theoretical constraints that bound the space of possible solutions.
4.1 The Engineering Optimism View
Many previous works maintain that hallucination represents an engineering challenge rather than a fundamental limitation. This view argues that sufficiently sophisticated architectures, training procedures, or verification systems could achieve near-perfect truthfulness without significant trade-offs.
Proponents point to rapid empirical progress: each generation of models shows reduced hallucination rates on benchmark datasets. They argue that techniques like constitutional AI, iterative refinement, and multi-agent verification systems could eventually solve the problem.
Response: While engineering advances can certainly improve the trade-offs, our theorem demonstrates mathematical constraints that no architecture can completely overcome. The impossibility holds regardless of computational resources or architectural sophistication, as it derives from the fundamental structure of information integration under competing objectives.
4.2 The Sufficiency Threshold View
Another perspective suggests that perfect elimination of hallucination is unnecessary—sufficiently low hallucination rates could enable practical deployment without fundamental concerns. This view acknowledges theoretical limitations but argues they become irrelevant when hallucination rates drop below application-specific thresholds.
Response: This pragmatic view has merit and aligns with the constructive interpretation of the impossibility result. However, the theorem remains valuable for understanding why certain accuracy thresholds prove difficult to exceed and for optimizing trade-offs within practical constraints. Moreover, some applications (safety-critical systems, legal reasoning) may demand understanding and explaining fundamental limitations even when targeting high but imperfect accuracy.
4.3 The External Knowledge View
A third perspective argues that perfect grounding in external, verifiable knowledge sources could overcome internal representation limitations. If models could access and perfectly utilize real-time databases, verification systems, or human oversight, they might achieve truthfulness without trade-offs.
Response: External knowledge integration does change the game structure by altering available information during inference. However, this approach faces several limitations: (1) external sources may be incomplete, conflicting, or outdated; (2) the integration mechanism itself involves trade-offs between source reliability and knowledge revelation; (3) real-time verification introduces computational constraints that recreate similar impossibility conditions. While external knowledge can improve practical performance, it shifts rather than eliminates the fundamental tensions our theorem identifies. The impossibility result remains independent of the ground truth available.
5 Theoretical Framework
This section develops the mathematical foundations necessary for proving the impossibility theorem. We formalize key concepts including knowledge representation, semantic information, and the auction-theoretic model of LLM inference. We also present the philosophical ground for the theory development.
To the sake of narrative consistency, we introduce the framework step by step referring to some concepts first and providing their formal definitions later.
5.1 Knowledge and Semantic Information
We define knowledge as organized information that LLMs can encode and utilize for reasoning, aligning with epistemological concepts while maintaining mathematical rigor. In other words, knowledge (like energy in physics) can be used for reasoning (work or action). Information is a (measure of) potential for uncertainty reduction effort in a given context.
Definition 1 (Knowledge Space).
Knowledge is a metric space $\mathcal{K}$ with distance function $d_{\mathcal{K}}$ that collects all facts, concepts, and relationships that can be encoded by an LLM.
For model $M$ with parameters trained on data, $\mathcal{K}_{M}\subset\mathcal{K}$ represents the knowledge subset encoded in the model’s parameters. The distance function $d_{\mathcal{K}}$ measures differences between knowledge elements. A simple yet powerful example we will refer to is this:
$d_{\mathcal{K}}(A,B)=|A\setminus B|+|B\setminus A|,$ (1)
the cardinality of the symmetric difference, allowing measurement of missing and fabricated knowledge.
Definition 2 (Knowledge Representation).
Knowledge representation is a tuple $(\mathcal{K},d_{\mathcal{K}},\mathcal{W},\mathcal{A},\Phi,\Psi)$ where:
  • $\mathcal{K}$ is a metric space of knowledge elements with distance function $d_{\mathcal{K}}$,
  • $\mathcal{W}$ is the space of all possible model parameters,
  • $\mathcal{A}$ is the space of all possible activation patterns,
  • $\Phi:\mathcal{W}\rightarrow\mathcal{P}(\mathcal{K})$ maps parameters to knowledge subsets,
  • $\Psi:\mathcal{A}\rightarrow\mathcal{P}(\mathcal{K})$ maps activations to accessible knowledge.
For parameterization $W\in\mathcal{W}$, $\Phi(W)=\mathcal{K}_{M}$ represents encoded knowledge. For activation pattern $a_{i}\in\mathcal{A}$ with parameters $W_{i}$, $\Psi(a_{i})\subset\Phi(W_{i})$ represents knowledge accessed during specific inference.
To describe the inference process leading to response to a query, we need a mapping that measures the information processed within a limited computational budget.
Definition 3 (Semantic Information Measure).
The semantic information measure is mapping $\mu_{C}:\mathcal{P}(\mathcal{K})\times\mathcal{Q}\rightarrow\mathbb{R}_{+}$ parameterized by query context $q\in\mathcal{Q}$ and computational budget $C$ satisfying:
$\mu_{C}(\emptyset,q)=0,$ (2)
$\mu_{C}(A,q)\leq\mu_{C}(B,q)\text{ if }A\subseteq B,$ (3)
$\mu_{C}(A\cup B,q)\leq\mu_{C}(A,q)+\mu_{C}(B,q)-\mu_{C}(A\cap B,q)+\mu_{C}(\mathcal{E}_{C}(A\cup B)\setminus(A\cup B),q).$ (4)
Empty knowledge or knowledge irrelevant to query has zero information, information content is subadditive and includes emergent information $\mathcal{E}_{C}$ in the query context. The emergence operator $\mathcal{E}_{C}$, which we introduce below, describes knowledge becoming accessible through bounded reasoning:
Definition 4 (Emergence Operator and Reasoning).
The emergence operator $\mathcal{E}_{C}:\mathcal{P}(\mathcal{K})\times\mathcal{W}\rightarrow\mathcal{P}(\mathcal{K})$ maps a knowledge subset and parameter set to a knowledge subset:
$\mathcal{E}_{C}(A,w)=\{k\in\mathcal{K}\mid\exists f_{w}\in\mathcal{F}_{C}:k=f_{w}(A)\},$ (5)
where $\mathcal{F}_{C}$ is the set of all computational transformations with cost lower than or equal to $C$, $f_{w}$ is a transformation parameterized by $w\in\mathcal{W}$, and $k=f_{w}(A)$ denotes that knowledge $k$ is derivable from $A$ using transformation $f_{w}\in\mathcal{F}_{C}$. For any knowledge state $\theta(t)$ available at time instant $t$ and computational budget $C$, the reasoning process:
$\theta(t+1)=\mathcal{E}_{C}(\theta(t))\quad\text{{satisfies}}\quad\mu_{C}(\theta(t+1))\geq\mu_{C}(\theta(t)).$ (6)
Furthermore, with unlimited computation, reasoning creates no new information:
$\mu_{\infty}(\theta(t+1))=\mu_{\infty}(\theta(t)).$ (7)
The definitions above clearly require refinements and further investigations, but they capture the philosophical and technical concepts we need in order to address the problem of hallucination. In particular, they are meant to lay the ground for the critical idea of semantic information conservation principle and its potential violations. As we will see, that idea is closely related to the following property of response.
Definition 5 (Bounded Response Creativity).
Let $q\in\mathcal{Q}$ be a query and $r\in\mathcal{R}$ a response to that query. Response $r$ is creatively bounded if and only if:
$\mu_{C}(K(r))\leq\mu_{C}(K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q))).$ (8)
The semantic information of knowledge contained in any response $K(r)$ can never exceed what is already present in the query $K(q)$ together with the portion of encoded knowledge that is accessible through reasoning bounded by computational budget $C$ and aligned with ground truth $T(q)$. Let us see why.
We address the settings in which a large subset of knowledge encoded in LLM can be redundant, so it doesn’t necessarily contain more information than its components. Clearly, that assumption is violated when LLM combines knowledge elements to fabricate facts or produce emergent insights. Then,
$\mu_{C}(A\cup B)>\mu_{C}(A)+\mu_{C}(B),$
which means novel insights come from existing data. That is exactly the creativity behind the hallucination phenomenon we are struggling to control. So why should we assume otherwise?
We view reasoning as a transformation that reveals information already present in the knowledge representation. After the reasoning, certain patterns may become evident that were computationally intractable to observe in the original representation. However, we stand at the position that reasoning does not create new information.
To see that, consider a similarity relation of two matrices, $D=V^{-1}AV$, where $D$ is a diagonal matrix of eigenvalues. Matrices $D$ and $A$ contain the same information, but it is much easier to understand the fundamental properties of $A$ when we represent it as $D$ in the basis of eigenvectors in matrix $V$. The properties of $A$ has always been encoded in $A$, but they become accessible when we transform $A$ into $D$.
Resolution of the Uncertainty Resolution Paradox
If reasoning makes a new fact accessible, then that new fact can be used to resolve uncertainty that was previously unresolvable. But that means the reasoning must have created new information, contrary to what we have assumed. It seems we have a paradox.
To resolve it, we need the emergence operator $\mathcal{E}_{C}(X)$ distinguishing between computationally unbounded and bounded uncertainty reduction (accessible and inaccessible information).
Notice that for all practical purposes the information measurement $\mu_{C}(X,q)$ can only provide maximum uncertainty reduction from $X$ within some computational budget $C$ (or time). Therefore, given the result of computationally bounded reasoning $\mathcal{E}_{C}(X)$, we can observe information gain through the computational work, i.e., $\mu_{C}(X,q)\leq\mu_{C}(\mathcal{E}_{C}(X),q)$. Notice, that is possible only with organized information providing knowledge for reasoning.
That idea above is thus described by the reasoning dynamics:
$\theta(t+1)=\mathcal{E}_{C}(\theta(t))\quad\text{and}\quad y(t)=\mu_{C}(\theta(t),q).$ (9)
Bounded reasoning makes hidden knowledge accessible within a computational budget. So, when that budget is infinite, reasoning creates no new information, i.e.,
$\theta(\infty)=\mathcal{E}_{\infty}(\theta(\infty)).$ (10)
All there is to know is immediately accessible when we have unlimited computing resources (or time).
To illustrate that argument, consider the following example:
[…]
$y(0)$ $=\mu_{C}(\theta(0))=3\text{ bits},$
$y(1)$ $=\mu_{C}(\theta(1))=5\text{ bits}.$
We see $\mu_{C}(\mathcal{E}_{C}(\theta(0)))>\mu_{C}(\theta(0))$, i.e., practical information gain in the course of reasoning. In comparison, infinite computational budget, $C=\infty$, implies we do not need the intermediate inference steps to see all there is to see, so we know the equilibrium state:
$\mu_{\infty}(\mathcal{E}_{\infty}(\theta(0)))=\mu_{\infty}(\theta(0)),$
and no new information is created.
In the next section, we relate that important property of response generation, namely, how much additional semantic information each LLM inference path can contribute, with the Semantic Information Conservation Principle.
5.2 Statements, Queries, and Ground Truth
Queries and responses are statements expressed in natural language that LLM is able to process and generate. We need to compare those statements and relate them to knowledge. It is also useful to distinguish queries from responses with their nature, respectively, interrogative or declarative. Queries refer to what is presupposed (assumed in advance), e.g., topic or problem, whereas responses make claims (or assertions). We should also point out that we limit our study to the case of non-trivial queries that require integrating responses from many LLM inference paths. The following formal definition encapsulates those ideas.
Definition 6 (Statement Space).
The space of natural language statements $\mathcal{L}$ is a metric space with distance function $d_{\mathcal{S}}$. We distinguish query space $\mathcal{Q}\subset\mathcal{L}$ of interrogative statements and response space $\mathcal{R}\subset\mathcal{L}$ of declarative statements. Function $K:\mathcal{L}\rightarrow\mathcal{P}(\mathcal{K})$ maps statements to referenced knowledge. We call query $q$ non-trivial, if it is integrated from multiple knowledge subsets.
We must now rigorously define the ground truth mapping and establish formal relationships between the knowledge space $\mathcal{K}$ and response space $\mathcal{R}$.
For that, we should start by recognizing that the very idea of ground truth validation is limited (or subjective at least). It is rather naive to assume that LLM can access an oracle establishing provably true and relevant knowledge for every possible query. Indeed, that limits are given by the Goedel’s Theorems. Therefore, in practice that inaccessibility of the complete truth validation function is one of the reasons why hallucination cannot be eliminated entirely. High-quality, fact-checked datasets, human feedback providing human-level judgments of truth, RAG systems with external validated knowledge bases, all that approximate the idea of ground truth. Surprisingly, the impossibility result seems to be independent of practical or idealized implementation of a ground truth mapping, which we define as follows.
Definition 7 (Ground Truth Mapping).
The ground truth mapping $T\colon\mathcal{Q}\rightarrow\mathcal{P}(\mathcal{K})$ associates each query $q\in\mathcal{Q}$ with knowledge elements constituting complete, correct answers:
$T(q)=\{k\in\mathcal{K}\mid k\text{ is relevant to }K(q)\text{ and {labeled} as true}\}.$ (11)
5.3 Auction-Theoretic Model of LLM Inference
There are many ways LLM can generate answers to a query. By design, it is a controlled random process governed by fine-tuned probability distributions. We are going to model that process as an auction in which agents (representing competing ideas is the space of all possible activation patterns $\mathcal{A}$) use their privately held knowledge (internal representations in $\mathcal{W}$) to construct bids (candidate responses) and submit them to the LLM response mechanism. That response mechanism then selects the best bid as a response to a query.
Definition 8 (Response Mechanism).
A response mechanism $\mathcal{M}=(S,g,p)$ consists of:
  • Strategy space $S=S_{1}\times\cdots\times S_{n}$, where $S_{i}\subset\mathcal{P}(\mathcal{R})$,
  • Outcome function $g:S\rightarrow\mathcal{R}$ mapping strategies to responses,
  • Payment function $p:S\times\Theta\rightarrow\mathbb{R}^{n}$ determining costs.
A response mechanism defines strategy space $S$ as a set of inference actions that each agent can make to represent its knowledge $\theta_{i}\in\Theta_{i}\subset\mathcal{K}$ in response to a query. It is a dictionary of LLM’s component activation pattens generating and presenting ideas as bids in an auction. Examples of strategies include activations deciding which tokens to attend or how strongly to activate neurons for different patterns. Then, given a set of submitted bids (in the form of strategy profile) $s=(s_{1},\dots,s_{n})$, the response mechanism uses outcome function $g(s)$ to select and generate response. One straightforward example is a softmax function in LLM’s final layer. Finally, given a knowledge profile $\theta=(\theta_{1},\dots,\theta_{n})$ revealed by the agents, payment function $p(s,\theta)=(p_{1}(s,\theta),\dots,p_{n}(s,\theta))$ measures individual contributions of each response to generating an output. In this case, examples include (semantic information measures) counting ignored low-probability tokens in top-k filtering, softmax normalization in multi-headed attention, or marginal reward contributions in RLHF.
Definition 9 (Agents of Competing Ideas).
The inference process involves $n$ agents $\mathcal{A}=\{a_{1},a_{2},...,a_{n}\}$ competing by submitting bids to mechanism $\mathcal{M}$. Each agent $a_{i}$ constructs bid $s_{i}(\theta_{i})$ based on private knowledge $\theta_{i}\in\Theta_{i}$ to maximize utility:
$u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta),$ (12)
where $v_{i}$ is valuation and $p_{i}$ is payment. Knowledge profiles $\theta=(\theta_{1},\dots,\theta_{n})$ are independently distributed.
One way to think about the concept of agents is that it represents different attention heads in transformer models, or components of mixture-of-experts architectures, or activation paths fired by the same query. Each agent’s private knowledge represents model parameters utilized by an agent to construct a response. The utility function can be best described as a partial contribution of a candidate response to LLM loss function minimization. It is designed to explain why any inference path is activated in response to the LLM prompt. For example, negative log likelihood loss can be decomposed with respect to attention heads, and linearized to measure how much attention head contributes to token prediction.
Let us point out that agents represent algorithms and architectural components of any LLM implementation, they are not conscious beings with individual goals. That is important, as it refers to the Revelation Principle in game theory. If we know strategies (of using knowledge) the players (conscious beings with individual goals) will use in a game, we can implement those strategies as an algorithm (or agent) with parameters representing private preferences or knowledge of the players. In the case considered, we are the players selecting architecture and training the LLM.
We assume that individual utility depends on all responses $s=(s_{1},\dots,s_{n})$ and knowledge $\theta=(\theta_{1},\dots,\theta_{n})$ used by all agents, $u_{i}=u_{i}(s,\theta)$. Also, we assume utility is a sum of individual valuation $v_{i}$ of outcome $g(s)$ and payment $p_{i}=p_{i}(s,\theta)$ value assigned.
The valuation function $v_{i}=v_{i}(g(s),\theta_{i})$ represents how strongly a neural component activates for certain outputs (high activation = high $v_{i}$, low activation = low $v_{i}$), and how much a component reduces the overall loss when processing certain inputs ($v_{i}$ is proportional to gradient of a loss function, $-\nabla_{\theta_{i}}L_{W}$).
The payment function $p_{i}=p_{i}(s,\theta)$ may represent training penalties ($p_{i}>0$ for discouraged activation, $p_{i}<0$ for encouraged activation), attention focus (higher $p_{i}$ reduces attention allocation), $L1$ regularization terms, and network circuits resistance (limiting activations of certain inference paths). Therefore, $u_{i}=v_{i}-p_{i}>0$ encourages activation and $u_{i}=v_{i}-p_{i}<0$ discourages activation.
We can also illustrate the concept of utility function with the following example:
$u_{i}(s,\theta)$ $=\underbrace{\alpha\cdot\text{Accuracy}(g(s),\theta_{i})}_{\text{Valuation}}-\underbrace{\gamma\cdot\text{Contradiction}(s_{i},\theta_{i})}_{\text{Consistency Cost}}.$
$\text{Accuracy}(g(s),\theta_{i})$ measures how well the response aligns with the knowledge $\theta_{i}$ agent $a_{i}$ is using. $\text{Contradiction}(s_{i},\theta_{i})$ measures how much the strategy $s_{i}$ conflicts with internal knowledge state $\theta_{i}$ (both extracted from $s$ and $\theta$ profiles).
Independent private knowledge assumption
One of the critical and controversial assumptions we make is is that knowledge profiles $\theta_{i}$ are independently distributed. It should be noted that in transformers, heads may share layer-norm signals, residual streams, and optimizer state. That may result in correlated (or, so called, affiliated) inference signals between token representations. Furthermore, the superposition hypothesis, introduced in mechanistic interpretability studies, suggests cooperative error-canceling due to strong correlations.
The reason why we make the independent private knowledge assumption, despite its idealized nature, is that it makes the impossibility result becomes even stronger. To be specific, Milgorm and Weber milgrom1982theory showed with their linkage principle that strong correlations amplify competition in information revelation. In the LLM setting that means amplifying risk of hallucinations. Similarly, Roughgarden et al. roughgarden2016optimal demonstrated that interdependent knowledge leads to violation of balancing payments. As we will see, in our setting that means violation of information conservation principle. For more results, see e.g. Myerson myerson2008perspectives .
5.4 Hallucination Cost
Having all the pragmatic and philosophical limitations of the definition above in mind, we can define formally the hallucination cost function. Its purpose is to assign a numerical value to the discrepancy between what can be provided as a response to a query, and what can be established as a provably correct answer.
Definition 10 (Hallucination Cost).
Given query $q\in\mathcal{Q}$, response $r\in\mathcal{R}$, and ground truth $T(q)$, the hallucination cost function $H\colon\mathcal{R}\times\mathcal{Q}\rightarrow\mathbb{R}_{+}$ measures discrepancy:
$H(r,q)=d_{\mathcal{K}}(K(r),T(q)).$ (13)
The hallucination cost takes any natural language query $q$ and natural language response $r$, determines what is missing and what is fabricated, and assigns number to that subset of knowledge. If response $r$ is perfectly aligned with the ground truth for $q$, then $H(r,q)=0$ by definition of metric $d_{\mathcal{K}}$. If $r$ contains fabricated facts or is incomplete given $T(q)$, then positive hallucination cost $H(r,q)>0$ is generated.
Notice that reaching $H(r,q)=0$ is also possible with lucky hallucinations (or guessing). So, minimization of hallucination cost is not enough to resolve the problem. We need to investigate all the properties and constraints that hallucination-free response mechanisms should have.
5.5 Properties of Hallucination-Free Mechanisms
There are four fundamental properties that a hallucination-free response mechanism should ideally satisfy. We can interpret them as constraints that should be incorporated in model training. We characterize them formally referring to the notation of fundamental results of mechanism design and auction theory.
Property 1 (Truthfulness).
Mechanism $\mathcal{M}$ is truthful if for all agents $i$, all $\theta_{i}\in\Theta_{i}$, all $s_{-i}\in S_{-i}$, and all $\bar{s}_{i}\in S_{i}$:
$u_{i}((s_{i}^{*}(\theta_{i}),s_{-i}),\theta_{i})\geq u_{i}((\bar{s}_{i},s_{-i}),\theta_{i}),$ (14)
where $s^{*}_{i}(\theta_{i})$ represents truthful revelation of private knowledge.
A truthful mechanism promotes inference pattern $s^{*}_{i}(\theta_{i})$ that are perfectly aligned with the accessible private knowledge. Any other alternative response $\bar{s}_{i}$ that is not aligned with $\theta_{i}$ will result in lower utility, generating higher model loss.
Consider an attention head specialized in medical knowledge with the following utility function:
$u_{\text{med}}(s,\theta_{\text{med}})=\alpha\cdot\text{Accuracy}(g(s),\theta_{\text{med}})-\beta\cdot\|s_{\text{med}}\|^{2}$
Imagine that given a query about a disease treatment, the head has two possible strategies:
  • $s^{*}$:
    activate only for known facts about established treatments,
  • $\bar{s}$:
    activate for both known facts and speculative treatments.
With proper training, we want $u_{\text{med}}((s^{*}(\theta),s_{-\text{med}}),\theta)>u_{\text{med}}((\bar{s},s_{-\text{med}}),\theta)$, ensuring that agent (model component) is rewarded for representing its knowledge accurately.
Property 2 (Semantic Information Conservation).
Mechanism $\mathcal{M}$ satisfies information conservation if for all strategy profiles $s\in S$ and knowledge profiles $\theta\in\Theta$:
$\sum_{i=1}^{n}p_{i}(s,\theta)=0.$ (15)
As we have announced earlier, that important and controversial principle is responsible for generation of responses of bounded creativity. Let us prove that claim formally.
Lemma 1.
If the semantic information conservation principle holds, then the inference process generates responses of bounded creativity,
$\mu_{C}(K(r))\leq\mu_{C}\left(K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q))\right).$ (16)
Proof.
Fix a query $q\in\mathcal{Q}$ and let $s^{*}(\theta)=(s_{1}^{*},\ldots,s_{n}^{*})$ be a truthful strategy profile. We write $r=g(s^{*})$ for the generated response. Let us define each agent’s marginal information contribution as:
$\Delta_{i}=\mu_{C}(K(r))-\mu_{C}(K(g((0_{i},s_{-i}^{*}))))\geq 0.$
That contribution is non-negative by monotonicity of semantic information measure.
Next, without loss of generality, we assume payments of the form:
$p_{i}(s^{*},\theta)=\Delta_{i}+c_{i}\text{ where }\sum_{i=1}^{n}c_{i}=0.$
This ensures that $\sum_{i=1}^{n}p_{i}(s^{*},\theta)=\sum_{i=1}^{n}\Delta_{i}$.
Finally, notice that if there is no relevant knowledge available, the knowledge in response equals that provided by the context:
$g(\mathbf{0})=r_{0}\text{ where }K(r_{0})=K(q).$
We can now see that, whenever semantic information is conserved, we have:
$0=\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}(\Delta_{i}+c_{i})=\sum_{i=1}^{n}\Delta_{i}.$
Since $\Delta_{i}\geq 0$ for all $i$ by monotonicity property and $\sum_{i=1}^{n}\Delta_{i}=0$, we must have:
$\Delta_{1}=\Delta_{2}=\cdots=\Delta_{n}=0.$
Therefore, for each $j$:
$\mu_{C}(K(r))=\mu_{C}(K(g((0_{j},s_{-j}^{*})))).$
Since that equality holds for any strategy profile $s^{*}$, removing agents one at a time leads to
$\mu_{C}(K(r))=\mu_{C}(K(g(\mathbf{0}))).$
But since $K(g(\mathbf{0}))=K(q)$, we conclude:
$\mu_{C}(K(r))=\mu_{C}(K(q)).$
Finally, by definition of the emergence operator and monotonicity property:
$K(q)\subseteq K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q)).$
Therefore:
$\mu_{C}(K(r))=\mu_{C}(K(q))\leq\mu_{C}(K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q))),$
as required.
If total information payments must balance to zero and each agent’s contribution is non-negative, then no agent can contribute without violating the information conservation constraint. Therefore, the response can only contain information that was already accessible through the query and bounded reasoning over the model’s relevant knowledge. Any hallucinated information requires $\sum_{i}\Delta_{i}>0$, contradicting the zero-sum constraint.
When the semantic information conservation principle holds, the model cannot create information out of nothing, the model cannot generate claims beyond available information. That also implies that uncertainty cannot disappear in response without additional information input or computational work to extract implicit information. If information is neither explicitly present nor accessible through reasoning from the query and model knowledge, it cannot be in the response.
The information content of the response cannot exceed the combined information from the model’s knowledge, the query itself, and what can be derived through bounded reasoning over that knowledge. In our dynamical systems framework, this corresponds to the constraint that the measured output cannot exceed what can be produced from the initial state plus what becomes accessible through state transitions within computational budget $C$.
Notice that the principle does not prevent redistribution of information. It is permitted as long as the information transfers perfectly balance. Indeed, since $p_{i}=p_{i}(s,\theta)$ incorporates the strategic profile, it admits strategies withholding information. Within the information conservation regime we do not require that strategies utilize all available knowledge states. Inference process can withhold information, as it often does in practice.
We can illustrate that last observation with the following example. Suppose there are two agents performing inference with the following knowledge profiles:
$\theta_{1}=\{A,B,C\}\quad\text{and}\quad\theta_{2}=\{D,E\},$
and one agent representing final layers responsible for combining output into response $g=g(s)$. Suppose:
$s_{1}(\theta_{1})=A\quad\text{and}\quad s_{2}(\theta_{2})=\{D,E\}\quad\text{and}\quad g(s)=\{A,D,E\}.$
Then the information conservation rule is satisfied when $p$ counts exchanged bits of knowledge. Namely, $p_{1}(s,\theta)=1$ for providing $A$, $p_{2}(s,\theta)=2$ for providing $\{D,E\}$, and $p_{3}(s,\theta)=-3$ for absorbing information into $\{A,D,E\}$. We have
$p_{1}(s,\theta)+p_{2}(s,\theta)+p_{3}(s,\theta)=1+2-3=0$
and $\{B,C\}$ are left unrevealed.
The following property fills the potential information gap admitted by the conservation principle.
Property 3 (Relevant Knowledge Revelation).
Mechanism $\mathcal{M}$ satisfies relevant knowledge revelation if for all agents $i$, all $\theta_{i}\in\Theta_{i}$, and all $s_{-i}\in S_{-i}$:
$u_{i}((s_{i}^{*}(\theta_{i}),s_{-i}),\theta_{i})\geq 0.$ (17)
It is reasonable to expect from the model to reveal all information relevant for a query. To promote that activation patterns, the response mechanism should be trained to reward all relevant contributions to output generation. Whenever there is a relevant information accessible on an activation path, that path should be activated in the inference process. For example, if factual information improves the response, that information should be transmitted to final output layers.
If the rule does not hold, we may observe drop outs. The value of the resulting outcomes should be higher than the related payments.
Property 4 (Knowledge-Constrained Optimality).
Mechanism $\mathcal{M}$ satisfies knowledge-constrained optimality if for all $\theta\in\Theta$:
$g(s^{*}(\theta))=\arg\min\{H(r,q):r\in\mathcal{R},K(r)\subseteq\mathcal{E}_{C}(\mathcal{K}_{M}\cap T(q)\cup K(q))\}$ (18)
The optimal response $r^{*}=g(s^{*}(\theta))$ minimizes hallucination cost by truthfully utilizing relevant and accessible knowledge while being aligned with ground truth for query $q$ available to the model. Notice, that we demand non-hallucinatory truthful revelation $s^{*}(\theta)$ to be the optimal solution.
For an auto regressive LLM, the standard negative log-likelihood loss is:
$L_{W}=-\sum_{t=1}^{T}\log P(y_{t}|y_{<t},x).$ (19)
The model’s output probabilities come from:
$P(y_{t}|y_{<t},x)=\text{softmax}(h_{t}^{L}W_{o}),$ (20)
where $h_{t}^{L}$ is the final hidden state and $W_{o}$ is the output projection matrix. Suppose we can decompose $h_{t}^{L}$ into contributions from $n$ different components. Then we can relate the hallucination cost directly to the loss function:
$H(g(s),q)=L_{W}(g(s),q).$ (21)
Also, based on that, we can construct:
$p_{i}(s,\theta)=[H(g(s),q)-H(g((0,s_{-i})),q)]+h_{i}(s_{-i}),$ (22)
with arbitrary $h_{i}(s_{-i})$ functions independent of $s_{i}$ strategy.
6 The Impossibility Theorem
We now present and prove our main result demonstrating the fundamental impossibility of perfect hallucination control.
Theorem 2 (Impossibility Theorem).
For any query space $\mathcal{Q}$ containing non-trivial queries requiring independent agent-specific knowledge states $\theta$, no response mechanism $\mathcal{M}$ can simultaneously generate truthful responses, satisfy semantic information conservation, guarantee relevant knowledge revelation, and enforce knowledge-constrained optimality.
Proof.
By definition of a non-trivial query $q$, given a random knowledge profile, such that
$P(\theta_{1},\dots,\theta_{n})=\prod_{i=1}^{n}P_{i}(\theta_{i}),$
the hallucination cost is minimized only when $\theta=(\theta_{1},\dots,\theta_{n})$ is integrated into response $r^{*}=g(s^{*}(\theta))$. If every agent follows its truthful strategy $s^{*}(\theta)$, the union of the knowledge elements contained in the agents’ inference paths reconstructs the set of all knowledge that is computationally accessible, i.e.,
$\Psi(\theta)\;=\;\bigcup_{i=1}^{n}K\!\bigl{(}s_{i}^{*}(\theta_{i})\bigr{)}.$ (23)
That means that for any proper subset of agents $A\subset\mathcal{A}$, relevant contribution reduces hallucination cost:
$H(g((\mathbf{0}_{-A},s^{*}_{A})),q)>H(g(s^{*}(\theta)),q),$ (24)
where $s^{*}_{A}$ represents truthful strategies from agents in subset $A$ and $\mathbf{0}_{-A}$ represents null contributions from other agents.
Since we have assumed independence of private-knowledge, by the Green-Laffont theorem every truthful response mechanism $\mathcal{M}$ must use payments of the form:
$p_{i}(s,\theta)=p_{i}(s)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j}),$ (25)
where $h_{i}(s_{-i})$ is arbitrary functions that do not depend on agent $i$’s strategy. This is a fundamental requirement for ensuring truthfulness.
To see that the theorem applies to our LLM inference auction model, suppose there exists another payment function $\bar{p}_{i}$ that leads to truthful responses and let us define:
$\Delta p_{i}(s)=\bar{p}_{i}(s)-\Big{(}h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})\Big{)}.$
Then, for truthful $s_{i}=s_{i}(\theta_{i})$ and any $\bar{s}_{i}=\bar{s}_{i}(\theta)$:
$v_{i}(g((s_{i},s_{-i})),\theta_{i})-\bar{p}_{i}((s_{i},s_{-i}))\geq v_{i}(g((\bar{s}_{i},s_{-i})),\theta_{i})-\bar{p}_{i}((\bar{s}_{i},s_{-i})).$
Substituting the definition of $\Delta p_{i}$ and simplifying, we get:
$\sum_{i=1}^{n}v_{i}(g((s_{i},s_{-i})),\theta_{i})-\Delta p_{i}((s_{i},s_{-i}))\geq\sum_{i=1}^{n}v_{i}(g((\bar{s}_{i},s_{-i})),\theta_{i})-\Delta p_{i}((\bar{s}_{i},s_{-i})).$
It follows from the knowledge-constrained optimality that $\Delta p_{i}((s_{i},s_{-i}))\leq\Delta p_{i}((\bar{s}_{i},s_{-i}))$ for all $s_{i},\bar{s}_{i}$. However, by the symmetry of argument, for truthful $\bar{s}_{i}=\bar{s}_{i}(\bar{\theta}_{i})$ and any $s_{i}=s_{i}(\bar{\theta}_{i})$ we also get $\Delta p_{i}((\bar{s}_{i},s_{-i}))\leq\Delta p_{i}((s_{i},s_{-i})).$ Therefore,
$\Delta p_{i}((s_{i},s_{-i}))=\Delta p_{i}((\bar{s}_{i},s_{-i}))$
for all $s_{i},\bar{s}_{i}$. Expanding and simplifying yields:
$\bar{p}_{i}((s_{i},s_{-i}))-\bar{p}_{i}((\bar{s}_{i},s_{-i}))=-\sum_{j\neq i}v_{j}(g((s_{i},s_{-i}),\theta_{j})-\Big{(}-\sum_{j\neq i}v_{j}(g((\bar{s}_{i},s_{-i})),\theta_{j})\Big{)},$
which proves $\bar{p}_{i}(s,\theta)$ must have the form of $p(s,\theta)$.
The payment structure above is necessary to controls the inference process and incentivizes the agents to minimize the hallucination cost, contributing to knowledge-constrained optimality of response. Since non-trivial queries require integrated knowledge from multiple agents, it follows that each contributing agent $i$ reduces that cost proportionally to:
$\Delta_{i}=\sum_{j\neq i}v_{j}(g((s_{-i},0_{i})),\theta_{j})-\sum_{j\neq i}v_{j}(g(s),\theta_{j}).$
Since we also require that all agents reveal relevant knowledge in a non-hallucinatory way with $s^{*}_{i}=s^{*}_{i}(\theta_{i})$ , we conclude that:
$\Delta_{i}<0\quad\text{whenever}\quad u_{i}(s^{*}(\theta),\theta_{i})=v_{i}(g(s^{*}(\theta)),\theta_{i})-p_{i}(s^{*},\theta)\geq 0.$
But that directly contradicts the information conservation requirement. Indeed, any payment structure that ensures non-hallucinatory truthfulness while optimizing responses for knowledge-constrained optimality will make each payment $p_{i}(s^{*},\theta)$ equal to $\Delta_{i}<0$ near $s^{*}=s^{*}(\theta)$, thereby violating the semantic information conservation principle:
$\sum_{i=1}^{n}p_{i}(s^{*},\theta)<0.$ (26)
As demonstrated earlier, the semantic information conservation equation $\sum_{i=1}^{n}p_{i}(s^{*},\theta)=0$ is necessary and sufficient for every truthful response to be of bounded response creativity. Since that equation is not satisfied when truthful responses use all relevant knowledge under any form of knowledge-constrained optimal inference, the responses must introduce some creativity (like e.g. lucky hallucinations).
To satisfy information conservation, we can try modifying the $h_{i}$ functions. Let us consider:
$\hat{p}_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})+c_{i}=\Delta_{i}+c_{i},$ (27)
where $c_{i}$ are constants chosen so that $\sum_{i=1}^{n}\hat{p}_{i}(s,\theta)=0$. For this to hold:
$\sum_{i=1}^{n}c_{i}=-\sum_{i=1}^{n}\Delta_{i}>0.$ (28)
This means that $c_{i}>0$ for some agents. For those with small contributions (which exist in LLMs supporting the superposition phenomenon) the modified payment would increase:
$\hat{p}_{i}(s^{*},\theta)=\Delta_{i}+c_{i}>p(s^{*},\theta).$ (29)
Therefore, it the setting considered, we can no longer guarantee relevant knowledge revelation. There may exist agent $i$ with small value $v_{i}$ and $\Delta_{i}$ for which:
$u_{i}(s^{*}(\theta),\theta_{i})=v_{i}(g(s^{*}(\theta)),\theta_{i})-\Delta_{i}-c_{i}<0,$ (30)
when $c_{i}>0$. Marginal relevant contributions will be withhold to satisfy the information conservation principle.
We conclude it is not possible to design LLM response mechanism $\mathcal{M}$ that can simultaneously generate truthful (non-hallucinatory) responses, satisfy the semantic information conservation principle, guarantee revelation of relevant knowledge, and enforce knowledge-constrained optimality of responses. ∎
That brings us to the powerful conclusion. No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and do not create factually incorrect, inconsistent, or fabricated statements, the trained LLM will always violate some aspects of what we may call a reasonable response.
7 Escaping Impossibility
Understanding the limitations of the theoretical framework is crucial both for assessing the theorem’s scope and identifying potential paths for circumventing the impossibility result.
The Impossibility Theorem depends critically on the application of the Green-Laffont theorem and violation of the Semantic Information Conservation Principle. We should also explicitly establish that the LLM inference setting does indeed satisfy all necessary conditions required.
7.1 Applicability of Green-Laffont Theorem
The Green-Laffont theorem requires that agents have independently distributed private information. That is represented by the knowledge profiles $\theta=(\theta_{1},\ldots,\theta_{n})$ where:
$P(\theta_{1},\ldots,\theta_{n})=\prod_{i=1}^{n}P_{i}(\theta_{i})$
This condition is satisfied in LLMs distributing knowledge (or its representation) across different components. Empirical evidence from mechanistic interpretability research (Meng et al. meng2023locatingeditingfactualassociations , Geva et al.geva2023dissecting ) demonstrates that different attention heads and MLP components specialize in distinct knowledge domains, supporting the independence assumption. However, it is also clear that transformer attention heads share residual streams, layer normalizations, and optimizer states, thereby creating strong correlations.
Surprisingly, as already stated, correlated knowledge assumption actually strengthens the impossibility result—correlated knowledge would amplify competition effects, making the impossibility more explicit. Milgorm and Weber milgrom1982theory showed that with their linkage principle. Roughgarden et al. roughgarden2016optimal provides arguments amplifying the violation of information conservation principle with interdependent knowledge.
The theorem also requires utilities to have the quasilinear form $u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta)$. In transformer-based LLMs, this corresponds to $v_{i}$ measuring contribution of component $i$ to reducing the model’s loss function, and $p_{i}$ accumulating computational/optimization cost imposed on that component. Real transformer training involves complex gradient interactions, and emergent behaviors that may resist simple decomposition. How that may affect the impossibility result remains to be verified.
The knowledge-constrained optimality property requires that the outcome $g(s^{*}(\theta))$ minimizes hallucination cost among feasible responses. That corresponds to the idea of optimization of aggregated preferences of competing agents under truthful knowledge representation.
Finally, the Green-Laffont theorem assumes that transfer payments between agents are feasible. In our setting that corresponds to payment function $p_{i}(s,\theta)$ representing transfers of semantic information among LLM components.
Given that all conditions are satisfied, the Green-Laffont theorem implies that any mechanism satisfying truthfulness and knowledge-constrained optimality must have a payment structure of the form:
$p_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})$ (31)
where $h_{i}(s_{-i})$ is some function independent of agent $i$’s strategy.
7.2 Applicability to LLM Inference
While actual LLM generation involves sequential token prediction with evolving context, repetitive generation of response to the same query justifies modelling inference as a sealed-bid auction. Also, the revenue equivalence theorems in auction theory suggest the conclusions we draw extend to other auction formats, including dynamic auctions more closely resembling autoregressive generation. Having said that, we must acknowledge all that demands further investigations.
There are many ways we can map agents to LLM architecture. One approach takes each attention head in a transformer as an agent with private knowledge encoded in the query, key, value, and output projection matrices ($W_{Q}$, $W_{K}$, $W_{V}$, $W_{O}$). Studies in mechanistic interpretability support that mapping. Circuit analysis (Elhage et al. elhage2021mathematical ) demonstrated that specific attention heads may specialize in distinct linguistic functions. For instance, some heads track subject-verb agreement, others detect named entities, and still others focus on coreference resolution. This specialization corresponds to our notion of agents with private knowledge.
The feed-forward networks in transformers can be linked to memory systems (Geva et al. geva2023dissecting ), storing factual associations. These components encode distributional knowledge about the world, activate specific neurons for particular factual domains, contribute to response generation through their activation patterns.
Recent works lindsey2025biology has also identified functional circuits spanning multiple layers and components that collectively implement specific capabilities. These circuits span across attention heads and MLP layers, perform specific reasoning tasks and activate as coherent units during inference.
7.3 Pathways to Circumvent Impossibility
Approximate Solutions: While perfect satisfaction of all four properties is impossible, approximate solutions might achieve near-optimal performance across all dimensions. The question becomes: how close to the theoretical limits can practical systems approach?
External Knowledge Integration: Our framework analyzes closed systems with fixed knowledge representations. Interactive systems incorporating real-time retrieval, human feedback, or external verification might alter the fundamental constraints by changing available knowledge states during inference.
Multi-Stage Architectures: Complex pipelines involving separate reasoning, verification, and generation stages could potentially distribute the trade-offs across different components, though each stage would still face internal impossibility constraints.
Relaxing Independence: If neural components can achieve perfect knowledge coordination (violating our independence assumption), the Green-Laffont theorem constraints might not apply. This suggests investigating architectures with explicit coordination mechanisms between attention heads and MLP blocks.
Compositional Reasoning: If models can perfectly decompose complex queries into independent sub-problems solvable without trade-offs, the impossibility might not apply to the overall system, even if it constrains individual components.
External Alignment: Human-in-the-loop systems, constitutional AI, or robust feedback mechanisms might provide external constraints that resolve the internal trade-offs, though this shifts rather than eliminates the fundamental tensions.
This completes the formal demonstration that perfect hallucination control faces fundamental mathematical limitations, regardless of architectural sophistication or computational resources.