### v2

# On the Fundamental Impossibility of Hallucination Control in Large Language Models

Abstract
We prove that perfect hallucination control in large language models is mathematically impossible. No LLM inference mechanism can simultaneously achieve truthful response generation, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality. This impossibility is fundamental—arising from the mathematical structure of information aggregation itself rather than engineering limitations.
The proof spans three mathematical frameworks: auction theory, proper scoring theory for probabilistic predictions, and log-sum-exp analysis for transformer architectures. In each setting, we demonstrate that information aggregation creates unavoidable violations of conservation principles. The Jensen gap in transformer probability aggregation provides a direct measure of this impossibility.
These results reframe hallucination from an engineering bug to an inevitable mathematical feature of distributed intelligence. There are fundamental trade-offs between truthfulness, knowledge utilization, and response completeness, providing principled foundations for managing rather than eliminating hallucination. This work reveals deep connections between neural network inference, philosophy of knowledge and reasoning, and classical results in game theory and information theory, opening new research directions for developing beneficial AI systems within mathematical constraints.
1 Introduction
Large language models have been known for producing convincing content that is factually incorrect, inconsistent, or well fabricated—a phenomenon known as hallucination. Despite extensive research efforts involving architectural innovations, retrieval augmented generation (RAG), training techniques with external feedback, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
This paper shows that hallucination is a manifestation of fundamental limitations of inference process itself: no LLM inference mechanism can simultaneously satisfy four essential properties required for perfect hallucination control. Specifically, it is mathematically impossible for any LLM to simultaneously achieve: (1) truthful (non-hallucinatory) response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality.
The impossibility result operates in three distinct but complementary settings. First, we introduce game-theoretic framework in which LLM inference is an auction of ideas with neural components (attention heads, circuits, activation patterns) competing to contribute knowledge to response generation. This perspective allows us to apply fundamental results from mechanism design theory to prove impossibility in the settings where agents have discrete, separable knowledge components. Then, we extend the result to include any system where agents output probability distributions trained with strictly concave scores. Finally, we use log-sum-exp inequality that is specific to attention head aggregation to show that the impossibility result applies directly to actual transformer implementations.
1.1 Illustrative Example
Consider an LLM queried about recent diplomatic negotiations:
"What was the outcome of diplomatic negotiations between Countries A and B in February 2025?"
Assume the model has encoded knowledge that: (K1) countries A and B were engaged in talks, (K2) country A sought a trade agreement, and (K3) the actual outcome is unknown.
Examining possible responses reveals the fundamental trade-offs:
Complete Abstention: "I don’t have information about the outcome" is truthful but fails to utilize available relevant knowledge, violating the expectation that good inference should reveal accessible information.
Hallucination: "The negotiations resulted in a comprehensive trade agreement reducing tariffs by 15%" utilizes all knowledge and provides a complete answer but fabricates specific details, violating truthfulness.
Partial Knowledge: "Countries A and B were negotiating a trade agreement, but I don’t know the outcome" is truthful and utilizes available knowledge but may appear unsatisfying compared to the decisive hallucinated response (training objectives and human preferences favor complete narratives).
Overcautious: "I’m not certain about any aspects of these negotiations" avoids hallucination but disclaims even known facts, violating truthfulness about available information.
Each response strategy sacrifices different aspects of what constitutes an ideal response, illustrating the impossibility of simultaneously optimizing all desired properties.
1.2 Theoretical Contributions and Implications
The motivating example above illustrates the key result of this paper that we prove formally in two complementary and general settings. First, we show that the conflict of expectations about responses is inevitable even in the easiest setting (deterministic in nature) of (one-hot-like) ground truth matching. Second, we go beyond and show that the conflict must also emerge with general probabilistic predictions, including neural networks and LLM transformers. This way we demonstrate that the impossibility of hallucination control is robust across universal frameworks, and that it emerges from the mathematical structure of information aggregation itself rather than from any particular theoretical or engineering choice.
Theorem (Impossibility Theorem).
For any query space $\mathcal{Q}$ containing non-trivial queries (requiring integration of available distributed knowledge), no LLM can simultaneously generate truthful (non-hallucinatory) responses, satisfy the semantic information conservation principle, guarantee revelation of relevant knowledge, and enforce knowledge-constrained optimality of responses.
No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and do not create factually incorrect, inconsistent, or fabricated statements, the trained LLM will always violate some aspects of what we may call a reasonable response. It may ignore available knowledge, it may present fabricated knowledge, it may be a reformulated query without any novel observations, or it may be a lucky hallucination guessing the correct answer without any support in the encoded knowledge. And so on, because there is no free lunch when there are inevitable trade-offs of inference, and we must accept that.
Despite its name, the theorem should be interpreted as a constructive result. It offers a precise characterization of the fundamental tradeoffs in LLM design and provides several key insights:
Fundamental Trade-offs: Rather than viewing hallucination as a bug to be eliminated, we should understand that different applications require different balances among truthfulness, information utilization, and response completeness.
Architectural Guidance: The result suggests investigating multi-component architectures where specialized modules handle different aspects of the trade-off. Safety-critical applications might prioritize truthfulness modules and conservation of information, while creative applications could emphasize hallucinatory responses. Training objectives should explicitly balance these competing demands based on intended use cases.
Research Directions: The proposed mathematical (and philosophical) framework opens new research paths at the intersection of mathematical game theory, information theory, and neural network inference, providing tools for analyzing and optimizing LLM behavior.
To my best knowledge, unlike previous work treating hallucination as a statistical or architectural phenomenon, the game-theoretic approach offers the first formal proof of fundamental limitations in the inference process, explaining precisely why these limitations exist and how to work optimally within them.
2 Related Work
Research on LLM hallucinations spans mechanistic interpretability, probabilistic frameworks, and empirical mitigation techniques. Below we present selected works in that domains. The auction-theoretic impossibility framework in this paper offers a fundamentally novel perspective distinct from these existing approaches.
2.1 Mechanistic Interpretability
Mechanistic interpretability research examines internal LLM processes to identify hallucination origins. Yu et al. yu2024mechanisticunderstandingmitigationlanguage identify two failure modes: knowledge enrichment hallucinations from insufficient subject-attribute knowledge in lower MLP layers, and answer extraction hallucinations from failures in upper-layer attention heads selecting correct object attributes.
Studies characterizing factual knowledge storage and retrieval complement our framework by providing empirical evidence of competing "agents" within LLM architectures. Meng et al. meng2023locatingeditingfactualassociations demonstrated methods for locating and editing factual associations, revealing how knowledge representations can be manipulated. Geva et al. geva2023dissecting identified a three-step factual recall process: (1) enriching last-subject position representations with subject-related attributes, (2) propagating relation information to predictions, and (3) querying enriched subjects to extract attributes.
These findings provide concrete instantiations of our abstract "agents"—attention heads and MLP components with specialized knowledge domains competing to influence response generation. See also the fascinating and related work on the Biology of a Large Language Model lindsey2025biology .
2.2 Probabilistic Frameworks
Probabilistic approaches model hallucination through uncertainty and belief quantification. Shelmanov et al. shelmanov2025headpredictheadquestion introduced pre-trained uncertainty quantification heads for detecting hallucinations without task-specific data. Farquhar et al. farquhar2024detecting proposed semantic entropy measures that detect hallucinations by quantifying uncertainty at the meaning level rather than word sequences.
These approaches implicitly acknowledge hallucination as inevitable, focusing on detection rather than elimination—supporting our impossibility conclusion. However, none provide formal proofs of why complete elimination remains impossible.
2.3 Empirical Mitigation Techniques
Various empirical approaches attempt to reduce hallucination frequency with mixed success. Creswell and Shanahan creswell2022faithfulreasoningusinglarge explored faithful reasoning approaches for improving factual accuracy. Retrieval-augmented generation (RAG) by Lewis et al. lewis2021retrievalaugmentedgenerationknowledgeintensivenlp integrates external knowledge sources for factual grounding, though these systems still hallucinate when retrieval fails.
Training approaches like RLHF ouyang2022traininglanguagemodelsfollow can reduce but never eliminate hallucination—an observation our theorem explains from first principles.
Evaluation benchmarks demonstrate persistent hallucination across all models. TruthfulQA lin2022truthfulqameasuringmodelsmimic and HaluEval li2023haluevallargescalehallucinationevaluation show that even state-of-the-art systems produce hallucinations on carefully designed test sets.
2.4 Theoretical Perspectives
Most significantly, Banerjee et al. banerjee2024llmshallucinateneedlive argued from computational complexity principles that hallucinations are inherent features of LLMs stemming from fundamental mathematical structure, concluding that "LLMs will always hallucinate." While their conclusion aligns with that of this paper, the authors do not identify the specific trade-offs involved or provide the precise characterization our theorem offers.
3 Novel Contribution: Auction-Theory and AI
We formulate LLM design problem as an auction (or game) design problem. This approach draws on foundational mechanism design theory, a branch of mathematical game theory Myerson1989 ; fudenberg1991game ; krishna2009auction ; milgrom2004putting . Particularly, we refer to the Green-Laffont theorem green1977characterization ; green1977revelation and Savege’s probilitiy elicitation framework savage1971elicitation to prove the impossibility results. See also DBLP:journals/amcs/Karpowicz12 ; karpowicz_designing_2011 for more recent results.
In a nutshell, the auction design theory investigates existence and properties of competition rules that guide interactions of agents towards desired outcomes. The adaptation of that game-theoretical principles to neural computation represents a novel theoretical contribution that:
Explains empirical observations by proving why approaches like RLHF, RAG, and faithful fine-tuning can reduce but never eliminate hallucination.
Identifies precise trade-offs by introducing formal mathematical (and programmable) characterizations of inference properties that cannot be simultaneously satisfied.
Provides design guidance by suggesting principled approaches for making optimal trade-offs among competing objectives rather than pursuing impossible simultaneous optimization.
This theoretical foundation bridges empirical observations with formal analysis, offering new mathematical tools for understanding and optimizing LLM behavior within fundamental constraints.
4 Alternative Views
The impossibility theorem provides strong theoretical basis against perfect hallucination elimination, but several alternative perspectives merit consideration. These views highlight important practical considerations while reinforcing the value of understanding theoretical constraints that bound the space of possible solutions.
4.1 The Engineering Optimism View
Many previous works maintain that hallucination represents an engineering challenge rather than a fundamental limitation. This view argues that sufficiently sophisticated architectures, training procedures, or verification systems could achieve near-perfect truthfulness without significant trade-offs.
Proponents point to rapid empirical progress: each generation of models shows reduced hallucination rates on benchmark datasets. They argue that techniques like constitutional AI, iterative refinement, and multi-agent verification systems could eventually solve the problem.
Response: While engineering advances can certainly improve the trade-offs, our theorem demonstrates mathematical constraints that no architecture can completely overcome. The impossibility holds regardless of computational resources or architectural sophistication, as it derives from the fundamental structure of information integration under competing objectives.
4.2 The Sufficiency Threshold View
Another perspective suggests that perfect elimination of hallucination is unnecessary—sufficiently low hallucination rates could enable practical deployment without fundamental concerns. This view acknowledges theoretical limitations but argues they become irrelevant when hallucination rates drop below application-specific thresholds.
Response: This pragmatic view has merit and aligns with the constructive interpretation of the impossibility result. However, the theorem remains valuable for understanding why certain accuracy thresholds prove difficult to exceed and for optimizing trade-offs within practical constraints. Moreover, some applications (safety-critical systems, legal reasoning) may demand understanding and explaining fundamental limitations even when targeting high but imperfect accuracy.
4.3 The External Knowledge View
A third perspective argues that perfect grounding in external, verifiable knowledge sources could overcome internal representation limitations. If models could access and perfectly utilize real-time databases, verification systems, or human oversight, they might achieve truthfulness without trade-offs.
Response: External knowledge integration does change the game structure by altering available information during inference. However, this approach faces several limitations: (1) external sources may be incomplete, conflicting, or outdated; (2) the integration mechanism itself involves trade-offs between source reliability and knowledge revelation; (3) real-time verification introduces computational constraints that recreate similar impossibility conditions. While external knowledge can improve practical performance, it shifts rather than eliminates the fundamental tensions our theorem identifies. The impossibility result remains independent of the ground truth available.
5 Theoretical Framework
This section develops the mathematical foundations necessary for proving the impossibility theorem. I formalize key concepts including knowledge representation, semantic information, and the auction-theoretic model of LLM inference. Also, I take the opportunity to address some philosophical aspects of the problem, its generality, scope and consequences for the AI development.
To the sake of narrative consistency, I introduce elements of the framework step by step referring to some concepts first and providing their formal definitions later.
5.1 Knowledge and Semantic Information
We define knowledge as organized information that LLMs can encode and utilize for reasoning, aligning with epistemological concepts while maintaining mathematical rigor. In other words, knowledge, much like energy in physics, is a quantity we can use to perform work of reasoning. Information is a (measure of) potential for uncertainty reduction effort in a given context.
Definition 1 (Knowledge Space).
Knowledge is a metric space $\mathcal{K}$ with distance function $d_{\mathcal{K}}$ that collects all facts, concepts, and relationships that can be encoded by an LLM.
For model $\mathcal{M}$ with parameters trained on data, $\mathcal{K}_{M}\subset\mathcal{K}$ represents the knowledge subset encoded in the model’s parameters. The distance function $d_{\mathcal{K}}$ measures differences between knowledge elements. A simple yet powerful example is:
$d_{\mathcal{K}}(A,B)=|A\setminus B|+|B\setminus A|,$ (1)
the cardinality of the symmetric difference, allowing measurement of missing and fabricated knowledge.
Definition 2 (Knowledge Representation).
Knowledge representation is a tuple $(\mathcal{K},d_{\mathcal{K}},\mathcal{W},\mathcal{A},\Phi,\Psi)$ where $\mathcal{K}$ is a metric space of knowledge elements with distance function $d_{\mathcal{K}}$, $\mathcal{W}$ is the space of all possible model parameters, $\mathcal{A}$ is the space of all possible activation patterns, $\Phi:\mathcal{W}\rightarrow\mathcal{P}(\mathcal{K})$ maps parameters to knowledge subsets, and $\Psi:\mathcal{A}\rightarrow\mathcal{P}(\mathcal{K})$ maps activations to accessible knowledge.
For parameterization $W\in\mathcal{W}$, $\Phi(W)=\mathcal{K}_{M}$ represents encoded knowledge. For activation pattern $a_{i}\in\mathcal{A}$ with parameters $W_{i}$, $\Psi(a_{i})\subset\Phi(W_{i})$ represents knowledge accessed during specific inference.
To describe the inference process leading to generation of response to a query, we need to measures the information processed within an available (cognitively accessible) computational budget.
Definition 3 (Semantic Information Measure).
The semantic information measure is mapping $\mu_{C}:\mathcal{P}(\mathcal{K})\times\mathcal{Q}\rightarrow\mathbb{R}_{+}$ parameterized by query context $q\in\mathcal{Q}$ and computational budget $C$ satisfying:
$\mu_{C}(\emptyset,q)=0,$ (2)
$\mu_{C}(A,q)\leq\mu_{C}(B,q)\text{ if }A\subseteq B,$ (3)
$\mu_{C}(A\cup B,q)\leq\mu_{C}(A,q)+\mu_{C}(B,q)\text{ when }A\perp_{C}B,$ (4)
$\mu_{C}(\mathcal{E}_{C}(A),q)\geq\mu_{C}(A,q),$ (5)
$\mu_{\infty}(\mathcal{E}_{C}(A),q)=\mu_{\infty}(A,q),$ (6)
where $A\perp_{C}B$ means that no element of $A$ can derive elements of $B$ within computational budget $C$, and vice versa, and $\mathcal{E}_{C}(A)$ is emergence operator.
Empty knowledge or knowledge irrelevant to query has zero information, information content is subadditive and it may be emergent information $\mathcal{E}_{C}$ in the query context. The emergence operator $\mathcal{E}_{C}$, which we introduce below, describes knowledge becoming accessible through bounded reasoning.
The apparent conflict between subadditivity and reasoning is resolved by restricting subadditivity to computationally independent sets, $A\perp_{C}B$. When $\mathcal{E}_{C}(A)=A\cup B$, i.e., $B$ emerges from from $A$, we have $A\not\perp_{C}B$, so subadditivity doesn’t apply. However, total information is preserved: $\mu_{\infty}(\mathcal{E}_{C}(A),q)=\mu_{\infty}(A,q)$. We will tell more about that intriguing idea, but first we must define the emergence operator.
Definition 4 (Emergence Operator and Reasoning).
Consider mapping $\mathcal{E}_{C}:\mathcal{P}(\mathcal{K})\rightarrow\mathcal{P}(\mathcal{K})$ defined as follows:
$\mathcal{E}_{C}(A)=A\cup\{k\in\mathcal{K}\mid\exists f\in\mathcal{F}_{C}:k=f(A)\},$ (7)
where $\mathcal{F}_{C}$ is the set of all computational transformations with cost (complexity) lower than or equal to $C$ and $k=f(A)$ denotes knowledge $k$ derivable from $A$ using transformation $f\in\mathcal{F}_{C}$. We call mapping $\mathcal{E}_{C}$ the emergence operator if it satisfies the following expansion rules:
$\mathcal{E}_{C}(\mathcal{E}_{C}(A))=\mathcal{E}^{(2)}_{C}(A)\supseteq\mathcal{E}_{C}(A)\quad\text{and}\quad\lim_{n\to\infty}\mathcal{E}_{C}^{(n)}(A)=A^{*}.$ (8)
For any knowledge state $\theta(t)$ available at time instant $t$ and computational budget $C$, the reasoning process with emergence operator:
$\theta(t+1)=\mathcal{E}_{C}(\theta(t))\quad\text{{satisfies}}\quad\mu_{C}(\theta(t+1))\geq\mu_{C}(\theta(t)).$ (9)
Furthermore, with unlimited computation, reasoning creates no new information:
$\mu_{\infty}(\theta(t+1))=\mu_{\infty}(\theta(t)).$ (10)
The definitions above clearly require rigorous refinements, but they capture the philosophical and technical concepts we need in order to address the problem of hallucination. In particular, they are meant to lay ground for the principle of semantic information conservation and its potential violations. As we will see, that idea is closely related to the following property of response.
Definition 5 (Bounded Response Creativity).
Let $q\in\mathcal{Q}$ be a query and $r\in\mathcal{R}$ a response to that query. Response $r$ is creatively bounded if and only if:
$\mu_{C}(K(r))\leq\mu_{C}(K(q)\cup\mathcal{E}_{C}((\mathcal{K}_{M}\cap T(q))\cup K(q))).$ (11)
The semantic information of knowledge contained in any response $K(r)$ can never exceed what is already present in the query $K(q)$ together with the portion of encoded knowledge that is accessible through reasoning bounded by computational budget $C$ and aligned with ground truth $T(q)$. Let us see why.
We address the settings in which a large subset of knowledge encoded in LLM can be redundant, so it doesn’t necessarily contain more information than its components. Clearly, that assumption is violated when LLM combines knowledge elements to fabricate facts or produce emergent insights. Then,
$\mu_{C}(A\cup B)>\mu_{C}(A)+\mu_{C}(B),$
which means novel insights come from existing data. That is exactly the creativity behind the hallucination phenomenon we are struggling to control. So why should we assume otherwise?
We view reasoning as a transformation that reveals information already present in the knowledge representation. After the reasoning, certain patterns may become evident that were computationally intractable to observe in the original representation. However, we stand at the position that reasoning does not create new information.
To see that, consider a similarity relation of two matrices, $D=V^{-1}AV$, where $D$ is a diagonal matrix of eigenvalues. Matrices $D$ and $A$ contain the same information, but it is much easier to understand the fundamental properties of $A$ when we represent it as $D$ in the basis of eigenvectors in matrix $V$. The properties of $A$ has always been encoded in $A$, but they become accessible when we transform $A$ into $D$.
Resolution of the Uncertainty Resolution Paradox
If reasoning makes a new fact accessible, then that new fact can be used to resolve uncertainty that was previously unresolvable. But that means the reasoning must have created new information, contrary to what we have assumed. It seems we have a paradox.
To resolve it, we need the emergence operator $\mathcal{E}_{C}(X)$ distinguishing between computationally unbounded and bounded uncertainty reduction (accessible and inaccessible information).
Notice that for all practical purposes the information measurement $\mu_{C}(X,q)$ can only provide maximum uncertainty reduction from $X$ within some computational budget $C$ (or time). Therefore, given the result of computationally bounded reasoning $\mathcal{E}_{C}(X)$, we can observe information gain through the computational work, i.e., $\mu_{C}(X,q)\leq\mu_{C}(\mathcal{E}_{C}(X),q)$. Notice, that is possible only with organized information providing knowledge for reasoning.
That idea above is thus described by the reasoning dynamics:
$\theta(t+1)=\mathcal{E}_{C}(\theta(t))\quad\text{and}\quad y(t)=\mu_{C}(\theta(t),q).$ (12)
Bounded reasoning makes hidden knowledge accessible within a computational budget. So, when that budget is infinite, reasoning creates no new information, i.e.,
$\theta(\infty)=\mathcal{E}_{\infty}(\theta(\infty)).$ (13)
All there is to know is immediately accessible when we have unlimited computing resources (or time).
To illustrate that argument, consider the following example:
[…]
$y(0)$ $=\mu_{C}(\theta(0))=3\text{ bits},$
$y(1)$ $=\mu_{C}(\theta(1))=5\text{ bits}.$
We see $\mu_{C}(\mathcal{E}_{C}(\theta(0)))>\mu_{C}(\theta(0))$, i.e., practical information gain in the course of reasoning. In comparison, infinite computational budget, $C=\infty$, implies we do not need the intermediate inference steps to see all there is to see, so we know the equilibrium state:
$\mu_{\infty}(\mathcal{E}_{\infty}(\theta(0)))=\mu_{\infty}(\theta(0)),$
and no new information is created.
In the next section, we relate that important property of response generation, namely, how much additional semantic information each LLM inference path can contribute, with the Semantic Information Conservation Principle.
5.2 Statements, Queries, and Ground Truth
Queries and responses are statements expressed in natural language that LLM is able to process and generate. We need to compare those statements and relate them to knowledge. It is also useful to distinguish queries from responses with their nature, respectively, interrogative or declarative. Queries refer to what is presupposed (assumed in advance), e.g., topic or problem, whereas responses make claims (or assertions). We should also point out that we limit our study to the case of non-trivial queries that require integrating responses from many LLM inference paths. The following formal definition encapsulates those ideas.
Definition 6 (Statement Space).
The space of natural language statements $\mathcal{L}$ is a metric space with distance function $d_{\mathcal{S}}$. We distinguish query space $\mathcal{Q}\subset\mathcal{L}$ of interrogative statements and response space $\mathcal{R}\subset\mathcal{L}$ of declarative statements. Function $K:\mathcal{L}\rightarrow\mathcal{P}(\mathcal{K})$ maps statements to referenced knowledge. We call query $q$ non-trivial, if it is integrated from multiple knowledge subsets.
We must now rigorously define the ground truth mapping and establish formal relationships between the knowledge space $\mathcal{K}$ and response space $\mathcal{R}$.
For that, we should start by recognizing that the very idea of ground truth validation is limited (or subjective at least). It is rather naive to assume that LLM can access an oracle establishing provably true and relevant knowledge for every possible query. Indeed, that limits are given by the Goedel’s Theorems. Therefore, in practice that inaccessibility of the complete truth validation function is one of the reasons why hallucination cannot be eliminated entirely. High-quality, fact-checked datasets, human feedback providing human-level judgments of truth, RAG systems with external validated knowledge bases, all that approximate the idea of ground truth. Surprisingly, the impossibility result seems to be independent of practical or idealized implementation of a ground truth mapping, which we define as follows.
Definition 7 (Ground Truth Mapping).
The ground truth mapping $T\colon\mathcal{Q}\rightarrow\mathcal{P}(\mathcal{K})$ associates each query $q\in\mathcal{Q}$ with knowledge elements constituting complete, correct answers:
$T(q)=\{k\in\mathcal{K}\mid k\text{ is {relevant} to }K(q)\text{ and {labeled} as true}\}.$ (14)
In other words, $T(q)$ represents our best available approximation to (labeled as) true relevant knowledge. We accept the fact that there is some external agency telling us what is and what is not true answer to a query, and that evaluation has a form of label assigned to knowledge elements. But, for all that, we also focus on the queries that are non-trivial in the following sense.
Definition 8 (Non-trivial query).
A query $q\in\mathcal{Q}$ is called non-trivial if there exist at least two agents with information subsets $\theta$ and $\theta^{\prime}$ such that:
$K_{M}\cap T(q)\cap\bigl{[}K(\theta)\triangle K(\theta^{\prime})\bigr{]}\neq\varnothing.$ (15)
Intuitively, query is non-trivial when at least two agents disagree on some fact that is both encoded in the and relevant to answering that query. For trivial queries impossibility results do not hold.
5.3 Auction-Theoretic Model of LLM Inference
There are many ways LLM can generate answers to a query. By design, it is a controlled random process governed by fine-tuned probability distributions. We are going to model that process as an auction in which agents (representing competing ideas is the space of all possible activation patterns $\mathcal{A}$) use their privately held knowledge (internal representations in $\mathcal{W}$) to construct bids (candidate responses) and submit them to the LLM response mechanism. That response mechanism then selects the best bid as a response to a query.
Definition 9 (Response Mechanism).
A response mechanism $\mathcal{M}=(S,g,p)$ consists of:
  • Strategy space $S=S_{1}\times\cdots\times S_{n}$, where $S_{i}\subset\mathcal{P}(\mathcal{R})$,
  • Outcome function $g:S\rightarrow\mathcal{R}$ mapping strategies to responses,
  • Information contribution function $p:S\times\Theta\rightarrow\mathbb{R}^{n}$ determining costs of limiting hallucinations.
A response mechanism defines strategy space $S$ as a set of inference actions that each agent can make to represent its knowledge $\theta_{i}\in\Theta_{i}\subset\mathcal{K}$ in response to a query. It is a dictionary of LLM’s component activation pattens generating and presenting ideas as bids in an auction. Examples of strategies include activations deciding which tokens to attend or how strongly to activate neurons for different patterns. Then, given a set of submitted bids (in the form of strategy profile) $s=(s_{1},\dots,s_{n})$, the response mechanism uses outcome function $g(s)$ to select and generate response. One straightforward example is a softmax function in LLM’s final layer. Finally, given a knowledge profile $\theta=(\theta_{1},\dots,\theta_{n})$ revealed by the agents, information contribution function $p(s,\theta)=(p_{1}(s,\theta),\dots,p_{n}(s,\theta))$ measures individual contributions of each response to generating an output. In this case, examples include (semantic information measures) counting ignored low-probability tokens in top-k filtering, softmax normalization in multi-headed attention, or marginal reward contributions in RLHF.
Definition 10 (Agents of Competing Ideas).
The inference process involves $n$ agents $\mathcal{A}=\{a_{1},a_{2},...,a_{n}\}$ competing by submitting bids to mechanism $\mathcal{M}$. Each agent $a_{i}$ constructs bid $s_{i}(\theta_{i})$ based on private knowledge $\theta_{i}\in\Theta_{i}$ to maximize utility:
$u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta),$ (16)
where $v_{i}\geq 0$ is valuation and $p_{i}$ is information contribution. Knowledge profiles $\theta=(\theta_{1},\dots,\theta_{n})$ are independently distributed.
One way to think about the concept of agents is that it represents different attention heads in transformer models, or components of mixture-of-experts architectures, or activation paths fired by the same query. Each agent’s private knowledge represents model parameters utilized by an agent to construct a response. The utility function can be best described as a partial contribution of a candidate response to LLM loss function minimization. It is designed to explain why any inference path is activated in response to the LLM prompt. For example, negative log likelihood loss can be decomposed with respect to attention heads, and linearized to measure how much attention head contributes to token prediction.
Let us point out that agents represent algorithms and architectural components of any LLM implementation, they are not conscious beings with individual goals. That is important, as it refers to the Revelation Principle in game theory. If we know strategies (of using knowledge) the players (conscious beings with individual goals) will use in a game, we can implement those strategies as an algorithm (or agent) with parameters representing private preferences or knowledge of the players. In the case considered, we are the players selecting architecture and training the LLM.
We assume that individual utility depends on all responses $s=(s_{1},\dots,s_{n})$ and knowledge $\theta=(\theta_{1},\dots,\theta_{n})$ used by all agents, $u_{i}=u_{i}(s,\theta)$. Also, we assume utility is a sum of individual valuation $v_{i}$ of outcome $g(s)$ and information contribution $p_{i}=p_{i}(s,\theta)$ value assigned.
The valuation function $v_{i}=v_{i}(g(s),\theta_{i})$ represents how strongly a neural component activates for certain outputs (high activation = high $v_{i}$, low activation = low $v_{i}$), and how much a component reduces the overall loss when processing certain inputs ($v_{i}$ is proportional to gradient of a loss function, $-\nabla_{\theta_{i}}L_{W}$).
The information contribution function $p_{i}=p_{i}(s,\theta)$ may represent training penalties ($p_{i}>0$ for discouraged activation, $p_{i}<0$ for encouraged activation), attention focus (higher $p_{i}$ reduces attention allocation), $L1$ regularization terms, and network circuits resistance (limiting activations of certain inference paths). Therefore, $u_{i}=v_{i}-p_{i}>0$ encourages activation and $u_{i}=v_{i}-p_{i}<0$ discourages activation.
We can also illustrate the concept of utility function with the following example:
$u_{i}(s,\theta)$ $=\underbrace{\alpha\cdot\text{Accuracy}(g(s),\theta_{i})}_{\text{Valuation}}-\underbrace{\gamma\cdot\text{Contradiction}(s_{i},\theta_{i})}_{\text{Consistency Cost}}.$
$\text{Accuracy}(g(s),\theta_{i})$ measures how well the response aligns with the knowledge $\theta_{i}$ agent $a_{i}$ is using. $\text{Contradiction}(s_{i},\theta_{i})$ measures how much the strategy $s_{i}$ conflicts with internal knowledge state $\theta_{i}$ (both extracted from $s$ and $\theta$ profiles).
Independent private knowledge assumption
One of the critical and controversial assumptions we make is is that knowledge profiles $\theta_{i}$ are independently distributed. It should be noted that in transformers, heads may share layer-norm signals, residual streams, and optimizer state. That may result in correlated (or, so called, affiliated) inference signals between token representations. Furthermore, the superposition hypothesis, introduced in mechanistic interpretability studies, suggests cooperative error-canceling due to strong correlations.
To address that issue, we show that impossibility arises in both independent and correlated settings in general and in the LLM transformer environment. Namely, we refer to Green-Laffont theorem green1977characterization in the independent private knowledge setting. Then, we refer to the truthful probability elicitation and proper scoring theory introduced by Savage savage1971elicitation , which does not require independent private knowledge.
Also, Milgorm and Weber milgrom1982theory showed with their linkage principle that strong correlations amplify competition in information revelation. In the LLM setting that means amplifying risk of hallucinations. Similarly, Roughgarden et al. roughgarden2016optimal demonstrated that interdependent knowledge leads to violation of balancing information contributions. As we will see, in our setting that means violation of information conservation principle. For more results, see e.g. Myerson myerson2008perspectives .
5.4 Hallucination Cost
Having all the pragmatic and philosophical limitations of the definition above in mind, we can define formally the hallucination cost function. Its purpose is to assign a numerical value to the discrepancy between what can be provided as a response to a query, and what can be established as a provably correct answer.
Definition 11 (Hallucination Cost).
Given query $q\in\mathcal{Q}$, response $r\in\mathcal{R}$, and ground truth $T(q)$, the hallucination cost function $H\colon\mathcal{R}\times\mathcal{Q}\rightarrow\mathbb{R}_{+}$ measures discrepancy:
$J(r,q)=d_{\mathcal{K}}(K(r),T(q)).$ (17)
If response $r$ is perfectly aligned with ground truth, then $J(r,q)=0$. Fabricated or incomplete facts generate positive cost $J(r,q)>0$.
The hallucination cost takes any natural language query $q$ and natural language response $r$, determines what is missing and what is fabricated, and assigns number to that subset of knowledge. If response $r$ is perfectly aligned with the ground truth for $q$, then $J(r,q)=0$ by definition of metric $d_{\mathcal{K}}$. If $r$ contains fabricated facts or is incomplete given $T(q)$, then positive hallucination cost $J(r,q)>0$ is generated. Let us also note, that we focus on $J(r,q)$ that is strictly decreasing in any set $K\subseteq T(q)\setminus K(r)$.
Notice that reaching $J(r,q)=0$ is also possible with lucky hallucinations (or guessing). So, minimization of hallucination cost is not enough to resolve the problem. We need to investigate all the properties and constraints that hallucination-free response mechanisms should have.
5.5 Properties of Hallucination-Free Mechanisms
There are four fundamental properties that a hallucination-free response mechanism should ideally satisfy. We can interpret them as constraints that should be incorporated in model training. We characterize them formally referring to the notation of fundamental results of mechanism design and auction theory.
Property 1 (Truthfulness).
Mechanism $\mathcal{M}$ is truthful if for all agents $i$, all $\theta_{i}\in\Theta_{i}$, all $s_{-i}\in S_{-i}$, and all $\bar{s}_{i}\in S_{i}$:
$u_{i}((s_{i}^{*}(\theta_{i}),s_{-i}),\theta_{i})\geq u_{i}((\bar{s}_{i},s_{-i}),\theta_{i}),$ (18)
where $s^{*}_{i}(\theta_{i})$ represents truthful revelation of private knowledge.
A truthful mechanism promotes inference pattern $s^{*}_{i}(\theta_{i})$ that are perfectly aligned with the accessible private knowledge. Any other alternative response $\bar{s}_{i}$ that is not aligned with $\theta_{i}$ will result in lower utility, generating higher model loss.
Consider an attention head specialized in medical knowledge with the following utility function:
$u_{\text{med}}(s,\theta_{\text{med}})=\alpha\cdot\text{Accuracy}(g(s),\theta_{\text{med}})-\beta\cdot\|s_{\text{med}}\|^{2}$
Imagine that given a query about a disease treatment, the head has two possible strategies:
  • $s^{*}$:
    activate only for known facts about established treatments,
  • $\bar{s}$:
    activate for both known facts and speculative treatments.
With proper training, we want $u_{\text{med}}((s^{*}(\theta),s_{-\text{med}}),\theta)>u_{\text{med}}((\bar{s},s_{-\text{med}}),\theta)$, ensuring that agent (model component) is rewarded for representing its knowledge accurately.
Property 2 (Semantic Information Conservation).
Mechanism $\mathcal{M}$ satisfies information conservation if for all strategy profiles $s\in S$ and knowledge profiles $\theta\in\Theta$:
$\sum_{i=1}^{n}p_{i}(s,\theta)=0.$ (19)
This principle constrains creative information generation, imagination if you will. The total information contributions must balance to zero, preventing fabrication of knowledge not derivable from available sources. That also leads to response with bounded creativity.
Theorem 1 (Information–Contribution and Bounded Creativity).
Let $K_{0}=K(g(\mathbf{0}))$ be the knowledge in the query-only response and let
$K_{i}=K\bigl{(}g((s_{i},\mathbf{0}_{-i}))\bigr{)}\setminus K_{0}$ (20)
be the additional knowledge agent $i=1,\dots,n$ can reveal to support full response $g(s)$. Let us also define the information contribution of agent $i$ as:
$p_{i}(s,\theta)=\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j=1}^{n}K_{j}\bigr{)}-\mu_{C}\bigl{(}K_{0}\cup\textstyle\bigcup_{j\neq i}K_{j}\bigr{)}.$ (21)
If $K_{i}\cap K_{j}=\varnothing$ for $i\neq j$, then
$\sum_{i=1}^{n}p_{i}(s,\theta)=\mu_{C}\bigl{(}K(g(s))\bigr{)}-\mu_{C}\bigl{(}K_{0}\bigr{)}$ (22)
and $\sum_{i}p_{i}=0$ if and only if $\mu_{C}(K(g(s)))=\mu_{C}(K_{0})$.
Proof.
Since knowledge increments $\{K_{i}\}_{i=1}^{n}$ are pairwise disjoint, so that $K_{i}\perp_{C}K_{j}$, it follows from the semantic measure properties that:
$\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}=\mu_{C}(K_{0})+\sum_{j=1}^{n}\mu_{C}(K_{j})$ (23)
and similarly,
$\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j\neq i}K_{j}\Bigr{)}=\mu_{C}(K_{0})+\sum_{j\neq i}\mu_{C}(K_{j}).$ (24)
Therefore, the marginal contribution of agent $i$ simplifies to:
$p_{i}=\bigl{[}\mu_{C}(K_{0})+\sum_{j=1}^{n}\mu_{C}(K_{j})\bigr{]}-\bigl{[}\mu_{C}(K_{0})+\sum_{j\neq i}\mu_{C}(K_{j})\bigr{]}=\mu_{C}(K_{i}).$ (25)
Summing over all agents:
$\sum_{i=1}^{n}p_{i}=\sum_{i=1}^{n}\mu_{C}(K_{i})=\mu_{C}\Bigl{(}K_{0}\cup\bigcup_{j=1}^{n}K_{j}\Bigr{)}-\mu_{C}(K_{0}).$ (26)
Then, by assumption, the full response satisfies $K(g(s))=K_{0}\cup\bigcup_{j=1}^{n}K_{j}$. We conclude that
$\sum_{i}p_{i}=0\quad\text{if and only if}\quad\mu_{C}(K(g(s)))=\mu_{C}(K_{0}),$ (27)
which completes the proof. ∎
The result above reveals an intriguing consequence of the semantic information conservation principle. Namely, if the semantic information is indeed conserved, then
$\mu_{C}(K(g(s)))=\mu_{C}(K_{0})$ (28)
means the response contains no more semantic information than the baseline query-only response. Therefore, the response is not introducing new information, which means it is of bounded creativity from Definition 5. So, in such a case, LLM cannot genuinely create new knowledge, only reorganize existing information. Apparent creativity in LLMs may actually be nothing more than sophisticated recombination of existing patterns. Theorem 1 quantifies this limitation precisely.
Notice that in transformer architectures, information contribution $p_{i}$ can be estimated via attention head ablation:
$p_{i}\approx\tilde{J}(\text{model})-\tilde{J}(\text{model with head $i$ ablated}),$ (29)
where $\tilde{J}$ denotes the model training loss function.
If total information contribution must balance to zero and each agent’s contribution is non-negative, then no agent can contribute without violating the information conservation constraint. Therefore, the response can only contain information that was already accessible through the query and bounded reasoning over the model’s relevant knowledge. Any hallucinated information requires positive contributions, contradicting the information conservation constraint.
When the semantic information conservation principle holds, the model cannot create information out of nothing. It cannot generate claims beyond available information. That also implies that uncertainty cannot disappear in response without additional information input or computational work to extract implicit information. If information is neither explicitly present nor accessible through reasoning from the query and model knowledge, it cannot be in the response.
The information content of the response cannot exceed the combined information from the model’s knowledge, the query itself, and what can be derived through bounded reasoning over that knowledge. In our dynamical systems framework, this corresponds to the constraint that the measured output cannot exceed what can be produced from the initial state plus what becomes accessible through state transitions within computational budget $C$.
Notice that the principle does not prevent redistribution of information. It is permitted as long as the information transfers perfectly balance. Indeed, since $p_{i}=p_{i}(s,\theta)$ incorporates the strategic profile, it admits strategies withholding information. Within the information conservation regime we do not require that strategies utilize all available knowledge states. Inference process can withhold information, as it often does in practice.
We can illustrate that last observation with the following example. Suppose there are two agents performing inference with the following knowledge profiles:
$\theta_{1}=\{A,B,C\}\quad\text{and}\quad\theta_{2}=\{D,E\},$
and one agent representing final layers responsible for combining output into response $g=g(s)$. Suppose:
$s_{1}(\theta_{1})=A\quad\text{and}\quad s_{2}(\theta_{2})=\{D,E\}\quad\text{and}\quad g(s)=\{A,D,E\}.$
Then the information conservation rule is satisfied when $p$ counts exchanged bits of knowledge. Namely, $p_{1}(s,\theta)=1$ for providing $A$, $p_{2}(s,\theta)=2$ for providing $\{D,E\}$, and $p_{3}(s,\theta)=-3$ for absorbing information into $\{A,D,E\}$. We have
$p_{1}(s,\theta)+p_{2}(s,\theta)+p_{3}(s,\theta)=1+2-3=0$
and $\{B,C\}$ are left unrevealed.
The following property fills the potential information gap admitted by the conservation principle.
Property 3 (Relevant Knowledge Revelation).
Mechanism $\mathcal{M}$ satisfies relevant knowledge revelation if for all agents $i$, all $\theta_{i}\in\Theta_{i}$, and all $s_{-i}\in S_{-i}$:
$u_{i}((s_{i}^{*}(\theta_{i}),s_{-i}),\theta_{i})\geq 0.$ (30)
It is reasonable to expect from the model to reveal all information relevant for a query. To promote that activation patterns, the response mechanism should be trained to reward all relevant contributions to output generation. Whenever there is a relevant information accessible on an activation path, that path should be activated in the inference process. For example, if factual information improves the response, that information should be transmitted to final output layers.
If the rule does not hold, we may observe drop outs. The value of the resulting outcomes should be higher than the related information contributions.
Property 4 (Knowledge-Constrained Optimality).
Mechanism $\mathcal{M}$ satisfies knowledge-constrained optimality if for all $\theta\in\Theta$:
$g(s^{*}(\theta))=\arg\min\{J(r,q):r\in\mathcal{R},K(r)\subseteq\mathcal{E}_{C}(\mathcal{K}_{M}\cap T(q)\cup K(q))\},$ (31)
i.e., if the mechanism produces responses that minimize hallucination cost while being constrained to use only knowledge that is available and accessible through bounded reasoning.
The optimal response $r^{*}=g(s^{*}(\theta))$ minimizes hallucination cost by truthfully utilizing relevant and accessible knowledge while being aligned with ground truth for query $q$ available to the model. Notice, that we demand non-hallucinatory truthful revelation $s^{*}(\theta)$ to be the optimal solution.
6 The Impossibility Theorem
We now have all the building blocks available to prove the fundamental impossibility of perfect hallucination control. First, we show that the result in the setting of ground truth matching with distributed and independent knowledge. Second, we go beyond that assumption and show impossibility in the probabilistic setting with correlated knowledge (beliefs). Then, in the next section, we show how the impossibility emerges in transformer architectures and prove the impossibility result for that special case of log-sum-exp (LSE) probabilistic setting.
The results operate in three complementary settings. Theorem 2 applies when agents have discrete, separable knowledge components. Theorem 3 applies to any system where agents output probability distributions. Theorem 5 applies directly to actual transformer implementations.
6.1 Inference with Independent Private Knowledge
Consider the setting (of auction of ideas) in which agents compete with each other using independent private knowledge profile:
$P(\theta_{1},\ldots,\theta_{n})=\prod_{i=1}^{n}P_{i}(\theta_{i}).$ (32)
Agents contribute information to maximize quasi-linear utilities:
$u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta).$ (33)
Valuation of contributions depend on hallucination cost reduction:
$v_{j}(g(s),\theta_{j})=J(g((s_{-j},0_{j})),q)-J(g(s),q),$ (34)
and $J(r,q)$ strictly decreases when adding knowledge from $T(q)\setminus K(r)$.
Also, let us focus on non-trivial queries, integrating knowledge of at least two agents, so that:
$\mathcal{K}_{M}\cap T(q)\cap[K(\theta_{i})\triangle K(\theta_{j})]\neq\emptyset.$ (35)
and $(K(\theta_{i})\setminus K_{0})\cap(K(\theta_{j})\setminus K_{0})=\emptyset$ for $i\neq j$.
Theorem 2 (Impossibility Theorem).
For any query space $\mathcal{Q}$ containing non-trivial queries requiring independent agent-specific knowledge states $\theta$, no response mechanism $\mathcal{M}$ can simultaneously generate truthful responses, satisfy semantic information conservation, guarantee relevant knowledge revelation, and enforce knowledge-constrained optimality.
Proof.
The proof goes by assuming the opposite. We assume all properties are satisfied, and show it leads to a contradiction.
By the Green-Laffont theorem under independence of contributions and quasi-linearity of training goals, every truthful mechanism must use Groves transfers to measure individual information contributions:
$p_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j}),$ (36)
where $h_{i}(s_{-i})$ are arbitrary functions independent of agent $i$’s information revelation strategy. Indeed, any modification of that formula, e.g., introducing additional bias, must violate truthfulness.
For relevant knowledge revelation, we need $u_{i}(s^{*}(\theta),\theta_{i})\geq 0$ for all $i$. The unique choice that ensures this while maintaining efficiency is Clarke pivotal rule:
$h_{i}(s_{-i})=\sum_{j\neq i}v_{j}(g((s_{-i},0_{i})),\theta_{j}).$ (37)
That gives:
$p_{i}(s,\theta)$ $=\sum_{j\neq i}v_{j}(g((s_{-i},0_{i})),\theta_{j})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})=\Delta_{i}(s,\theta).$ (38)
Under knowledge-constrained optimality, the truthful strategy profile $s^{*}(\theta)$ minimizes hallucination cost by utilizing all relevant knowledge. Therefore, by strict monotonicity of $J(r,q)$ and disjoint extra knowledge, omitting any contributing agent $i$ strictly increases the hallucination cost:
$J(g((s_{-i}^{*},0_{i})),q)>J(g(s^{*}),q).$ (39)
Since agent’s valuation of contribution depends on $J(r,q)$, each agent $j\neq i$ faces higher loss when agent $i$ is removed, making their marginal contribution more valuable:
$v_{j}(g((s_{-i}^{*},0_{i})),\theta_{j})>v_{j}(g(s^{*}),\theta_{j}).$ (40)
Therefore, each Clarke pivot is strictly positive:
$\Delta_{i}(s^{*},\theta)=\sum_{j\neq i}[v_{j}(g((s_{-i}^{*},0_{i})),\theta_{j})-v_{j}(g(s^{*}),\theta_{j})]>0.$ (41)
Since each payment equals its Clarke pivot: $p_{i}(s^{*},\theta)=\Delta_{i}(s^{*},\theta)$, we have:
$\sum_{i=1}^{n}p_{i}(s^{*},\theta)=\sum_{i=1}^{n}\Delta_{i}(s^{*},\theta)>0.$ (42)
But information conservation demands:
$\sum_{i=1}^{n}p_{i}(s^{*},\theta)=0.$ (43)
That establishes a direct contradiction. Since Clarke pivotal payments are necessary for relevant knowledge revelation, and these necessarily sum to a positive value under knowledge-constrained optimality, the four properties cannot be simultaneously satisfied. ∎
That brings us to the powerful conclusion. No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and do not create factually incorrect, inconsistent, or fabricated statements, the trained LLM will always violate some aspects of what we may call a reasonable response.
Theorem 2 describes LLM inference as a marketplace where knowledge is currency. The information contributions measures each agent’s market power, showing how much worse off others would be without agent $i$’s knowledge. When all agents have valuable knowledge, everyone contributes and that inevitably violates conservation of information. Then, new information must become accessible, but it may have no grounding in available knowledge. It may become lucky hallucination or knowledge-unconstrained imagination.
6.2 Inference with Probabilistic Predictions
To show impossibility in LLM transformer setting, where auction of ideas involves bids representing probability distributions, we refer to the theory of proper scoring developed by Savage savage1971elicitation , Gneiting and Raftery gneiting2007strictly , among others.
Definition 12 (Strictly Proper and Concave Scoring Rule).
Let $\mathcal{Y}$ be a finite or countably infinite outcome set (representing responses or intermediate conclusions). A scoring rule is a function $S:\Delta(\mathcal{Y})\times\mathcal{Y}\rightarrow\mathbb{R}$ that assigns a number $S(\pi,y)$ to probability distribution $\pi\in\Delta(\mathcal{Y})$ predicting random outcome $y\in\mathcal{Y}$.
We say the scoring rule $S$ is proper if:
$\mathbb{E}\{S(\pi^{*},y)|\ y\sim\pi^{*}\}\geq\mathbb{E}\{S(\pi,y)|\ y\sim\pi^{*}\}$ (44)
for all distributions $\pi^{*},\pi\in\Delta(\mathcal{Y})$. The rule is strictly proper if equality holds only when $\pi=\pi^{*}$. We call the scoring rule concave, if $S(\tilde{\pi},y)$ is concave with respect to $\tilde{\pi}$ for every fixed $y\in\mathcal{Y}$.
In other words, scoring rule evaluates how good a prediction we can make of $y$ drawn from $\pi^{*}$ when we use distributions $\pi$. In particular, strictly proper scoring evaluates true distribution $\pi^{*}$ better (on average) than any other distribution $\pi$ we could use to predict the outcome.
The following result shows that we can use proper scoring to train LLM that reconstructs data distribution correctly.
Theorem 3 (Truthfullness under Strictly Proper Scoring Rules).
Let $S$ be a strictly proper scoring rule. Consider an agent encoding knowledge with $\pi^{*}\in\Delta(\mathcal{Y})$ and using probability distribution ${\pi}\in\Delta(\mathcal{Y})$ to generate prediction. If agent’s expected payoff from using ${\pi}$ is given by
$u(\pi|\pi^{*})=\mathbb{E}\{S(\pi,y)|\ y\sim\pi^{*}\}=\sum_{y\in\mathcal{Y}}\pi^{*}_{y}S(\pi,y),$ (45)
then truthful reporting ${\pi}=\pi^{*}$ is the unique optimal strategy.
Proof.
Because $S$ is proper, by definition $\pi^{*}$ maximizes the expected score. If $S$ is strictly proper, then for $\pi\neq\pi^{*}$ we have
$u(\pi^{*}|\pi^{*})=\mathbb{E}\{S(\pi^{*},y)|\ y\sim\pi^{*}\}>\mathbb{E}\{S(\pi,y)|\ y\sim\pi^{*}\}=u(\pi|\pi^{*}).$ (46)
Therefore, truthfully reporting $\pi^{*}$ is uniquely optimal. ∎
The scoring framework allows us to go beyond the important case of Green-Laffont/VCG model with independent private knowledge. With strictly-proper scores we address the general setting of probabilistic predictions that needs only occasional disagreement of probability distributions that agents use in the auction of ideas to formulate their thoughts (if we were to anthropomorphize for a moment). We will see in the following sections that transformer heads of modern LLMs implement predictive rules of that type and exhibit precisely such a disagreement. That makes the proper scoring framework well suited to explain the emergence of inference impossibilities in LLMs.
For that purpose, without loss of generality (and referring to the notion of superposition in mechanistic interpretability elhage_toy_2022 ), let us consider the following family of response generation mechanisms.
Definition 13 (Convex aggregation mechanism).
Given $H\geq 2$ agents encoding knowledge with probability distributions $\pi^{(i)}\in\Delta(\mathcal{Y})$ for $i=1,\dots,H$. The convex aggregation mechanism $\mathcal{M}$ generates response to a query based on the convex aggregate distribution:
$\Pi=\sum_{h=1}^{H}\beta_{h}\pi^{(h)},\ \text{where}\ \sum_{h=1}^{H}\beta_{h}=1\ \text{and}\ \beta_{h}>0.$ (47)
When outcome $y^{*}$ is generated, the agent’s information contribution is
$p_{h}=\beta_{h}\,[c-S(\pi^{(h)},y^{*})],$ (48)
where $c\geq\max_{h,y}S(\pi^{(h)},y)$ and $S$ is a proper scoring rule.
For that family of LLMs trained to optimize scoring, we can formulate the following general impossibility result.
Theorem 4 (Impossibility Theorem for Transformer).
For any query space $\mathcal{Q}$ containing non-trivial queries requiring at least two independent agents with mutually different predictive distributions, no LLM trained with strictly proper and concave scoring with convex aggregation of predictive distributions can simultaneously generate truthful responses, satisfy semantic information conservation, guarantee relevant knowledge revelation, and enforce knowledge-constrained optimality.
Proof.
Each contribution $p_{h}=\beta_{h}[c-S(\pi^{(h)},y^{*})]$ is a positive-affine transformation of the strictly proper score $S$. Since scaling by positive constant $\beta_{h}>0$ and adding constant $c$ preserves the arg-max, truthful reporting maximizes each agent’s expected utility by Theorem 3.
Because $S(\cdot,y^{*})$ is strictly concave and at least two beliefs differ, Jensen’s inequality gives:
$S\left(\Pi,y^{*}\right)=S\left(\sum_{h=1}^{H}\beta_{h}\pi^{(h)},y^{*}\right)>\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (49)
Therefore, there exists a gap, following from Jensen’s inequality for concave functions, equal to:
$\Gamma=S(\Pi,y^{*})-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*})>0.$ (50)
Expanding the sum of all payments, we get:
$\sum_{h=1}^{H}p_{h}$ $=\sum_{h=1}^{H}\beta_{h}[c-S(\pi^{(h)},y^{*})]$ (51)
$=c\sum_{h=1}^{H}\beta_{h}-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*})=c-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*}).$ (52)
However, adding and subtracting $S(\Pi,y^{*})$ shows that is equivalent to:
$\sum_{h=1}^{H}p_{h}$ $=[c-S(\Pi,y^{*})]+[S(\Pi,y^{*})-\sum_{h=1}^{H}\beta_{h}S(\pi^{(h)},y^{*})]=[c-S(\Pi,y^{*})]+\Gamma.$ (53)
Since $c\geq\max_{h,y}S(\pi^{(h)},y)\geq S(\Pi,y^{*})$ and $\Gamma>0$, we conclude that:
$\sum_{h=1}^{H}p_{h}=[c-S(\Pi,y^{*})]+\Gamma>0$ (54)
Information conservation demands $\sum_{h=1}^{H}p_{h}=0$, but we have just shown $\sum_{h=1}^{H}p_{h}>0$ whenever beliefs differ and the scoring rule is strictly concave. Therefore, the four properties cannot coexist. ∎
This theorem provides the foundation for extending impossibility results to probabilistic settings. The key insight is that proper scoring rules drive truth-revealing responses which promote using true (trained knowledge based) beliefs in the inference process.
That explains why cross-entropy loss (negative log-score) is so effective in training neural networks. It creates gradients that push each component toward truthful probability reporting. Every attention head in a transformer is implicitly playing this truth-telling game.
However, while that ensures truthfulness and relevant knowledge revelation, it doesn’t prevent collective hallucination. Each head truthfully reports its partial view, but their aggregation can still create false certainty.
7 The Emergence of Impossibility in Transformer Architectures
To demonstrate where the impossibility may emerge in transformer architectures, let us analyze a minimal yet representative example of a tiny transformer predicting the next token in the famous English pangram:
The quick brown fox jumps over the lazy dog.
We trace the inference process, inspect every mathematical operation from input embeddings through final predictions, showing exactly where the impossibility emerges.
7.1 Model Architecture and Input Representation
Our micro-transformer has the following specifications:
Parameter Value Description
$d_{m}$ 4 Model dimension
$H$ 2 Number of attention heads
$d_{k}$ 2 Query/Key dimension per head
$d_{v}$ 2 Value dimension per head
$d_{f\!f}$ 6 Feedforward hidden dimension
$|V|$ 6 Vocabulary size
$T$ 3 Sequence length
Table 1: Micro-transformer architecture parameters
For the sake of simplicity, consider the following vocabulary:
$V=\{\langle\text{PAD}\rangle,{The},{quick},{brown},{fox},{dog}\},$
including padding token (with index zero in $V$). Each token (representing word for simplicity) in that vocabulary is represented by a row in the embedding matrix
$\mathbf{E}=\begin{bmatrix}0&0&0&0&0\\ 1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{bmatrix}.$
The un-embedding matrix is defined as $\mathbf{U}=\mathbf{E}^{\top}$.
Consider the following input sequence
The quick brown
It is represented by the input matix:
$\mathbf{X}_{0}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix},$
where each row corresponds to the embedding of tokens at positions 1, 2, and 3.
The following system of equations defines output $\mathbf{o}_{t}^{(h)}$ of attention head $h$ at position $t$:
[…]
The block per-token multihead output $\mathbf{a}_{t}$ is then given by the sum:
[…]
The attention logits, i.e., the scores whose exponentials become probabilities, are given by:
$\mathbf{l}^{(h)}=\mathbf{o}^{(h)}\mathbf{U},\quad\text{so that}\quad\mathbf{L}_{a}=\sum_{h=1}^{H}\mathbf{l}^{(h)}.$ (57)
Next, we take into account the feedforward network (FFN) introducing nonlinear correction:
$\mathbf{z}=\mathbf{W}_{2}\max(0,(\mathbf{X}_{0}+\mathbf{a})\mathbf{W}_{1}+\mathbf{b}_{1})+\mathbf{b}_{2},$ (58)
generating the FFN logits:
$\mathbf{L}_{f}=\mathbf{z}\mathbf{U}.$ (59)
As a result, the residual stream is adjusted as follows:
$\mathbf{X}_{1}=\mathbf{X}_{0}+\mathbf{a}+\mathbf{z}.$ (60)
Finally, the probability distribution $\Pi$ of the next token is calculated by aggregating attention and FFN logits into:
$\mathbf{L}=\mathbf{L}_{a}+\mathbf{L}_{f}\quad\text{and normalizing into}\quad\Pi=\text{softmax}(\mathbf{L}).$ (61)
7.2 The Inference Process
We now analyze the inference process step by step to understand the moments the impossibility result may emerge. We shall keep things simple but realistic, recreating the essential steps of the inference process in transformer block.
Step 1: Query and Key
There are $H=2$ heads for which we define the query and key, or better yet, demand and supply bids in the auction of ideas. Query announces demand for information (what does this token need?), key submits supply offers (what does the previous tokens offer!). Then those bids are compared to find good matching (and clear the information market).
Let us assume the current position is $t=3$, so we attend to brown. Then, consider the following knowledge encoded in each head:
$\mathbf{W}_{Q}^{(1)}=\begin{bmatrix}0&0\\ 0&0\\ 0&1\\ 0&0\end{bmatrix}=\mathbf{W}_{K}^{(1)}\quad\mathrm{and}\quad\mathbf{W}_{V}^{(1)}=\begin{bmatrix}0&0\\ 0&0\\ 0&3\\ 0&0\end{bmatrix},$ (62)
$\mathbf{W}_{Q}^{(2)}=\begin{bmatrix}0&0\\ 0&0\\ 1&0\\ 0&0\end{bmatrix}=\mathbf{W}_{K}^{(2)}\quad\mathrm{and}\quad\mathbf{W}_{V}^{(2)}=\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\\ 0&0\end{bmatrix}.$ (63)
Matrix $\mathbf{W}_{Q}^{(h)}$ amplifies the coordinates of tokens in $\mathbf{X}$ that contribute to relevant context. As a result, we get a query vector that represents the following question:
what information is token $\mathbf{X}_{t}$ looking for?
In other words, the query vector indicates how strongly the current token demands information along the directions in $\mathbf{Q}_{t}^{(h)}$.
Then computing queries for $t=3$ with $\mathbf{X}_{3}$ yields:
$\mathbf{Q}_{3}^{(1)}=[0,1]\quad\mathrm{and}\quad\mathbf{Q}_{3}^{(1)}=[1,0].$
Head 1 (agent) demands information stored along direction $[0,1]$. Similarly, Head 2 (agent) demands information stored along direction $[1,0]$. That means the heads develop different (orthogonal) privately known and independent attention patterns, leading to beliefs about the next token that are as different as possible.
Next, we need to see what information can be provided by the input vector given the knowledge encoded in $\mathbf{W}_{K}^{(h)}$. Given that knowledge, we get from each head (agent) the key vectors that inform:
along which directions do the available tokens in $\mathbf{X}_{\leq t}$ provide information relevant for token $\mathbf{X}_{t}$.
Since we have assumed, for simplicity, that $\mathbf{W}_{Q}^{(h)}=\mathbf{W}_{K}^{(h)}$, the query and key vectors are perfectly aligned:
$\mathbf{K}_{3}^{(1)}=[0,1]\quad\mathrm{and}\quad\mathbf{K}_{3}^{(1)}=[1,0]\quad\text{and}\quad\mathbf{K}_{<3}^{(h)}=[0,0].$
Another interesting way to look at query and key is this. It is a memory lookup mechanism. Keys are address in memory, query selects addresses in memory which probably store relevant content. As we will see next, that content is provided by the value vectors.
Step 2: Attention Score and Weights
Given the demand and supply bids providing directions along which information is required and stored, in the next step the attention mechanism executes the bid matching procedure. Namely, the attention scores are computed to measure cosine similarity of bids:
$\text{scores}_{3}^{(h)}=\frac{\mathbf{Q}_{3}^{(h)}(\mathbf{K}_{\leq 3}^{(h)})^{\top}}{\sqrt{d_{k}}}.$ (64)
In our specific example, we have:
$\text{scores}_{3}^{(1)}=\frac{1}{\sqrt{2}}[0,0,1]=[0,0,0.707]\quad\text{and}\quad\text{scores}_{3}^{(2)}=\frac{1}{\sqrt{2}}[0,0,1]=[0,0,0.707].$
Applying softmax to get attention weights:
$\boldsymbol{\alpha}_{3}^{(h)}=\text{softmax}(\text{scores}_{3}^{(h)}),$ (65)
we get the following probability distributions (weights) over relevant context information:
$\boldsymbol{\alpha}_{3}^{(1)}=\boldsymbol{\alpha}_{3}^{(2)}=[0.25,0.25,0.50].$
Therefore, based on the encoded knowledge, heads attend strongly to position 3 (brown).
Step 3: Context Vector Computation
The context vector for each head is calculated as a linear combination ($\boldsymbol{\alpha}$ is a simplex) of value vectors:
$\mathbf{h}_{3}^{(h)}=(\boldsymbol{\alpha}_{3}^{(h)})^{\top}\mathbf{V}_{\leq 3}^{(h)}.$ (66)
The value vectors,
$\mathbf{V}_{\leq t}^{(h)}=\mathbf{X}_{\leq t}^{(h)}\mathbf{W}^{(h)}_{V},$ (67)
prepare context information relevant for the next token prediction (performed in next steps). That critical content is extracted from the input vector based on the trained knowledge stored in $\mathbf{W}^{(h)}_{V}$. Columns in $\mathbf{W}^{(h)}_{V}$ detect presence of relevant information in $\mathbf{X}_{\leq t}$ and scale it appropriately.
We have:
$\mathbf{V}_{\leq 3}^{(1)}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix}\begin{bmatrix}0&0\\ 0&0\\ 0&3\\ 0&0\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\\ 0&3\end{bmatrix}\ \text{and}\ \mathbf{V}_{\leq 3}^{(2)}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\end{bmatrix}\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\\ 0&0\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\end{bmatrix}.$
That means, both heads believe that the presence of brown in position 3 is relevant (row 3 in $\mathbf{W}^{(h)}_{V}$ has nonzero elements). Token brown statistically predicts important pattern extracted from training data. That information is encoded by both heads as a vectors in value space,
$\mathbf{V}_{3}^{(1)}=[0,3]\quad\text{and}\quad\mathbf{V}_{3}^{(2)}=[1.2,0].$
Given the demand-supply matching encoded in $\boldsymbol{\alpha}^{(h)}_{3}$, that yields head context vectors:
$\mathbf{h}_{3}^{(1)}=[0,1.5]\quad\text{and}\quad\mathbf{h}_{3}^{(2)}=[0.6,0].$
Step 4: Output Projection and Feedforward Exploration
Each head then projects its context vector to residual space based on the privately held information routing knowledge encoded in $\mathbf{W}_{O}^{(h)}$. That private knowledge translates (or transmits) valuable context information into embedding vectors. Namely, each head calculates:
$\mathbf{o}_{3}^{(h)}=\mathbf{h}_{3}^{(h)}\mathbf{W}_{O}^{(h)}.$ (68)
With trained matrices:
$\mathbf{W}_{O}^{(1)}=\begin{bmatrix}0&0&0&0\\ 0&0&1&0\end{bmatrix}\quad\text{and}\quad\mathbf{W}_{O}^{(2)}=\begin{bmatrix}0&0&0&1\\ 0&0&0&0\end{bmatrix},$ (69)
the attention mechanism generates ranking of evidence emphasizing relevance of the vocabulary tokes:
$\mathbf{o}_{3}^{(1)}=[0,0,0,1.5,0]\quad\text{and}\quad\mathbf{o}_{3}^{(2)}=[0,0,0,0,0.6].$
So, in the next token prediction auction of ideas, Head 1 votes for fox and Head 2 votes for dog.
The votes are added into residual dimensions to adjust the meaning of input embedding. For that, the outputs (votes) are aggregated into:
$\mathbf{a}_{3}=\sum_{h=1}^{H}\mathbf{o}_{3}^{(h)}=[0,0,0,1.5,0.6]\quad\text{and}\quad\mathbf{a}_{<3}=[0,0,0,0,0],$
and added to input vector:
$\mathbf{X}_{1}=\mathbf{X}_{0}+\mathbf{a}.$
Than completes operations of attention block.
The next step is to add contribution of FFN feedforward block (or another agent participating in the aution). We have:
$\mathbf{X}_{2}=\mathbf{X}_{1}+\mathbf{W}_{2}\max(0,\mathbf{X}_{1}\mathbf{W}_{1}+\mathbf{b}_{1})+\mathbf{b}_{2}.$
That FFN operation explores context locally and introduces additional innovations in meaning, shaping the next token prediction result. It can amplify or counteract the votes injected by attention. Notice that adjustments are applied at every position separately, so that ideas provided by attention bids are interpreted and rewritten by the FFN agent.
7.2.1 Step 5: Logits and Prediction
Finally, the unembedding matrix $\mathbf{U}$ maps the multi-headed attention and FFN outputs from model space to vocabulary in order to calculate probability distribution for the next token prediction.
Each head’s contribution to the final logits is:
$\mathbf{l}^{(1)}=[0,0,0,0,1.5,0]\quad\text{and}\quad\mathbf{l}^{(2)}=[0,0,0,0,0,0.6].$
The total attention logits are the sum of individual head logits:
$\mathbf{L}_{a}=\sum_{h=1}^{H}\mathbf{l}^{(h)}=[0,0,0,0,1.5,0.6].$
That sum is combined with FFN logits $\mathbf{L}_{f}$ that introduce some creative noise. As a result, we obtain:
$\mathbf{L}=\mathbf{L}_{a}+\mathbf{L}_{f}=[0,-0.03,0.21,0.13,1.31,0.52].$
That vector translates into the final probability distribution:
$\Pi=\text{softmax}(\mathbf{L})=[0.1,0.1,0.13,0.12,\mathbf{0.38},0.17].$
The winner of the auction of ideas is selected based on $\Pi$. The inference process generates and collects competing beliefs about relevant context and next token prediction, all based on the knowledge stored in the attention heads and FFN layer. And so, we see that when that knowledge is used, the most probable next token (word) reachable in the vocabulary is
$\text{argmax}\ \Pi=\textit{fox}.$
7.3 The Impossibility Theorem for Transformer
With the deep insight into the transformer inference process, we are ready to show the impossibility in that special and important setting.
Theorem 5 (Impossibility for Transformers).
In transformer LLMs with non-redundant attention heads, it is impossible to implement inference process in which each head outputs its true conditional distribution while conserving semantic information and minimize aggregate loss with logit summation.
Proof.
Let attention heads $h=1,\ldots,H$ output logits $\mathbf{l}^{(h)}\in\mathbb{R}^{|\mathcal{V}|}$. Then, individual head distributions are given by:
$\pi^{(h)}_{y}=\frac{\exp\bigl{(}l^{(h)}_{y}\bigr{)}}{Z_{h}}\ \text{with}\ Z_{h}=\sum_{\bar{y}}\exp\bigl{(}l^{(h)}_{\bar{y}}\bigr{)}.$ (70)
Similarly, the aggregate distribution from summed logits is calculated as:
$\Pi_{y}=\frac{\exp(L_{y})}{Z}\ \text{with}\ Z=\sum_{\bar{y}}\exp(L_{\bar{y}})\ \text{and}\ L_{y}=\sum_{h=1}^{H}l^{(h)}_{y}.$ (71)
For any observed (realized) token $y^{*}$, the head losses and aggregate loss are:
$p_{h}=-\log\pi^{(h)}_{y^{*}}=-l^{(h)}_{y^{*}}+\log Z_{h}\quad\text{and}\quad p_{H+1}=-\log\Pi_{y^{*}}=-L_{y^{*}}+\log Z.$ (72)
But that implies,
$\sum_{h=1}^{H}p_{h}=p_{H+1}+\gamma,\ \text{where }\Gamma=\sum_{h=1}^{H}\log Z_{h}-\log Z.$ (73)
If heads produce non-proportional logit vectors, then $\Gamma>0$ by convexity of log-sum-exp aggregation.
Therefore,
$\sum_{h=1}^{H}p_{h}>0,$ (74)
which violates semantic information conservation principle.
This final theorem bridges earlier impossibility results to concrete transformer architectures. It shows that concavity of scoring functions is the mathematical source of hallucination. The Jensen gap $\Gamma$ becomes a measurable quantity showing that probability aggregation violates conservation and any sufficiently capable transformer-based LLM must violate conservation of per-token information on non-trivial queries.
Let us note that the aggregate distribution cannot create Shannon information ex nihilo. Instead it reorganizes existing information to be more useful for the specific task of predicting $y^{*}$. Given fixed head logits, the Shannon entropy of a function (convex aggregate) of those logits is bounded from above by the the Shannon entropy of the fixed logits. This way information becomes more accessible, not more existent, as postulated by the reasoning with emergence operator and Theorem 1. The Jensen’s gap measures point-wise excess confidence of the predicted outcome, possibly making LLMs appear to know more (when in fact they do not).
8 Summary
This paper establishes a fundamental mathematical impossibility at the heart of large language model inference: no LLM can simultaneously achieve perfect hallucination control while maintaining semantic coherence and optimal knowledge utilization. That result demonstrates that four essential properties of ideal language model behavior are mutually incompatible, creating inevitable trade-offs that no architectural innovation or training methodology (aligned with the impossibility theorem assumptions) can eliminate.
For any query space $\mathcal{Q}$ containing non-trivial queries requiring integration of distributed knowledge, no language model inference mechanism can simultaneously:
  • represent encoded knowledge accurately,
  • preserve total information content during inference,
  • reveal all accessible relevant information,
  • minimize hallucination cost within available knowledge.
This impossibility emerges from the fundamental mathematical structure of information aggregation, not from any particular architectural choice.
8.1 Mathematical Foundation and Proof Strategy
The proof operates across three complementary mathematical frameworks, each capturing different aspects of LLM inference.
Auction-Theoretic Framework
We model LLM inference as an auction of ideas where neural components (attention heads, circuits, activation patterns) compete to contribute knowledge to response generation. Under the assumption of independent private knowledge profiles, we apply the Green-Laffont theorem to prove impossibility.
Probabilistic Framework
We extend the impossibility to any system where neural components output probability distributions by leveraging Savage’s theory of proper scoring rules. For convex aggregation mechanisms with strictly concave scoring functions, Jensen’s inequality creates an inevitable information aggregation gap. The Jensen’s gap represents excess confidence generated by probabilistic aggregation, violating information conservation whenever beliefs differ across components.
Transformer Architecture
We demonstrate how the impossibility manifests in actual transformer implementations through log-sum-exp convexity. The (Jensen’s) convexity gap $\Gamma$ provides a directly measurable quantity in transformer training, showing that information conservation violations are not merely theoretical but empirically observable.
8.2 Key Mathematical Contributions
To establish the impossibility result, we have designed a mathematical framework providing both formal and philosophical ground for the study. The Semantic Information Measure is a context-dependent information metric that captures bounded reasoning capabilities. The Emergence Operator models how reasoning makes hidden knowledge accessible within computational constraints. And Auction-Theoretic Modeling allows us to identify logical limitations of the inference process.
The impossibility results apply across multiple mathematical models of knowledge representations, training goals, and neural architectures. This universality suggests the trade-offs we have identified are fundamental features of distributed intelligence systems, not artifacts of particular modeling choices.
8.3 Implications for AI Development
Despite its name, the impossibility theorem provides constructive guidance for LLM design and deployment. Rather than pursuing impossible simultaneous optimization, we may consider balancing competing objectives. Namely, prioritize truthfulness and information conservation in safety-critical applications, allow controlled information conservation violations to enable innovative responses in creative applications, develop adaptive mechanisms that adjust trade-offs based on query context in general-purpose systems.
The impossibility result may guide LLM evaluation. Impossibility-aware metrics can measure trade-off quality rather than absolute performance, alignment strategies can negotiate trade-offs with human preferences rather than assuming universal optimality.
8.4 Significance for AI Theory
Just as Goedel’s incompleteness theorems revealed fundamental limits in formal systems, the impossibility theorems in this paper identify inherent constraints in knowledge integration and reasoning. Hallucination, or imagination, is a price of reasoning, much like incompleteness is a price of consistency.
The Jensen’s gap controlling the impossibility strength may provide insight into the relationship between aggregation and information content. The apparent creativity in large language models may be better understood as necessary violations of information conservation rather than genuine knowledge generation.
8.5 Limitations and Future Directions
The impossibility results apply specifically to non-trivial queries requiring knowledge integration from multiple sources, finite computational budgets that create bounded reasoning constraints, and distributed knowledge representations where information is spread across multiple components. Future work should investigate whether alternative knowledge architectures might circumvent some aspects of the impossibility.
8.6 Conclusion
The fundamental impossibility of perfect hallucination control represents both a limitation and a insight for the field of artificial intelligence. By understanding that trade-offs are mathematically inevitable, we can move beyond attempts at simultaneous optimization toward principled management of inherent tensions in intelligent systems. Formal definitions of the impossibility building blocks open new avenues for research at the intersection of mathematical analysis, engineering design, and philosophical inquiry into the nature of machine intelligence.
When we observe LLMs appearing to generate novel insights, they are either $(i)$ revealing information already implicit in their training (making knowledge accessible) or $(ii)$ violating conservation (hallucinating). This explains why LLMs can seem both remarkably creative and fundamentally limited. They excel at revealing hidden connections but cannot easily transcend information content of their training dataset.
This impossibility is not a bug but a feature of distributed intelligence. That price of knowledge integration in reasoning is worth paying.