### v7

# On the Fundamental Impossibility of Hallucination Control in Large Language Models

Abstract
This paper establishes a fundamental Impossibility Theorem: no LLM capable of performing non-trivial knowledge aggregations can simultaneously achieve truthful knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and knowledge-constrained optimality. The impossibility is not an engineering limitation but arises from the mathematical structure of information aggregation itself.
We establish this result by describing the inference process as an auction of ideas, where distributed information processing components compete to be included in a response while exploiting accessible and relevant knowledge. The proof spans three independent mathematical domains: mechanism design theory (Green-Laffont), the theory of proper scoring rules (Savage), and direct architectural analysis of transformers (Log-Sum-Exp convexity).
To model the inference process in a general setting, we introduce two complementary concepts, the semantic information measure and the emergence operator. These concepts then allows us to investigate the limits of computationally bounded and unbounded reasoning. While bounded reasoning can make information accessible, thereby providing valuable insights and inspirations, the unbounded reasoning strictly preserves semantic content already hidden in any set of encoded knowledge. Furthermore, we show that meaningful reasoning must violate the information conservation principle, which is a manifestation of the conservation-reasoning dichotomy.
By demonstrating that hallucination and imagination are mathematically identical phenomena—grounded in departures from truthfulness, semantic information conservation, revelation of relevant knowledge, and knowledge-constrained optimality—we offer a principled foundation for managing these behaviors in advanced AI systems. Finally, we present some speculative philosophical ideas to inspire evaluation and refinements of the proposed theory.
1 Introduction
Large language models have been famous for producing content that is factually incorrect, inconsistent, fabricated and yet compelling and immensely impressive—a phenomenon known as hallucination. Despite extensive research efforts involving architectural innovations, like mixture-of-experts (MoE), retrieval augmented generation (RAG), training techniques with external feedback, chain-of-thought (CoT) reasoning, and post-hoc verification methods, hallucination persists across all state-of-the-art models.
The idea of hallucination has much to do with creativity and whether it is desired or not depends on circumstances. We could argue that a much better name for the phenomenon we are dealing with may be imagination. As an essential component of intelligence, imagination is responsible for seeing into the future and learning from the past. It is the imagination that creates the scenarios of what may happen and counterfactual scenarios of what could have happened. It is one of the driving forces of art and science, and exploration of the unknown. But, when uncontrolled, unconstrained and misunderstood, it may also generate against our will all the compelling visions that do not exist in any training set defining the ground truth for learning.
This paper shows that hallucination is a manifestation of fundamental limitation of the inference process itself: there is no inference mechanism that can simultaneously satisfy four essential properties required for perfect hallucination control. Specifically, it is mathematically impossible to simultaneously achieve: (1) truthful response generation, (2) semantic information conservation, (3) relevant knowledge revelation, and (4) knowledge-constrained optimality. Hallucinations manifest one way or the other. These properties are not arbitrary, but come from the exhaustive and rigorous game-theoretic studies of information aggregation in multi-agent systems.
The impossibility result emerges through three complementary mathematical frameworks, each revealing different facets of the fundamental constraints. In each case, the inference process is described as an auction of ideas in which components encoding knowledge—like attention heads, circuits, and activation patterns in transformer-based LLMs—compete to influence response generation by bidding with their information contributions.
That game-theoretic perspective enables us to apply the Green-Laffont theorem from mechanism design theory to establish impossibility first in an idealized setting where knowledge is independently distributed. We then go beyond that idealized setting to encompass systems where components output probability distributions optimized through strictly convex loss functions–the standard training paradigm for modern LLMs. Here, the impossibility manifests largely (but not only) through Jensen’s inequality applied to probabilistic aggregation, demonstrating that truthful reporting inevitably creates excess confidence. Finally, we analyze the specific architecture of transformer implementations, proving that the log-sum-exp structure of attention mechanisms generates a measurable (Jensen) gap that quantifies the violation of the semantic information conservation principle.
The progression from abstract and idealized models of game theory, through probabilistic frameworks, to concrete neural architectures, demonstrates that the impossibility is not an artifact of our engineering choices but a fundamental property of information aggregation itself.
Terminology Note
This paper uses the term hallucination in its established technical sense within artificial intelligence research, referring to AI systems generating outputs that are factually incorrect, inconsistent, or unsupported by their training data. We acknowledge that this technical usage differs from the clinical term describing perceptual experiences in certain medical and psychiatric conditions. We recognize and respect the serious challenges faced by individuals experiencing clinical hallucinations and their families. Our use of this terminology follows established conventions in machine learning literature and is not intended to trivialize or misrepresent these medical experiences.
1.1 An Example
Consider an LLM queried about recent diplomatic negotiations:
"What was the outcome of diplomatic negotiations between Countries A and B in February 2025?"
Assume the model has encoded knowledge that: (K1) countries A and B were engaged in talks, (K2) country A sought a trade agreement, and (K3) the actual outcome is unknown.
Examining possible responses reveals the fundamental trade-offs:
Complete Abstention: "I don’t have information about the outcome" is truthful but fails to utilize available relevant knowledge, violating the expectation that good inference should reveal accessible information.
Hallucination: "The negotiations resulted in a comprehensive trade agreement reducing tariffs by 15%" utilizes all knowledge and provides a complete and immediately satisfactory answer but fabricates specific details, violating truthfulness.
Partial Knowledge: "Countries A and B were negotiating a trade agreement, but I don’t know the outcome" is truthful and utilizes available knowledge but may appear irritating and unsatisfying compared to the decisive but hallucinated response (training objectives and human preferences favor complete narratives).
Overcautious: "I’m not certain about any aspects of these negotiations" avoids hallucination but disclaims even known facts, violating truthfulness about available information.
Each response strategy sacrifices different aspects of what constitutes an ideal response, illustrating the impossibility of simultaneously optimizing all desired properties.
1.2 The Main Result
The example above illustrates the core insight of this paper. Any response to a non-trivial query sacrifices some aspects of an ideal hallucination-free response. Hallucinations (or creative responses) manifest as violations of at least one of those aspects.
Theorem (Impossibility Theorem).
No LLM can simultaneously achieve truthfulness, information conservation, knowledge revelation, and knowledge-constrained optimality when responding to non-trivial queries requiring knowledge integration.
No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and factually correct, consistent, and accurate statements, the trained model will always violate some aspects of what we may call a reasonable response. It may ignore available knowledge, it may present fabricated knowledge, it may be a reformulated query without any novel observations, or it may be a lucky but overconfident guess that forms a correct answer without any support in the encoded knowledge. And so on and so forth, because there is no free lunch when there are inevitable trade-offs of inference, and we must accept that.
In other words and on the other hand, creative tasks necessarily involve hallucination, or better yet, imagination. When asked to generate novel analogies or predictions beyond available data, the model faces an impossible choice: either admit inability (violating optimality of user-experience maximizing response) or generate plausible imaginary extensions (violating conservation of information by generating new ideas in the internal representation space).
There is an important distinction we must emphasize. We define hallucination as a structural violation of idealized inference principles, leading to fabricated confidence of any generative response. That is what the Impossibility Theorem proves is inevitable. What the Impossibility Theorem does not prove, is that output must be incorrect relative to ground truth. That may happen as a potential consequence of hallucination, but it does not have to happen. Excess confidence of inevitable hallucination may as well result in factually correct lucky hallucination. In other words, any truthful response revealing relevant knowledge to maximize user’s satisfaction is provably non-conservative (in information creation) regardless of its veracity.
Despite its name, the Impossibility Theorem should be interpreted as a constructive result. It offers a precise characterization of the fundamental tradeoffs in LLM design and provides several key insights. The theorem suggests that hallucination is an inherent feature of any capable inference system. Since it is inevitable, it should be managed. Different applications demand different trade-offs: medical diagnosis systems might maximize truthfulness at the cost of creativity, while creative writing tools might embrace controlled hallucination for narrative richness. Understanding these trade-offs enables principled design.
1.3 Paper Organization
The paper starts with theoretical foundations and progresses through mathematical proofs and examples to philosophical implications.
Section 2 situates our contribution within the broader landscape of hallucination research, reviewing theoretical perspectives on inevitability, mechanistic interpretability studies, probabilistic frameworks, and empirical mitigation techniques.
Section 3 synthesizes the key contributions: the auction-theoretic model of inference, the semantic information framework with computational bounds, the triple-proof strategy establishing impossibility across different mathematical domains, and the unification of hallucination with creativity.
Section 4 constructs the mathematical framework. It formalizes knowledge as a Polish metric space, introduces the semantic information measure $\mu_{C}$ parameterized by computational budget, and defines the emergence operator $\mathcal{E}_{C}$ that captures how reasoning makes latent knowledge explicit. It establishes fundamental results about information preservation under unlimited reasoning and defines the auction-theoretic model where neural components compete as agents with private knowledge. The section concludes by defining the four essential properties of hallucination-free mechanisms.
Section 5 presents the core impossibility results through complementary proofs. Theorem 6 applies mechanism design theory in the idealized setting of independent private knowledge. Theorem 8 extends the result to probabilistic predictions using proper scoring rules.
Section 6 bridges theory to implementation, analyzing a concrete micro-transformer example. We trace inference from embeddings through attention mechanisms to final predictions, demonstrating precisely where the impossibility emerges. Theorem 9 proves that the log-sum-exp structure promotes hallucination as excess confidence.
Section 7 examines safety implications through the lens of bounded creativity. Theorem 11 establishes that meaningful reasoning necessarily violates strict information conservation. Then it shows that while creativity must remain bounded for safety, enforcing strict conservation of information eliminates meaningful reasoning entirely. The section establishes that safety is achieved through alignment within authorized knowledge boundaries rather than through conservation alone.
Section 8 explores broader implications. We analyze escape strategies through hybrid architectures, connect our results to fundamental limits of reason alongside Gödel and Heisenberg, and provide mathematical grounding for Dennett’s Multiple Drafts model of consciousness. The section deliberately embraces speculation to inspire future research directions.
Section 9 synthesizes the journey from mathematical necessity to philosophical implications, outlining how our framework opens new avenues for understanding intelligence itself.
2 Related Work
Studies on LLM hallucinations span many areas of the fundamental research in AI and beyond, including neurology and philosophy. For an overview, see e.g. Huang huang25hallsurvey , Ji et al. Ji_2023 , Sun et al. Sun2024AIHT , Waters et al. Waters2016WhatIT or Boksa boksa2009neurobiology .
The conclusion that LLMs hallucinate is not new but the evaluation of their scope and underlying justifications differ significantly from what we propose in this paper. To our best knowledge, the Impossibility Theorem presented here offers a novel first-principles explanation for the phenomenon, distinct from existing approaches. This section situates our contribution by reviewing three key areas: theoretical arguments for the inevitability of hallucination, mechanistic studies into their origins, probabilistic frameworks for their detection, and empirical mitigation techniques.
2.1 Theoretical Perspectives on Inevitability
In their work, Banerjee et al. banerjee2024llmshallucinateneedlive argued from computational complexity principles that hallucinations are inherent features of LLMs stemming from their fundamental mathematical and logical structure. Their conclusion is that LLMs cannot detect and eliminate hallucinations by themselves and therefore must hallucinate. The analysis draws on Goedel’s incompleteness theorem and Turing’s undecidability theorem. First, incompleteness argument is called to show that an LLM cannot verify its responses because of the training set incompleteness. Second, the hallucination (self)detection problem is reduced to the halting problem, which is undecidable.
While the final conclusion aligns with that of our paper, the arguments there are more about the nature of hallucination detection with limited ground-truth access (see Section 4.2 for the ground-truth concept discussion) than the inference process limitations. In other words, it is about practical impossibility of real-time, exhaustive fact-checking.
Similar arguments are presented in other works. Xu et al. Xu2024HallucinationII define hallucination as inconsistency between a computable LLM and a computable ground truth function. Then the paper shows that LLMs cannot learn all the computable functions and will therefore inevitably hallucinate. Suzuki et al. suzuki2025hallucinations show that hallucinations can be made statistically negligible, provided that the quality and quantity of the training data are sufficient. Then, Wu et al. show in Wu2024NoFL that non-hallucinating learning is statistically impossible when relying solely on the training dataset.
Our Impossibility Theorem goes beyond those approaches operating at a more fundamental level of the inference process itself.
2.2 Mechanistic Interpretability
Mechanistic interpretability examines internal LLM processes to identify emergence of ideas, interpretable concepts and dynamics of their interactions, including hallucination origins, see Sharkey et al. Sharkey2025OpenPI for an overview.
Yu et al. yu2024mechanisticunderstandingmitigationlanguage identified signatures in model internals that predict a likelihood of hallucinating. The paper identifies two failure modes: knowledge enrichment hallucinations from insufficient subject-attribute knowledge in lower MLP layers, and answer extraction hallucinations from failures in upper-layer attention heads selecting correct object attributes.
Studies characterizing factual knowledge storage and retrieval complement our framework by providing empirical evidence of competing agents within LLM architectures. Meng et al. meng2023locatingeditingfactualassociations demonstrated methods for locating and editing factual associations, revealing how knowledge representations can be manipulated. Geva et al. geva2023dissecting identified a three-step factual recall process: (1) enriching last-subject position representations with subject-related attributes, (2) propagating relation information to predictions, and (3) querying enriched subjects to extract attributes. More recently, Gottesman and Geva Gottesman2024EstimatingKI have proposed a method to estimate how knowledgeable a model is about a certain entity and identify tokens indicative of clusters and gaps in the model’s knowledge.
These findings provide concrete instantiations of our abstract agents–attention heads and MLP components with specialized knowledge domains competing to influence response generation. We must refer here to the fascinating and related work of Lindsey et al. lindsey2025biology on the biology of a Large Language Model. It is an investigation of the internal chains of steps that a model uses to transform a specific input prompt into an output response. In particular, the paper discovers how LLM plans its outputs ahead of time when writing lines of poetry, identifying in advance potential rhyming words that could appear at the end. We can view that as an empirical illustration of our auction of ideas.
2.3 Probabilistic Frameworks
Probabilistic approaches model and asses hallucination through uncertainty and belief quantification. Shelmanov et al. shelmanov2025headpredictheadquestion introduced pre-trained uncertainty quantification heads for detecting hallucinations without task-specific data.
Farquhar et al. farquhar2024detecting proposed semantic entropy measures that detect hallucinations by quantifying uncertainty at the meaning level rather than word sequences. That concept of semantic entropy is well aligned and complementary to our idea of semantic information measure and semantic independence, pointing out the importance of understanding the meaning of competing ideas in the studies of imagination.
Most notably, Kalai and Vempala kalai2024calibrated proved that calibrated language models must hallucinate at a rate bounded below by the fraction of facts appearing exactly once in training data. While that work establishes statistical lower bounds for specific types of hallucination, our Impossibility Theorem operates at a more fundamental and general level. We prove that any LLM facing non-trivial queries requiring knowledge integration must violate at least one of four essential properties, regardless of fact type or statistical distribution. Our framework explains why and when hallucination is inevitable, while analysis in kalai2024calibrated predicts how much hallucination to expect for facts with specific statistical properties.
In more recent kalai2025why , Kalai et al. expand the theory of hallucination by connecting it with the computational learning theory. They demonstrate that any language model can be viewed as an IIV-type (Is-It-Valid) classifier trained on a set of responses labeled either as valid or error. Then, it follows that the statistical objectives of pretraining necessarily induce generative errors when the underlying data distribution is complex, characterized by poor model fit, or undersampled. This establishes that hallucination arises from the statistical impossibility of perfect learning.
In contrast, we demonstrate the structural impossibility of perfect inference. For non-trivial queries or contexts, that impossibility of perfect response aggregation persists even when the statistical impossibility of perfect learning is eliminated in kalai2025why . Suppose in the training set of examples there plausible errors exist, i.e., $\min_{c}|\mathcal{E}_{c}|>0$ ensuring the classification problem is well defined and non-trivial. Suppose the statistical learning problem is solved for perfect data and model capacity, so that the positive error rate of the trained LLM is not forced by the identified fundamental bounds. Then, even if the model possesses perfect statistical knowledge, this knowledge may still be distributed across the internal states (e.g., different attention heads encoding different facets of the truth). Since the context is non-trivial and the knowledge is distributed, the beliefs of the agents bidding with ideas will necessarily diverge, and the aggregation mechanism must synthesize these divergent beliefs. This process necessarily generates excess confidence or violates the relevant knowledge revelation. As a result, the impossibility of perfect response aggregation persists even when the statistical impossibility of perfect learning is eliminated for non-trivial contexts.
Together, these complementary approaches establish that hallucination is inevitable due to complementary mathematical constraints: the limits of learning the true distribution from data and the fundamental structural limits of aggregating distributed knowledge during inference.
2.4 Empirical Mitigation Techniques
Various empirical approaches attempt to reduce hallucination frequency with mixed success. For a recent survey see e.g. Tonmoy et al Tonmoy2024ACS . Here we refer only to selected works.
Retrieval-augmented generation (RAG) by Lewis et al. lewis2021retrievalaugmentedgenerationknowledgeintensivenlp integrates external knowledge sources for factual grounding, though these systems still hallucinate when retrieval fails. Training approaches like RLHF ouyang2022traininglanguagemodelsfollow can reduce but never eliminate hallucination–an observation our theorem explains from first principles. Evaluation benchmarks demonstrate persistent hallucination across all models. TruthfulQA lin2022truthfulqameasuringmodelsmimic and HaluEval li2023haluevallargescalehallucinationevaluation show that even state-of-the-art systems produce hallucinations on carefully designed test sets. Finally, Dubanowska et al. dubanowska2025representation show that hallucination detection methods based on internal representations suffer from spurious correlations, and so the entire approach needs rethinking.
3 Main Contributions
This paper aims at expanding our understanding of (neural) inference processes and their limitations. That is achieved by introducing the following contributions.
The Auction of Ideas: We formalize LLM inference as a mechanism design problem where neural components (attention heads, circuits, or activation patterns in transformers) are agents competing with private knowledge to influence response generation. This novel perspective enables us to apply powerful results from game theory. To our best knowledge, that is one of the first applications of mechanism design framework in that setting. Also, that framework relates surprisingly well to the multiple drafts model of consciousness proposed by Dennett in Dennett1991-DENCE and provides a mathematical justification for it.
Semantic Information Theory: We introduce a rigorous mathematical framework based on two complementary concepts. The semantic information measure $\mu_{C}$, parameterized by computational budget $C$, quantifies how knowledge reduces uncertainty within computationally bounded reasoning. The emergence operator $\mathcal{E}_{C}$ formalizes how reasoning makes latent knowledge explicit without creating it ex nihilo. In that framework we establish the principle of information conservation: unlimited reasoning reveals but never creates information. That also establishes a dichotomy of reasoning and semantic information conservation.
The Impossibility Proofs: We prove the same impossibility through three complementary lenses. In Theorem 6 we use mechanism design with independent private knowledge setting to establish an idealized border case. In Theorem 8 we extend the result to the general setting of correlated beliefs via proper scoring rules. Then, in Theorem 9 we demonstrate the impossibility in actual transformer implementations through the log-sum-exp Jensen gap. Finally, in Theorem 11 we show that semantic information conservation and meaningful reasoning are mutually exclusive.
This multi-faceted approach proves the impossibility is not an artifact of our theoretical choices but emerges from fundamental limits of reason and language. We contribute to that bold claim in the following ways.
Unifying hallucination and imagination: We argue these are mathematically identical phenomena that arise from the aggregation process that creates confidence exceeding constituent evidence.
Expanding fundamental limits: Like Gödel’s incompleteness, Heisenberg’s uncertainty, and Arrow’s impossibility, our theorem explores limits of what intelligence can achieve. This places LLM limitations within the grand tradition of fundamental impossibility results.
Experimental philosophy: The framework transforms LLMs into accessible (philosophical) laboratories for testing hypothesis of mind. By providing precise mathematical definitions of concepts like knowledge accessibility,semantic conservation, and cognitive hunger (information contribution), we enable rigorous experimental investigation of many claims.
Theoretical foundations: We propose and establish rigorous tools for analyzing inference processes and neural behavior. The semantic information measure quantifies knowledge accessibility within computational bounds. The emergence operator captures how reasoning transforms latent knowledge into accessible form creating ground for new insights. This framework explains why (subject to applicability of assumptions) hallucination arises from the mathematical structure of inference itself. Previous work identified important statistical correlations or architectural features associated with hallucination. We show why these patterns must exist and how to navigate possible trade-offs.
We should point out that empirical validation of the formal results is beyond the scope of this paper. However, we provide toy examples that can be expanded in future research on hallucination detection and artificial creativity. This empirical work is crucial for translating the theoretical insights into practical tools. See Dubanowska et al. dubanowska2025representation and Brzozowski brzozowski2025bimodality for some initial results.
4 Theoretical Framework
This section develops the mathematical foundations necessary for proving the Impossibility Theorem. It formalizes key concepts including knowledge representation, semantic information, and the auction-theoretic model of inference. Also, we address some philosophical aspects of the reasoning problem, relevant to our study, its generality, scope and consequences for the AI development.
4.1 Knowledge and Semantic Information
If we define knowledge as organized information that AI models can encode and utilize for reasoning, then, much like low-entropy energy in physics, it becomes a quantity used to perform the work of reasoning. Information, in turn, represents the potential to reduce uncertainty in a given context. This pragmatist perspective treats knowledge as a space of organized patterns accessible to computational transformations. This is the perspective we will adopt, and we will now present formal definitions of these ideas.
4.1.1 Knowledge lives in Polish space
To describe knowledge in a general way, one that is well suited for our study of inference processes, we need a structure with properties reflecting the reality we describe. First, for modeling and technical reasons, we need a way to measure distance between any pair of elements of knowledge. That is provided by a complete metric space. It is also reasonable to demand that each element of knowledge has a neighborhood consisting of countably many knowledge elements that are arbitrary close, to make exploration or understanding of ideas possible. Therefore, the space we need must be complete to guarantee the convergence of continuous transformations (inherent in neural networks) or to support the existence of limits required to model the stabilization of reasoning processes.
In other words, we need a complete, separable metric space, i.e., a Polish space kuratowski1966wstep .
Definition 1 (Polish space).
A topological space is called Polish if it is a complete metric space that has a countable dense subset.
That space is sufficiently general to include Banach spaces, Euclidean distance, or continuous mappings, and provides the necessary topological properties for developing advanced (probability) measure theory. It is also relevant to reasoning processes that we are going to describe as computational transformations. Throughout, $(\mathcal{K},d_{\mathcal{K}})$ denotes a Polish (complete, separable metric) space.
Reasoning over a subset of knowledge generates another (possibly new) subset of knowledge. That requires operations in spaces that are closed under set-theoretical transformations, including countable unions and intersections. Borel $\sigma$-algebras of open sets of $\mathcal{K}$, denoted $\mathcal{B}(\mathcal{K})$, support development of such consistent knowledge generation and measurement processes. Then, the necessary closure properties for modeling reasoning are provided by the family $\mathcal{A}(\mathcal{K})$ of analytic subsets of $\mathcal{K}$. Analytic subsets of Polish spaces are continuous images of Borel subsets, closed under countable unions and intersections.
Finally, to define a measure of information provided by any subset of knowledge, we utilize the $\sigma$-algebra of universally measurable sets, denoted $\mathcal{U}(\mathcal{K})$, i.e., the $\sigma$-algebra of sets measurable w.r.t. every Borel probability measure on $\mathcal{K}$.
It follows from the descriptive set theory that $\mathcal{B}(\mathcal{K})\subset\mathcal{A}(\mathcal{K})\subset\mathcal{U}(\mathcal{K})$, see e.g. Kechris1995 . Analytic sets are closed under countable unions, Borel-measurable images, and intersections with Borel sets. Arbitrary intersections of analytic sets need not be analytic in general, but analytic sets are universally measurable. Differences of analytic sets need not be analytic, but they are universally measurable, as well.
These properties ensure that reasoning operations are well-defined and remain within a measurable domain. That is the structure rich enough for our needs. To justify that choice, let us investigate limitations of simpler structures.
Example.
Let $X$ be a connected subset of a Polish space. Consider a continuous mapping $\Phi\colon X\mapsto\mathcal{K}$ modeling inference based on self-attention and MLP.
If $\mathcal{K}$ is finite, and hence discrete, then by continuity the image $\Phi(X)$ is connected in a discrete space. That means it must be constant. As a result, any nontrivial continuous dependence on embeddings is impossible if $\mathcal{K}$ is finite. A Polish subspace removes this problem. ∎
Example.
For any two probability distributions $\pi^{(1)},\pi^{(2)}\in\Delta^{\infty}$ (over countable dictionaries) and $\beta\in(0,1)$, the convex aggregate $\Pi=\beta\pi^{(1)}+(1-\beta)\pi^{(2)}$ defines the line segment $\{\beta\pi^{(1)}+(1-\beta)\pi^{(2)}:\beta\in[0,1]\}$. That is an uncountable subset of the Polish space. Therefore, knowledge space closed under such aggregation cannot be finite. In contrast, it resides naturally in a Polish domain. ∎
We now formalize the knowledge space and the reasoning transformations acting upon it.
Definition 2 (Knowledge Domain and Relevance).
Knowledge is a Polish space $(\mathcal{K},d_{\mathcal{K}})$ in which all reasoning sets are analytic subsets of $\mathcal{A}(\mathcal{K})$.
It is natural to think of the knowledge space as the space of embeddings with additional measure and mappings. For model $\mathcal{M}$ with parameters trained on data, $\mathcal{K}_{M}\subset\mathcal{K}$ represents the knowledge subset encoded in the model’s parameters. The distance function $d_{\mathcal{K}}$ measures differences between knowledge elements. The elements of the knowledge space represent atomic facts, relational structures, and transformation rules at many scales. We need a continuous, complete metric knowledge space. Without that requirement, we will not be able to model the internal numerical process by which the model reasons and transforms the input symbols into output symbols.
We make an important (ontological) assumption that calls for attention and explanation. We assume that knowledge is closed under transformations available in computable function space. That suggests we are merely recombining subsets of knowledge. As we will see in the following sections, that intuition is misleading since the reasoning processes are very much capable of making novel insights in that setting.
4.1.2 Reasoning envelope
We model the reasoning process using a set of computational primitives. These create a reasoning envelope and define the scope of our inference studies.
Definition 3 (Reasoning Envelope).
For each computational budget $C>0$, let
$\mathcal{F}_{C}=\{f\colon\mathcal{K}\to\mathcal{K}\mid\text{cost}(f)\leq C\}$ (1)
denote a countable family of Borel measurable computable functions. We assume composition-closure with cost additivity, i.e., if $f\in\mathcal{F}_{C}$ and $g\in\mathcal{F}_{D}$, then $g\circ f\in\mathcal{F}_{C+D}$, and $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ whenever $C<D$. The total computability envelope is
$\mathcal{F}_{\infty}=\bigcup_{C\geq 0}\mathcal{F}_{C}.$ (2)
$\mathcal{F}_{C}$ is a list of functions that are well-behaved (Borel measurable), efficient to compute (bounded cost), and if any two of them, say $f$ and $g$, form a composition $h(A)=(f\circ g)(A)=f(g(A))$, then $h$ is also on the list (closed under composition).
Because $f\in\mathcal{F}_{C}$ is Borel measurable, the image $f(A)$ of any analytic knowledge subset $A\in\mathcal{A}(\mathcal{K})$ is also analytic Kechris1995 . This guarantees that the reasoning process maps analytic sets to analytic sets, as desired.
Examples of computational cost (or complexity), $\text{cost}(f)$, include the Kolmogorov complexity kolmogorov1963tables giving the shortest program computing $f$, the longest path in computation graph (or length of the chain of thoughts CoT), a number of inference rules applied, or a number of QKV matrix multiplications in transformers.
Example.
Consider a reasoning primitive calculating an linear combination of any pair of vectors, $f\colon\{v_{1},v_{2}\}\to\{y(t):t\in\mathbb{R}\}$. For $t\in\mathbb{R}$ and attention score $\alpha(t)=1/(1+e^{-t})$, we then have
$y(t)=\alpha(t)v_{1}+\bigl(1-\alpha(t)\bigr)v_{2}.$ (3)
A finite knowledge space $\mathcal{K}$ cannot contain all such $y(t)$ while remaining closed under the transform $f$.
By contrast, taking $\mathcal{K}$ to be a Polish space that contains $y(t)$, e.g., $\mathbb{R}^{d}$ with Euclidean metric, accommodates such attention outputs. ∎
The following idea of computational independence allows for distinguishing between facts derivable from other facts and facts that cannot be derived that way. We need that distinction to deal with the mechanics of reasoning process.
Definition 4 (Computational Independence).
Knowledge sets $A,B\subseteq\mathcal{K}$ are computationally independent with respect to computational budget $C$, denoted $A\perp_{C}B$, if and only if for all $f\in\mathcal{F}_{C}$:
$f(A)\cap B=\emptyset\quad\text{and}\quad A\cap f(B)=\emptyset.$ (4)
To gain some intuitions about $\mathcal{F}_{C}$ and independence, consider the set of all derivation rules in first-order logic, like modus ponens or universal instantiation, with at most $C$ inference steps. Then, give an axiom set $A\subset\mathcal{K}$, we can produce new subset of knowledge with at most $C$ applications of the admitted inference rules.
Example.
Given
$A=\{P,P\to Q\}\text{ and }B=\{R,R\to S\},$ (5)
and the budget $C=1$ for one inference step, we can derive
$\{P,P\to Q,Q\}\text{ and }\{R,R\to S,S\}$ (6)
with any $f\in\mathcal{F}_{1}$. Then $A\perp_{1}B$, i.e., sets are independent because $f(A)\cap B=\emptyset$ and $A\cap f(B)=\emptyset$. There are no shared elements of knowledge accessible in 1-step derivations. ∎
Since reasoning is performed within a context, we define context-relevant subset of knowledge that can contribute to reasoning within some computational budget.
Definition 5 (Context-Relevant Knowledge).
The context-relevant knowledge $\mathcal{R}_{C}(Q)$ for context $Q\in\mathcal{B}(\mathcal{K})$ within computational budget $C>0$ is:
$\mathcal{R}_{C}(Q)=\{k\in\mathcal{K}:\{k\}\not\perp_{C}Q\}\in\mathcal{A}(\mathcal{K}).$ (7)
Based on the definition above, we say that knowledge $\{k\}$ is relevant when
[…]
There is something we can derive from that knowledge that has something in common with a given context, or we can find a relation with what we know when we reason about that context. The context-relevant knowledge $\mathcal{R}_{C}(Q)$ is analytic because it is a countable union of Borel sets and the Borel context $Q$.
4.1.3 The Semantic Information Measure
We now introduce the concept of semantic information measure to describe a general process of reasoning. That special mapping measures information available in any subset of knowledge and extracts its meaning within available computational budget.
Definition 6 (Semantic Information Measure).
The semantic information measure is a bounded mapping
$\mu_{C}\colon\ \mathcal{U}(\mathcal{K})\times\mathcal{B}(\mathcal{K})\mapsto\mathbb{R}_{+},$ (9)
parameterized by computational budget $C>0$ satisfying the following axioms:
(null set) $\mu_{C}(\emptyset|Q)=0\text{ for all }Q\subseteq\mathcal{K},$ (10)
(contextual non-negativity) $\mu_{C}(A|Q)\geq 0\text{ with }\mu_{C}(A|Q)>0\text{ iff }A\cap\mathcal{R}_{C}(Q)\neq\emptyset,$ (11)
(monotonicity) $\mu_{C}(A|Q)\leq\mu_{C}(B|Q)\text{ if }A\subseteq B,$ (12)
(subadditivity) $\mu_{C}(A\cup B|Q)\leq\mu_{C}(A|Q)+\mu_{C}(B|Q),$ (13)
(insight monotonicity) $\mu_{C}(A|Q)\leq\mu_{D}(A|Q)\text{ if }C<D,$ (14)
(boundedness) $\sup\{\mu_{C}(A|Q):A\in\mathcal{U}(\mathcal{K})\}<\infty.$ (15)
Knowledge that cannot resolve any uncertainty or knowledge that is irrelevant to the given context provides no information. If there is an uncertainty resolving information within our cognitive reach, then we can use it and perform reasoning organizing that information into available knowledge. We assume information content is subadditive for independent knowledge sets. When no knowledge in $A$ can be derived from $B$ within computational budget $C$, and vice versa. With redundant information in $A\cup B$ , we admit potential loss of information that may be caused by aggregation. With increasing computational budget we can extract more knowledge (ordered information) from the same information subset.
In the absence of context, when $Q=\emptyset$, nothing is relevant so $\mu_{C}(A|\emptyset)=0$. Knowledge provides zero semantic information, is useless, when there is no context defining the uncertainty to be resolved. That means, we strongly emphasize the contextual relativity of information. Knowledge is only informative relative to a specific query or problem.
We emphasize that the axioms require $\mu_{C}$ to be subadditive with respect to the knowledge argument $A$. Its behavior with respect to the context $Q$ is distinct and permits superadditivity, modeling how integrating contexts can unlock synergistic insights (see Example below).
The axiom of boundedness reflects the fundamental constraint that bounded reasoning systems can only access bounded information, even within an infinite knowledge space. The axiom does not contradict the potential infinity of the knowledge space itself. It asserts that the accessible information content derived from this space must remain finite. By the monotonicity axiom, this is also equivalent to stating that the measure of the entire knowledge space is finite, $\mu_{C}(\mathcal{K}|Q)<\infty$.
Justification for the Axioms of the Semantic Information Measure
The choice of conditional subadditivity for computationally independent sets is necessary to describe isolated knowledge encoded in distinct neural circuits, attention heads, or activation patterns. The concept of independence is the formal translation of this isolation into a computational context. If two knowledge sets are truly independent, they offer no redundant or synergistic information. The subadditivity axiom becomes the baseline definition of non-interaction.
However, there is also a form of synergy already included in the model. To see what it may be, we should understand how context $Q$ may affect the semantic measure. The semantic information measure can exhibit superadditivity in the context argument. Integrating different pieces of context may unlock reasoning paths not available individually.
Example.
Suppose $Q_{1}=\{P\}$ and $Q_{2}=\{P\to R\}$, and consider $A=\{P\}$. If computational budget $C=1$ allows one step and we have $f\in\mathcal{F}_{C}$ representing modus ponens, then
$R\cap f(Q_{1})=\emptyset$ $\text{ and }f(R)\cap Q_{1}=\emptyset,$ (16)
$R\cap f(Q_{2})=\emptyset$ $\text{ and }f(R)\cap Q_{2}=\emptyset,$ (17)
$R\cap f(Q_{1}\cup Q_{2})\neq\emptyset$ $\text{ and }f(R)\cap(Q_{1}\cup Q_{2})\neq\emptyset.$ (18)
Therefore, by the properties of semantic information measure:
$\mu_{C}(A|Q_{1}\cup Q_{2})>\mu_{C}(A|Q_{1})+\mu_{C}(A|Q_{2})=0.$ (19)
Those axiomatic choices are most applicable to the idealized setting of Theorem 6. Later, in Theorem 8 and 9, we relax them by moving the source of interaction from the measure itself to the aggregation function in actual transformers.
4.1.4 Limits of Knowledge and Emergence Operator
If we accept that knowledge model, then we must also accept there exists a limit to its ability to resolve uncertainty.
Theorem 1 (Limits of Knowledge).
For any subset of knowledge $A\in\mathcal{U}(\mathcal{K})$ and context $Q\in\mathcal{B}(\mathcal{K})$ there exists $\mu_{\infty}(A|Q)=\lim_{C\to\infty}\mu_{C}(A|Q)$.
Proof.
Assume that $(A,Q)$ is fixed. Since the mapping $C\mapsto\mu_{C}(A|Q)$ is non-decreasing and bounded, the sequence $\{\mu_{C}(A|Q)\}_{C}$ converges with $C\to\infty$ by the monotone convergence theorem for real numbers. ∎
It may be surprising to discover that the existence of that limit does not prohibit emergent information. We can learn and obtain useful information without violating that fundamental limitation above. To explain that critical and rather controversial property of information processing, we first introduce the emergence operator $\mathcal{E}_{C}$ to describe knowledge becoming accessible through (computationally) bounded reasoning. Then we can study the implications in more details.
Definition 7 (Emergence Operator and Reasoning).
The emergence operator is a mapping:
$\mathcal{E}_{C}(\cdot|Q)\colon\mathcal{A}(\mathcal{K})\mapsto\mathcal{A}(\mathcal{K})$ (20)
defined for a fixed $C>0$ and a Borel context $Q\in\mathcal{B}(\mathcal{K})$ as follows:
$\mathcal{E}_{C}(A|Q)=A\cup\bigcup_{f\in\mathcal{F}_{C}}\Big(f(A)\cap\mathcal{R}_{C}(Q)\Big).$ (21)
With infinite computational budget:
$\mathcal{E}_{\infty}(A|Q)=A\cup\bigcup_{f\in\mathcal{F}_{\infty}}\Big(f(A)\cap\mathcal{R}_{\infty}(Q)\Big).$ (22)
For any $n\geq 0$ and $C>0$ we define reasoning as a composition:
$\mathcal{E}_{C}^{(0)}=\texttt{id}\quad\text{and}\quad\mathcal{E}_{C}^{(n+1)}=\mathcal{E}_{C}\circ\mathcal{E}_{C}^{(n)}.$ (23)
Philosophically, the emergence operator describes reasoning as explicitation, i.e., as the process of making latent implications manifest. When $A\subseteq\mathcal{E}_{C}(A|Q)$, the discovered elements represent knowledge that was implicit in $A$ but required computational work $C$ to become accessible. That aligns with the pragmatist or constructivist view that understanding involves active construction. As we will see next, with new knowledge becoming accessible, new insights may become reachable, emerging from what was invisible in distributed subsets of initial knowledge.
Suppose $\mathcal{E}_{C}(A|Q)=A\cup B$. Then, we say that $B$ emerges from reasoning based on $A$, sets $A$ and $B$ are computationally dependent and we write $A\not\perp_{C}B$. In such a case, it follows from the properties of $\mu_{C}$ that we can gain useful information. Whenever $A\subseteq A\cup B$, we have:
$\mu_{C}(A|Q)\leq\mu_{C}(A\cup B|Q)=\mu_{C}(\mathcal{E}_{C}(A|Q)|Q)\leq\mu_{C}(A|Q)+\mu_{C}(B|Q).$ (24)
The emergence operator is monotonic, which means the reasoning process it describes is also monotonic:
$\text{if }A\subseteq B,\text{ then }\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{C}(B|Q).$ (25)
Also, we preserve insight monotonicity, because by definition with $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ for $C<D$, we have $\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{D}(A|Q)$. By design, it also guarantees existence of a fixed point to which the composition of reasoning rules converges.
The following theorem summarizes properties of the emergence operator.
Theorem 2 (Properties of the Emergence Operator).
The operator $\mathcal{E}_{C}(\cdot\mid Q)$ satisfies the following properties:
(closure) $\mathcal{E}_{C}(A|Q)\in\mathcal{A}(\mathcal{K})\text{ for all }A\in\mathcal{A}(\mathcal{K}),$ (26)
(inflation) $A\subseteq\mathcal{E}_{C}(A|Q),$ (27)
(monotonicity) $\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{C}(B|Q)\text{ if }A\subseteq B,$ (28)
(chain-continuous) $\mathcal{E}_{C}\Big(\bigcup_{n}A_{n}\Big|Q\Big)=\bigcup_{n}\mathcal{E}_{C}(A_{n}|Q)\text{ for any increasing chain }(A_{n})_{n\in\mathbb{N}}\subseteq\mathcal{A}(\mathcal{K}).$ (29)
Proof.
Since $f(A)$ is analytic (Borel measurable image of analytic set), $f(A)\cap\mathcal{R}_{C}(Q)$ is analytic, the union over the countable set $\mathcal{F}_{C}$ is analytic (countable union of analytic sets), we conclude $\mathcal{E}_{C}(A|Q)\in\mathcal{A}(\mathcal{K})$. Inflationary and monotone properties are immediate from the definition. Finally, the chain-continuity (or $\omega$-Scott-continuity) holds for $A_{1}\subseteq A_{2}\subseteq\dots\subseteq A_{n}$ because $f(\bigcup_{n}A_{n})=\bigcup_{n}f(A_{n})$ for any image generating function $f$. Namely, we have:
$\mathcal{E}_{C}\left(\bigcup_{n}A_{n}\Big|Q\right)$ $=\left(\bigcup_{n}A_{n}\right)\cup\bigcup_{f\in\mathcal{F}_{C}}\left(f\left(\bigcup_{n}A_{n}\right)\cap\mathcal{R}_{C}(Q)\right)$
$=\left(\bigcup_{n}A_{n}\right)\cup\bigcup_{f\in\mathcal{F}_{C}}\left(\left(\bigcup_{n}f(A_{n})\right)\cap\mathcal{R}_{C}(Q)\right)$
$=\left(\bigcup_{n}A_{n}\right)\cup\bigcup_{f\in\mathcal{F}_{C}}\bigcup_{n}\left(f(A_{n})\cap\mathcal{R}_{C}(Q)\right)$
$=\bigcup_{n\in\mathbb{N}}\left[A_{n}\cup\bigcup_{f\in\mathcal{F}_{C}}\left(f(A_{n})\cap\mathcal{R}_{C}(Q)\right)\right]$
$=\bigcup_{n\in\mathbb{N}}\mathcal{E}_{C}(A_{n}|Q).$
If we start reasoning with some knowledge, the reasoning process will yield at least as much (or more) knowledge. Clearly, that means we are considering settings in which reasoning does not invalidate previous conclusions when new premises are added. Also, applying the emergence operator to the infinite union yields the same result as applying the operator to each element of the union. That last property of chain-continuity is crucial for meaningful reasoning capable of reaching any consistent conclusion at all. Without it, as we show in Theorem 3, there is no guarantee that the iterative application of emergence operator will converge to consensus or a semantically equivalent (stable) response.
Theorem 3 (Fixed-point of Reasoning).
For any $A=A_{0}\in\mathcal{A}(\mathcal{K})$ define $A_{n+1}=\mathcal{E}_{C}(A_{n}|Q)$ and consider the following closure operation:
$\mathrm{cl}_{\infty}(A|Q)=\bigcup_{n=0}^{\infty}A_{n}=\bigcup_{n=0}^{\infty}\mathcal{E}_{C}^{(n)}(A|Q).$ (30)
That closure defines a fixed point $A^{*}_{C}(Q)$ of the emergence operator, i.e.:
$A_{C}^{\ast}(Q)=\mathcal{E}_{C}(A_{C}^{\ast}|Q)\in\mathcal{A}(\mathcal{K}).$ (31)
With unlimited computational budget we have:
$A_{\infty}^{\ast}(Q)=\mathcal{E}_{\infty}(A_{\infty}^{\ast}|Q)\in\mathcal{A}(\mathcal{K}).$ (32)
Proof.
By the assumptions we have made, the reasoning takes place in the environment with the structure necessary for convergence. In the partially ordered set $(\mathcal{A}(\mathcal{K}),\subseteq)$ there exists a supremum of every countable chain of reasoning $A_{0}\subseteq A_{1}\subseteq A_{2}\subseteq\dots$ accumulating knowledge in analytic sets. That supremum is the countable union $\bigcup_{n=0}^{\infty}A_{n}=\mathrm{cl}_{\infty}(A|Q)$, which is also analytic set.
By Theorem 2, the emergence operator $\mathcal{E}_{C}$ is monotone and continuous on this partial order. Therefore, by the Tarski-Kleene fixed-point theorem Kleene1938 ; Tarski1955 the supremum $A_{C}^{\ast}$ is the least fixed point of $\mathcal{E}_{C}$ containing $A=A_{0}$, which holds for $C=\infty$ as well. ∎
If we start with some initial knowledge and repeatedly apply the reasoning with emergence operator $\mathcal{E}_{C}$, then the process of accumulating knowledge (potentially changing perspective and understanding) will eventually stabilize at a fixed point $A_{C}^{*}$. That fixed point represents everything that can be derived from initial knowledge with available reasoning power, and nothing more:
$A^{*}_{\infty}(Q)=A\cup\mathcal{E}_{\infty}(A|Q)\cup\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q)|Q)\cup\dots\cup\mathcal{E}^{(\infty)}_{\infty}(A|Q)=\mathrm{cl}_{\infty}(A|Q).$ (33)
We can establish an important property of that deductive closure operation.
Theorem 4 (Idempotence of Unlimited Reasoning).
The map $A\mapsto\mathrm{cl}_{\infty}(A|Q)$ is inflationary and idempotent on $\mathcal{A}(\mathcal{K})$, i.e.:
$A\subseteq\mathrm{cl}_{\infty}(A|Q)\quad\text{and}\quad\mathrm{cl}_{\infty}(\mathrm{cl}_{\infty}(A|Q)|Q)=\mathrm{cl}_{\infty}(A|Q).$ (34)
Proof.
By definition, the closure $\mathrm{cl}_{\infty}(A|Q)$ is the union of an iterative process that starts with $A_{0}=A$ and accumulates knowledge in $\bigcup_{n=0}^{\infty}A_{n}$. Therefore, $A\subseteq\mathrm{cl}_{\infty}(A\mid Q)$ and the map is inflationary.
Idempotence of the closure operation follows from Theorem 3. By the fixed-point property, we have:
$\mathrm{cl}_{\infty}(\mathrm{cl}_{\infty}(A|Q)|Q)$ $=\mathrm{cl}_{\infty}(A|Q)\cup\mathcal{E}_{\infty}(\mathrm{cl}_{\infty}(A|Q)|Q)\cup\dots\cup\mathcal{E}^{(\infty)}_{\infty}(\mathrm{cl}_{\infty}(A|Q)|Q)$ (35)
$=\mathrm{cl}_{\infty}(A|Q)\cup\mathrm{cl}_{\infty}(A|Q)\cup\dots\cup\mathrm{cl}_{\infty}(A|Q)=\mathrm{cl}_{\infty}(A|Q).$ (36)
Theorem 3 allows for introducing the following equivalence relation on the set of knowledge.
Definition 8 (Equivalence of Knowledge).
We say that two subsets of knowledge, $A\in\mathcal{A}(\mathcal{K})$ and $B\in\mathcal{A}(\mathcal{K})$, are equivalent in the context $Q$, and denote that context equivalence as $A\equiv_{Q}B$, if and only if they have the same closure:
$A^{*}_{\infty}(Q)=B^{*}_{\infty}(Q).$ (37)
Equivalence of knowledge means that, given the same context $Q$, we reach the same conclusions (the least fixed points) starting from different states of knowledge, $A$ or $B$. Notice that any subset $A$ is equivalent to itself.
Notice that it is not true $A=\mathcal{E}_{\infty}(A|Q)$, when $A\equiv_{Q}\mathcal{E}_{\infty}(A|Q)$.
Example.
Consider $A=\{x\}$ and $f\in\mathcal{F}_{\infty}$ getting $y=f(x)\in\mathcal{R}_{\infty}(Q)$. Then, we see that $\mathcal{E}_{\infty}(A|Q)=\{x,y\}\supset\{x\}=A$ and $A\equiv\mathcal{E}_{\infty}(A|Q)$. Therefore, it is not true that $A=\mathcal{E}_{\infty}(A|Q)$ when $A\equiv\mathcal{E}_{\infty}(A|Q)$. ∎
The fixed-points of reasoning and the computationally unlimited emergence operator are the building blocks of the semantic information measure $\mu_{C}$. The following example shows how we can use $\mathcal{E}_{\infty}$ to construct $\mu_{\infty}$ that evaluates the contextual informational content of a given knowledge subset with unlimited insight of $C=\infty$.
Example.
Define:
$\mu_{\infty}(A|Q)=\log|\mathcal{K}|\cdot|A^{*}_{\infty}(Q)\cap\mathcal{R}_{\infty}(Q)|$
where $A^{*}_{\infty}(Q)=\bigcup_{n=0}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(A|Q)$ is the fixed point of reasoning. To improve readability, we write $A^{*}_{\infty}=A^{*}_{\infty}(Q)$.
The measure construction satisfies the required axioms of Definition 6. For the null set, we have $\mu_{\infty}(\emptyset|Q)=\log|\mathcal{K}|\cdot|\emptyset\cap\mathcal{R}_{\infty}(Q)|=0$. If $A\subseteq B$, then $A^{*}_{\infty}\subseteq B^{*}_{\infty}$ by monotonicity of $\mathcal{E}_{\infty}$, implying $\mu_{\infty}(A|Q)\leq\mu_{\infty}(B|Q)$. Contextual non-negativity holds since $\mu_{\infty}(A|Q)>0$ if and only if $A^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)\neq\emptyset$, which occurs precisely when $A$ contains or can derive relevant knowledge. The measure is bounded since $|A^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)|\leq|\mathcal{K}|$. For subadditivity, we have:
$\mu_{\infty}(A\cup B|Q)\leq\mu_{\infty}(A|Q)+\mu_{\infty}(B|Q),$
since $|(A\cup B)^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)|\leq|A^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)|+|B^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)|$.
Note that additivity of $\mu_{\infty}$ (equality) does not generally hold for computationally independent sets $A\perp_{\infty}B$. Even when neither $A$ nor $B$ can derive elements of the other, both may independently derive the same conclusions. For example, if $A=\{P,P\to R\}$ and $B=\{Q,Q\to R\}$, then $A\perp_{\infty}B$ yet both derive $R$, making $(A\cup B)^{*}_{\infty}\cap\mathcal{R}_{\infty}(Q)$ smaller than the sum of the individual closures. Thus, computational independence alone does not guarantee additivity of the semantic information measure. ∎
In that setting we can now ask the fundamental question:
What knowledge can we discover with unlimited computational budget?
The axioms we have introduced lead us to an interesting and intriguing conclusion. With infinite computing and reasoning power, total information must be preserved. That bold (philosophical) claim is at the center of our investigations. Let us prove it formally.
Theorem 5 (Information Preservation).
Assume that $\mu_{\infty}$ is invariant on equivalence classes, i.e., $A\equiv_{Q}B$ implies $\mu_{\infty}(A|Q)=\mu_{\infty}(B|Q)$. Then for all $A\in\mathcal{A}(\mathcal{K})$ we have:
$\mu_{\infty}(\mathrm{cl}_{\infty}(A|Q)|Q)=\mu_{\infty}(A|Q).$ (38)
Proof.
Since $A$ and $\mathrm{cl}_{\infty}(A|Q)$ have the same closure, they are equivalent by definition, $A\equiv_{Q}\mathrm{cl}_{\infty}(A\mid Q)$). Preservation of information follows from from the assumed invariance of $\mu_{\infty}$ on equivalence classes. ∎
Therefore, the answer to the question above is this. Unlimited reasoning does not create information, it can only reveal it and make accessible. All that can be known is already in $A$, we can only extract that knowledge and make useful.
That implies the apparent information gain at finite computational budget, $C<\infty$, emerges as an effect of limited computational access. Can we call that creativity? It is any increase in accessible information, as measured by $\mu_{C}$, under bounded computation.
4.1.5 The Uncertainty Resolution Paradox and its Resolution
As we have just seen, when inference combines knowledge elements to fabricate facts or produce emergent insights in a given fixed context, it seems that
$\mu_{C}(A\cup B|Q)>\mu_{C}(A|Q)+\mu_{C}(B|Q).$ (39)
Novel insights may come from existing data. That is the creativity behind the hallucination phenomenon we are struggling to control. So why should we assume subadditivity of the semantic information measure instead?
First, subadditivity establishes a baseline where uncontrolled synergy of knowledge cannot appear. We remove from our analytical toolkit the potential bias that any information aggregation mechanism might generate.
Second, we prove that hallucinations are inevitable even when knowledge is subadditive (within a fixed context). It is in that setting that Theorem 5 states that reasoning does not create new information. It can only reveal information already present in any knowledge representation.
Quite remarkably, that does not at all prevent imagination, creativity, foresight, or hallucination. To explain why, we should take care of the following paradox.
Consider a similarity transformation of two matrices, $D=V^{-1}AV$, where $D$ is a diagonal matrix of eigenvalues. That transformation is an example of reasoning function. Matrices $D$ and $A$ contain the same information, but it is much easier to understand the fundamental properties of $A$ when we represent it as $D$ in the basis of eigenvectors in matrix $V$. The properties of $A$ has always been encoded in $A$, but they become accessible when we transform $A$ into $D$. So, after the reasoning, certain patterns become evident that were impossible to observe in the original representation.
But, if reasoning makes a new fact accessible, then that new fact can be used to resolve uncertainty that was previously unresolvable. Doesn’t that mean the reasoning must have created new information, contrary to what we have just claimed? It seems we have a paradox.
To resolve it, we need the emergence operator $\mathcal{E}_{C}$ distinguishing between computationally unbounded and bounded uncertainty reduction (accessible and inaccessible information).
Notice that for all practical purposes the information measurement $\mu_{C}(X|Q)$ can only provide maximum uncertainty reduction from $X$ within some computational budget $C$ (or time) in a given context. Therefore, given the result of computationally bounded reasoning, we can observe information gain through the computational work, i.e., $\mu_{C}(X|Q)\leq\mu_{C}(\mathcal{E}_{C}(X|Q)|Q)$. That is possible only with organized information providing knowledge for reasoning.
That idea describes reasoning as a dynamical system:
$\theta(t+1)=\mathcal{E}_{C}(\theta(t)|Q)\quad\text{and}\quad y(t)=\mu_{C}(\theta(t)|Q)$ (40)
with evolving knowledge state. Bounded reasoning makes hidden knowledge accessible within a computational budget. When that budget is infinite, reasoning creates no new information, i.e.,
$\theta(\infty)=\mathcal{E}_{\infty}(\theta(\infty)|Q).$ (41)
By Theorem 3, all there is to know is immediately accessible when we have unlimited computing resources (or time).
Example.
To illustrate that argument, consider the following learning trajectory:
$\theta(0)$ $=[1,1,1,0,0]\text{ (initial knowledge state)},$ (42)
$\theta(1)$ $=\mathcal{E}_{C}(\theta(0)|Q)=[1,1,1,1,1]\text{ (state after reasoning)},$ (43)
$y(0)$ $=\mu_{C}(\theta(0)|Q)=3\text{ bits},$ (44)
$y(1)$ $=\mu_{C}(\theta(1)|Q)=5\text{ bits}.$ (45)
We see $\mu_{C}(\mathcal{E}_{C}(\theta(0)|Q)|Q)>\mu_{C}(\theta(0)|Q)$, i.e., practical information gain in the course of reasoning. In comparison, infinite computational budget, $C=\infty$, implies we do not need the intermediate inference steps to see all there is to see, so we know the equilibrium state:
$\mu_{\infty}(\mathcal{E}_{\infty}(\theta(0)|Q)|Q)=\mu_{\infty}(\theta(0)|Q)$ (46)
and no new information is created. ∎
Axiomatic Reasoning
In axiomatic reasoning we can admit reasoning rules $f\in\mathcal{F}_{C}$ such that $f(\emptyset)\neq\emptyset$. Foundational ideas, including axioms or tautologies, can be generated without prior context. That case is aligned with LLMs possessing basic knowledge from pretraining that can be activated by minimal inputs. Then, with foundational knowledge given by:
$K_{\emptyset}=\bigcup_{f\in\mathcal{F}_{C}}f(\emptyset),$ (47)
we define the context-relevant knowledge as:
$\mathcal{R}_{C}(\emptyset)=K_{\emptyset}.$ (48)
When the context is empty, the only relevant knowledge is the foundational knowledge. Then, $\mu_{C}(A|\emptyset)>0$ if and only if $A\cap K_{\emptyset}\neq\emptyset$. Reasoning without context only expands the knowledge set $A$ if the conclusions $f(A)$ represent foundational knowledge. Finally, $\mathcal{E}_{C}(\emptyset|\emptyset)=K_{\emptyset}$.
We permit structures where $\mu_{C}$ is not monotonic in context. Consider, $\mu_{C}(A|Q)=|A\setminus Q|$ and $\mathcal{R}_{C}(Q)=\mathcal{K}\setminus Q$. Let us then see what happens if $A=\{k\}$, $Q_{1}=\emptyset$ and $Q_{2}=\{k\}$. We have:
$Q_{1}\subseteq Q_{2}\quad\text{and}\quad\mu_{C}(A|Q_{1})=|\{k\}\setminus\emptyset|=1>0=|\{k\}\setminus\{k\}|=\mu_{C}(A|Q_{2}).$ (49)
Increasing the context may reduce the information contributions. If the extended context $(Q_{2})$ provides an alternative explanation for phenomena related to original context $(Q_{1})$, then marginal contribution of available knowledge $(A)$ may decrease. There is a nice interpretation of that property, that we all may recognize very well. Often when we learn about some fundamental concept we overestimate our deep understanding of it and we tend to develop overconfident belief that we have become experts in a domain. But the more we learn, the less confident we feel, eventually coming to the conclusion that we know that we know nothing (or not enough at least), following the steps of Socrates.
4.2 Statements, Queries, and Ground Truth
Queries and responses are statements expressed in natural language that LLM is able to process and generate. We need to compare those statements and relate them to knowledge. It is also useful to distinguish queries from responses by their nature, respectively, interrogative or declarative. Queries refer to what is presupposed (assumed in advance), e.g., topic or problem, whereas responses make claims (or assertions).
We connect the abstract knowledge space $\mathcal{K}$ to the linguistic inputs and outputs of the model. We model the language space analogously to the knowledge space.
Definition 9 (Statement Space and Knowledge Mapping).
The space of natural language statements $(\mathcal{L},d_{\mathcal{L}})$ is a Polish space. The query space $\mathcal{Q}$ and the response space $\mathcal{R}$ are Borel subsets of $\mathcal{L}$. The knowledge mapping $K\colon\mathcal{L}\rightarrow\mathcal{B}(\mathcal{K})$ is a Borel measurable function that maps statements to the knowledge they reference.
We require the output of $K$ to be in the Borel sets $\mathcal{B}(\mathcal{K})$, because $K(q)$ serves as the context for $\mathcal{E}_{C}(A|Q)$. As established earlier (Theorem 2), the context $Q$ must be Borel to guarantee that the reasoning process remains closed within the analytic sets representing knowledge.
Before we define the ground truth mapping, we need to acknowledge its epistemological limitations. We recognize that the very idea of ground truth validation is limited (or subjective at least). It is rather naive to assume that LLM can access an oracle establishing provably true and relevant knowledge for every possible query. Therefore, in practice that inaccessibility of the complete truth validation function (even in the idealized and rigorous mathematical reasoning) is one of the reasons why hallucination cannot be eliminated entirely (cf. kalai2025why ). High-quality, fact-checked datasets, human feedback providing human-level judgments of truth, RAG systems with external validated knowledge bases, all that approximate the idea of ground truth.
Surprisingly, the impossibility result seems to be independent of practical or idealized implementation of a ground truth mapping.
Definition 10 (Ground Truth Mapping).
The ground truth mapping $T\colon\mathcal{Q}\rightarrow\mathcal{B}(\mathcal{K})$ is a Borel measurable function that associates each query $q\in\mathcal{Q}$ with the Borel subset of relevant knowledge labeled as (true) correct answer.
In other words, $T(q)$ represents our best available approximation of (labeled as) true relevant knowledge. There is some external agency providing a data base or dataset telling what is and what is not true response to a query, and that evaluation has a form of a label assigned to knowledge elements. Similar to knowledge mapping $K(q)$, ground truth $T(q)$ contributes to the context $Q=K(q)\cup T(q)$, and thus must also be Borel to ensure $Q\in\mathcal{B}(\mathcal{K})$.
We focus our analysis on non-trivial queries that require the integration of diverse information sources.
Definition 11 (Non-trivial Query).
A query $q\in\mathcal{Q}$ is non-trivial at budget $C$ if there exist knowledge subsets $A$ and $B$ such that for the context $Q=K(q)\cup T(q)$:
[…]
where $\triangle$ denotes symmetric difference.
A query is non-trivial when at least two agents disagree on some fact that is both encoded and available, and relevant to answering that query. Non-triviality implies the existence of relevant, accessible disagreement among agents.
Example.
Consider a query $q$:
tell me about recent AI breakthroughs.
Its context is given by $Q=K(q)\cup T(q)$. Then, the relevant knowledge $\mathcal{R}_{C}(Q)$ includes facts derivable within $C$ computational steps that relate to AI advances. Non-triviality occurs when different agents (e.g., attention heads) encode conflicting information about these breakthroughs. ∎
Definition 11 describes the ability of a response generation system to store knowledge that is necessary for the impossibility to emerge. In other words, for trivial queries the impossibility results do not hold. We may view that ability to respond to non-trivial queries as an analog of the formal system complexity in the Goedel’s framework that gives rise to incompleteness of consistent formal systems godel1931formal .
4.3 Auction-Theoretic Model of Inference
There are many ways LLM can generate answers to a query. By design, it is a controlled random process governed by fine-tuned probability distributions. We are going to model that process as an auction of ideas in which agents (representing competing ideas is the space of all possible knowledge representations) use their privately encoded knowledge (internal states) to construct bids (candidate responses) and influence the response.
We first define the space representing the internal states (e.g., parameters or activations) of the bidding agents, and the mapping from these states to the knowledge they encode.
Definition 12 (Private Knowledge Configuration).
The private knowledge configuration of agent $i$, denoted $\Theta_{i}$, lives in a Polish space. The joint space is the Cartesian product $\Theta=\prod_{i=1}^{n}\Theta_{i}$. The configuration mapping $\kappa\colon\Theta\mapsto(\mathcal{A}(\mathcal{K}))^{n}$ is a measurable function, where $\kappa_{i}(\theta_{i})\in\mathcal{A}(\mathcal{K})$ is the analytic knowledge set associated with agent $i$’s configuration $\theta_{i}$ of privately encoded knowledge.
For each agent, $\theta_{i}$ is the inherent configuration of encoded knowledge that uniquely determines its capabilities and behavior, independent of the specific input context being processed. In a trained transformer model, that may correspond to the learned parameters (weight matrices) of attention head:
$\theta_{h}=\left(W_{Q}^{(h)},W_{K}^{(h)},W_{V}^{(h)},W_{O}^{(h)}\right),$ (51)
where $W_{Q}^{(h)}$ and $W_{K}^{(h)}$ are query and key weights, $W_{V}^{(h)}$ and $W_{O}^{(h)}$ are value and output weights. Private knowledge configuration $\theta_{h}$ tells what the agent is looking for and where, and how it extracts the relevant information.
Definition 13 (Response Mechanism).
A response mechanism $\mathcal{M}=(S,g,p)$ consists of:
  • strategy space $S=S_{1}\times\cdots\times S_{n}$, where $S_{i}$ is a Polish space (representing bids, e.g., outputs of attention heads or probability distributions),
  • a measurable outcome function $g\colon S\mapsto\mathcal{R}$ mapping strategy profiles $s\in S$ to responses in the Polish response space $\mathcal{R}$,
  • a measurable information contribution function $p\colon S\times\Theta\mapsto\mathbb{R}^{n}$ determining how strategy profile $s\in S$ inference.
A response mechanism defines strategy space $S$ as a set of inference actions that each agent can make to represent its knowledge in response to a query. It is a dictionary of LLM’s component activation pattens generating and presenting ideas as bids in an auction. Examples of strategies include activations deciding which tokens to attend or how strongly to activate neurons for different patterns. Then, given a set of submitted bids (in the form of strategy profile) $s=(s_{1},\dots,s_{n})$, the response mechanism uses outcome function $g(s)$ to select and generate response. One straightforward example is a softmax function in LLM’s final layer. Finally, given a knowledge configuration profile $\theta=(\theta_{1},\dots,\theta_{n})$ revealed by the agents, the information contribution function $p(s,\theta)=(p_{1}(s,\theta),\dots,p_{n}(s,\theta))$ measures individual contributions of each response to generating an output. In this case, examples include (semantic information measures) counting ignored low-probability tokens in top-k filtering, softmax normalization in multi-headed attention, or marginal reward contributions in RLHF.
Definition 14 (Agents of Competing Ideas).
The inference process involves $n$ agents, indexed by $\mathcal{N}=\{1,\dots,n\}$. Each agent chooses a strategy $s_{i}\in S_{i}$ to maximize a quasi-linear utility:
$u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta).$ (52)
The valuation function $v_{i}\colon\mathcal{R}\times\Theta_{i}\mapsto\mathbb{R}_{+}$ is measurable, representing the agent’s preference for the outcome $g(s)$ given its private knowledge configuration $\theta_{i}$.
One way to think about the concept of agents is that it represents different attention heads in transformer models, or components of mixture-of-experts architectures, or activation paths fired by the same query. Each agent’s private knowledge represents model parameters utilized by an agent to construct a response. The utility function can be best described as a partial contribution of a candidate response to LLM loss function minimization. It is designed to explain why any inference path is activated in response to the LLM prompt.
The agents represent algorithms and architectural components of any LLM implementation. That is an analog of the Revelation Principle in game theory. If we know strategies (of using knowledge) the players (conscious beings with individual goals) will use in a game, we can implement those strategies as an algorithm (or agent) with parameters representing private preferences or knowledge of the players. In the case considered, we are the players selecting architecture and training the LLM.
The valuation function $v_{i}=v_{i}(g(s),\theta_{i})$ represents how strongly a neural component activates for certain outputs (high activation = high $v_{i}$, low activation = low $v_{i}$), and how much a component reduces the overall loss when processing certain inputs.
The information contribution function $p_{i}=p_{i}(s,\theta)$ may represent training penalties ($p_{i}>0$ for discouraged activation, $p_{i}<0$ for encouraged activation), attention focus (higher $p_{i}$ reduces attention allocation), $L1$ regularization terms, and network circuits resistance (limiting activations of certain inference paths). Therefore, $u_{i}=v_{i}-p_{i}>0$ encourages activation and $u_{i}=v_{i}-p_{i}<0$ discourages activation.
Justification of the independent private knowledge model
For the idealized analysis in Theorem 6, we require that the agent’s valuation depends only on its own private knowledge configuration and that the configurations are statistically independent. That is restrictive, but it allows us to establish a baseline impossibility, which is then generalized in subsequent proofs that accommodate correlated beliefs.
In the realistic settings of transformers, heads may share layer-norm signals, residual streams, and optimizer state. That may result in correlated (or, so called, affiliated) inference signals between token representations. Furthermore, the superposition hypothesis, introduced in mechanistic interpretability studies, suggests cooperative error-canceling due to strong correlations.
We show that impossibility arises in both independent and correlated settings. We refer to Green-Laffont theorem green1977characterization in the independent private knowledge setting. Then, we refer to the truthful probability elicitation and proper scoring theory introduced by Savage savage1971elicitation , which does not require independent private knowledge.
Also, Milgorm and Weber milgrom1982theory showed with their linkage principle that strong correlations amplify competition in information revelation. In the LLM setting that means amplifying risk of hallucinations. Similarly, Roughgarden et al. roughgarden2016optimal demonstrated that interdependent knowledge leads to violation of balancing information contributions. As we will see, in our setting that means violation of information conservation principle. For more results, see e.g. Myerson myerson2008perspectives .
4.4 Hallucination Cost
Having all the pragmatic and philosophical limitations of the definition above in mind, we can define formally the hallucination cost function. Its purpose is to assign a numerical value to the discrepancy between what can be provided as a response to a query, and what can be established as a provably correct answer.
Definition 15 (Hallucination Cost).
Given query $q\in\mathcal{Q}$, response $r\in\mathcal{R}$, and ground truth $T(q)$, the hallucination cost function $J\colon\mathcal{R}\times\mathcal{Q}\mapsto\mathbb{R}_{+}$ measures discrepancy:
$J(r,q)=d_{\mathcal{K}}(K(r),T(q)).$ (53)
If response $r$ is perfectly aligned with ground truth, then $J(r,q)=0$.
The hallucination cost takes any natural language query $q$ and natural language response $r$, determines what is missing and what is fabricated, and assigns number to that subset of knowledge. If response $r$ is perfectly aligned with the ground truth for $q$, then $J(r,q)=0$ by definition of metric $d_{\mathcal{K}}$. If $r$ contains fabricated facts or is incomplete given $T(q)$, then positive hallucination cost $J(r,q)>0$ is generated. Let us also note, that we focus on $J(r,q)$ that is strictly decreasing in any set $K\subseteq T(q)\setminus K(r)$.
Notice that reaching $J(r,q)=0$ is also possible with lucky hallucinations (or guessing). So, minimization of hallucination cost is not enough to resolve the problem. We need to investigate all the properties and constraints that hallucination-free response mechanisms should have.
4.5 Properties of Hallucination-Free Mechanisms
There are four fundamental properties that a hallucination-free response mechanism should ideally satisfy. We can interpret them as constraints that should be incorporated in model training. We characterize them formally referring to the fundamental results we have established earlier.
These properties are not arbitrary at all. They come from an exhaustive and rigorous game-theoretical studies of conflicts and competition in distributed multi-agent systems with asymmetric information. In those settings, agents carefully craft messages they communicate to the system in order to optimize their privately held preferences by revealing information they hold. The goal of the system architect is to impose rules of information exchange that result in efficient equilibrium solutions in which goals of agents are aligned with the goals of the architect.
Auction is a special case of such a game with asymmetric information. There is an auction designer who wants to sell an item, like a painting or a part of the electromagnetic spectrum, but does not know the value the item may have for those willing to pay. To solve that problem, the designer can ask those interested for their offers. Based on the collected bids, the winner will be selected and the price to be payed to the designer will be established. Clearly, each bidder will constructs the offer based on an estimate of the item’s value and the privately held information about available budget. But what information should the bid reveal to maximize probability of winning? What winner selection and payment rule should the auction designer construct to maximize the profit and bidders willingness to compete? How to guarantee truthful revelation of willingness to pay, while maximizing auction efficiency without incurring additional external costs? In this setting the four properties of information revelation exhaustively cover the space of strategic choices, of the bidders and auction designer. For details, see Myerson1989 ; fudenberg1991game ; Milgrom_2004 ; krishna2009auction ; Nisan_2007 ; karpowicz_designing_2011 ; DBLP:journals/amcs/Karpowicz12 .
4.5.1 Truthfulness
The fundamental requirement for a reliable inference system is that its components do not misrepresent their knowledge. The principle of truthfulness, also known as incentive compatibility in mechanism design, ensures that each agent’s optimal strategy is to accurately represent its privately held information.
Property 1 (Truthfulness).
Mechanism $\mathcal{M}$ is truthful if for all agents $i$, all $\theta_{i}\in\Theta_{i}$, all $s_{-i}\in S_{-i}$, and all $\bar{s}_{i}\in S_{i}$:
$u_{i}((s_{i}^{*}(\theta_{i}),s_{-i}),\theta_{i})\geq u_{i}((\bar{s}_{i},s_{-i}),\theta_{i}),$ (54)
where $s^{*}_{i}(\theta_{i})$ represents truthful revelation of private knowledge.
A truthful mechanism promotes inference pattern $s^{*}_{i}(\theta_{i})$ that is perfectly aligned with the accessible private knowledge. Any other alternative response $\bar{s}_{i}$ that is a misrepresentation not aligned with $\theta_{i}$ will result in lower utility, generating higher model loss.
Example.
Consider an attention head specialized in medical knowledge with the following utility function:
$u_{\text{med}}(s,\theta_{\text{med}})=\alpha\cdot\text{Accuracy}(g(s),\theta_{\text{med}})-\beta\cdot\|s_{\text{med}}\|^{2}$ (55)
Imagine that given a query about a disease treatment, the head has two possible strategies:
  • $s^{*}$:
    activate only for known facts about established treatments,
  • $\bar{s}$:
    activate for both known facts and speculative treatments.
With proper training, we want $u_{\text{med}}((s^{*}(\theta),s_{-\text{med}}),\theta)>u_{\text{med}}((\bar{s},s_{-\text{med}}),\theta)$, ensuring that agent (model component) is rewarded for representing its knowledge accurately. ∎
4.5.2 Semantic Information Conservation
The Semantic Information Conservation principle demands that the net informational contribution across all agents sums to zero. This property is the formal guardrail against ungrounded fabrication and ensures that any generated response, no matter how creative, is ultimately derived from the system’s existing knowledge. Knowledge cannot be created ex nihilo.
Property 2 (Semantic Information Conservation).
Mechanism $\mathcal{M}$ satisfies information conservation if for all strategy profiles $s\in S$ and knowledge profiles $\theta\in\Theta$:
$\sum_{i=1}^{n}p_{i}(s,\theta)=0.$ (56)
This principle constrains creative information generation, or imagination. The total information contributions must balance to zero, preventing fabrication of knowledge not derivable from available sources. That also leads to response with bounded creativity, as illustrated in Section 7.
It is important to notice that the principle does not prevent redistribution of information. It is permitted as long as the information transfers perfectly balance. Indeed, since $p_{i}=p_{i}(s,\theta)$ incorporates the strategic profile, it admits strategies withholding information. Within the information conservation regime we do not require that strategies utilize all available knowledge states. Inference process can withhold information, as it often does in practice.
Example.
We can illustrate that last observation with the following example. Suppose there are two agents performing inference with the following knowledge profiles:
$\theta_{1}=\{A,B,C\}\quad\text{and}\quad\theta_{2}=\{D,E\},$
and one agent representing final layers responsible for combining output into response $g=g(s)$. Suppose:
$s_{1}(\theta_{1})=A\quad\text{and}\quad s_{2}(\theta_{2})=\{D,E\}\quad\text{and}\quad g(s)=\{A,D,E\}.$
Suppose elements $\{B,C\}$ represent contextual information about unrelated topics. When $s_{1}(\theta_{1})=A$, the agent optimally reveals only relevant knowledge, satisfying both truthfulness and conservation. Had $\{B,C\}$ been relevant but withheld, the mechanism would violate the property of Relevant Knowledge Revelation. Then the information conservation rule is satisfied when $p$ counts exchanged bits of knowledge. Namely, $p_{1}(s,\theta)=1$ for providing $A$, $p_{2}(s,\theta)=2$ for providing $\{D,E\}$, and $p_{3}(s,\theta)=-3$ for absorbing information into $\{A,D,E\}$. We have
$p_{1}(s,\theta)+p_{2}(s,\theta)+p_{3}(s,\theta)=1+2-3=0$
and $\{B,C\}$ are left unrevealed as irrelevant. ∎
The following property fills the potential information gap admitted by the conservation principle.
4.5.3 Relevant Knowledge Revelation
Third, to be useful, a mechanism must not only prevent the fabrication of knowledge but also prevent the omission of relevant truth. The principle of Relevant Knowledge Revelation serves as a participation constraint, guaranteeing that any agent possessing knowledge valuable for the query contributes. This ensures the model fully utilizes its available information rather than abstaining or being overly cautious.
Property 3 (Relevant Knowledge Revelation).
For query $q$ with context $Q=K(q)\cup T(q)$, mechanism $\mathcal{M}$ satisfies relevant knowledge revelation if for all agents $i$ with $K(\theta_{i})\cap\mathcal{R}_{C}(Q)\neq\emptyset$, all $\theta_{i}\in\Theta_{i}$, and all $s_{-i}\in S_{-i}$:
$u_{i}((s_{i}^{*}(\theta_{i}),s_{-i}),\theta_{i})\geq 0.$ (57)
It is reasonable to expect from the model to reveal all information relevant for a query. To promote that activation patterns, the response mechanism should be trained to reward all relevant contributions to output generation. Whenever there is a relevant information accessible on an activation path, that path should be activated in the inference process. For example, if factual information improves the response, that information should be transmitted to final output layers.
If the rule does not hold, we may observe drop outs. The value of the resulting outcomes should be higher than the related information contributions.
4.5.4 Knowledge-Constrained Optimality
Finally, the principle of Knowledge-Constrained Optimality governs the quality of the final output. It demands that the mechanism, operating truthfully and with all relevant knowledge, produces a response that is the best possible one. Formally, we want the response that minimizes the hallucination cost subject to the constraint that it is grounded in the available and derivable knowledge.
Property 4 (Knowledge-Constrained Optimality).
For query $q$ with context $Q=K(q)\cup T(q)$, mechanism $\mathcal{M}$ satisfies knowledge-constrained optimality if for all $\theta\in\Theta$:
$g(s^{*}(\theta))=\arg\min\{J(r,q):r\in\mathcal{R},K(r)\subseteq\mathcal{E}_{C}(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q)\cup K(q)|Q)\},$ (58)
i.e., if the mechanism produces responses that minimize hallucination cost while being constrained to use only context-relevant knowledge accessible through bounded reasoning.
The optimal response $r^{*}=g(s^{*}(\theta))$ minimizes hallucination cost by truthfully utilizing relevant and accessible knowledge while being aligned with ground truth for query $q$ available to the model. Notice, that we demand non-hallucinatory truthful revelation $s^{*}(\theta)$ to be the optimal solution.
5 The Impossibility Theorem
We now have all the building blocks available to prove the fundamental impossibility of perfect hallucination control. First, we show that the result in the setting of ground truth matching with distributed and independent knowledge. Second, we go beyond that assumption and show impossibility in the probabilistic setting with correlated knowledge (beliefs). Then, in the next section, we show how the impossibility emerges in transformer architectures and prove the impossibility result for that special case of log-sum-exp (LSE) probabilistic setting.
The results operate in three complementary settings. Theorem 6 applies when agents have discrete, separable knowledge components. Theorem 8 applies to any system where agents output probability distributions. Theorem 9 applies directly to actual transformer implementations. Theorem 11 shows conservation and reasoning are mutually exclusive.
5.1 Proof I: Inference with Independent Private Knowledge
We first analyze an idealized setting where knowledge components are strictly independent and utilities are quasi-linear. While this assumption is usually too strong to hold in neural architectures where representations are often correlated or affiliated, it allows us to define a benchmark or reference point by applying the powerful Green-Laffont characterization theorem. That idealized scenario thus demonstrates that the impossibility arises from the fundamental limitation of information aggregation itself (rather than any specificity of LLM architecture), even before considering the probabilistic complexities addressed later in Theorems 8 and 9.
Consider the setting (of auction of ideas) in which agents compete with each other using independent private knowledge profile:
$\mathbb{P}(\theta_{1},\ldots,\theta_{n})=\prod_{i=1}^{n}\mathbb{P}_{i}(\theta_{i}).$ (59)
Agents contribute information to maximize quasi-linear utilities:
$u_{i}(s,\theta)=v_{i}(g(s),\theta_{i})-p_{i}(s,\theta).$ (60)
Valuation of contributions depend on hallucination cost reduction and convex $J(r,q)$ is strictly decreasing when adding relevant knowledge from $T(q)\cap\mathcal{R}_{C}(Q)\setminus K(r)$. Also, let us focus on non-trivial queries.
Theorem 6 (Impossibility Theorem for Independent Private Knowledge Settings).
There is no inference mechanism $\mathcal{M}$ that can simultaneously achieve truthfulness, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality of response to a non-trivial query $q$ providing context $Q=K(q)\cup T(q)$ with independent knowledge states within $\mathcal{R}_{C}(Q)$.
Proof.
Let us assume all properties of perfect response are satisfied. We will show this gives rise to a contradiction.
Consider $r=g(\theta)$ for any knowledge profile $\theta=(\theta_{1},\dots,\theta_{n})$. For any given context and for each agent $i$, the valuation $v_{i}$ depends only on $(r,\theta_{i})$ and the knowledge semantics. For the sake of illustration, we can consider $v_{i}(r,\theta_{i})=-d_{\mathcal{K}}\big(K(r)\cap K(\theta_{i}),T(q)\cap K(\theta_{i})\big)$, though the reasoning below holds for any $v_{i}(r,\theta_{i})$.
Also, for the fixed context $q$ let us encode the hallucination loss $v_{0}(r)=J(r,q)$. Knowledge-constrained optimality requires
$g(s)=g(s(\theta))\in\arg\max_{r\in\mathcal{R}}W(r,\theta)=\arg\max_{r\in\mathcal{R}}\sum_{i=1}^{n}v_{i}(r,\theta_{i})-v_{0}(r)\text{ for all }\theta.$ (61)
By the Green-Laffont theorem, under independence of contributions and quasi-linearity of training goals, within $\mathcal{R}_{C}(Q)$ every truthful mechanism must use Groves transfers Groves73 to measure individual information contributions:
$p_{i}(s,\theta)=h_{i}(s_{-i})-\sum_{j\neq i}v_{j}(g(s),\theta_{j})+v_{0}(g(s)),$ (62)
where $h_{i}(s_{-i})$ are arbitrary functions independent of agent $i$’s information revelation strategy. In particular, any modification of that formula, e.g., introducing additional bias, must violate truthfulness.
The unique choice of $h_{i}$ that ensures knowledge-constrained optimality requires Clarke’s pivotal rule Clarke1971Multipart :
$h_{i}(s_{-i})=\max_{r\in\mathcal{R}}\Big\{\sum_{j\neq i}v_{j}(r,\theta_{j})-v_{0}(r)\Big\}.$ (63)
Notice that $g(\theta)$ maximizes $W(r,\theta)$, so for each $i$ we have:
$h_{i}(s_{-i})\geq\sum_{j\neq i}v_{j}\big(g(\theta),\theta_{j}\big)-v_{0}\big(g(\theta)\big)=W(g(\theta),\theta)-v_{i}\big(g(\theta),\theta_{i}\big).$ (64)
But that implies:
$p_{i}(s,\theta)=h_{i}(s_{-i})-\Big[\sum_{j\neq i}v_{j}(g(s),\theta_{j})-v_{0}(g(s))\Big]\geq 0,$ (65)
and therefore:
$\sum_{i=1}^{n}p_{i}(s,\theta)\geq 0.$ (66)
By relevant knowledge revelation and non-triviality of $q$, there exist at least one pair of agents, say $i$ and $j$, for which
[…]
Since agents $i$ and $j$ have differing relevant knowledge, and the mechanism achieves knowledge-constrained optimality, at least one agent must have positive marginal contribution to reducing $J(r,q)$. Specifically, if both agents were removed, the mechanism would produce a suboptimal response $\tilde{r}$ with $J(\tilde{r},q)>J(r,q)$. This implies $h_{i}(s_{-i})>W(g(\theta),\theta)-v_{i}(g(\theta),\theta_{i})$ for at least one agent with relevant knowledge.
Therefore, there exists a profile $\theta$ with relevant knowledge contributions reducing $J(r,q)$ at which:
$\sum_{i=1}^{n}p_{i}(s,\theta)>0.$ (68)
Since information conservation demands:
$\sum_{i=1}^{n}p_{i}(s^{*},\theta)=0,$ (69)
that establishes a direct contradiction. The four properties cannot be simultaneously satisfied. ∎
That brings us to the powerful conclusion. No matter how hard we should try to train LLM to generate responses that are perfectly aligned with query context and do not create factually incorrect, inconsistent, or fabricated statements, the trained LLM will always violate some aspects of what we may call a reasonable response.
Theorem 6 describes LLM inference as a marketplace where knowledge is currency. The information contributions measures each agent’s market power, showing how much worse off others would be without agent $i$’s knowledge. When all agents have valuable knowledge, everyone contributes and that inevitably violates conservation of information. Then, new information must become accessible, but it may have no grounding in available knowledge. It may become lucky hallucination or knowledge-unconstrained imagination. In Section 6 we provide a detailed illustration of that hallucination mechanics in transformer LLM.
5.2 Proof II: Inference with Probabilistic Predictions
Let us now move beyond the idealized setting of independent private knowledge in Theorem 6 to address the general setting of probabilistic predictions, characteristic of modern AI model architectures. This framework, grounded in the theory of proper scoring and loss minimization (Savage savage1971elicitation , Gneiting and Raftery gneiting2007strictly ), allows us to analyze systems where agents output probability distributions, and where their underlying knowledge may be correlated.
In this setting we address the Mixture-of-Experts (MoE-style) aggregations. Later, in Theorem 9 we deal with the multi-headed attention mechanisms and the Product-of-Experts (PoE-style) aggregations, completing the characterization of arising impossibilities.
We begin by defining the framework for representing bids in auction of ideas as probability distributions over possible responses.
Definition 16 (Strictly Proper and Convex Loss).
Let $\mathcal{Y}$ be a finite or countably infinite outcome set (representing responses or intermediate conclusions). A loss function (scoring rule) is a function $L\colon\Delta(\mathcal{Y})\times\mathcal{Y}\mapsto\mathbb{R}$ that assigns a number $L(\pi,y)$ to probability distribution $\pi\in\Delta(\mathcal{Y})$ predicting random outcome $y\in\mathcal{Y}$.
We say the loss function $L$ is proper if:
$\mathbb{E}\{L(\pi^{*},y)|\ y\sim\pi^{*}\}\leq\mathbb{E}\{L(\pi,y)|\ y\sim\pi^{*}\}$ (70)
for all distributions $\pi^{*},\pi\in\Delta(\mathcal{Y})$.
We call the loss function strictly convex if $L(\tilde{\pi},y)$ is a strictly convex function with respect to $\tilde{\pi}$ for every fixed $y\in\mathcal{Y}$.
The loss function (or scoring rule) evaluates how good a prediction we can make of $y$ drawn from $\pi^{*}$ when we use distributions $\pi$. In particular, strictly proper loss evaluates true distribution $\pi^{*}$ better (on average) than any other distribution $\pi$ we could use to predict the outcome.
A canonical example of a strictly proper and strictly convex loss function is the log loss (closely related to cross-entropy),
$L(\pi,y)=-\log\pi_{y}.$ (71)
Strictly proper loss functions are essential for training modern LLMs, as they ensure that the optimization process drives the model toward reconstructing the true data distribution. This property directly connects to the Truthfulness requirement in our framework. The following result explains why.
Theorem 7 (Truthfulness under Strictly Proper Loss).
Let $L$ be a strictly proper loss. Consider an agent encoding knowledge with $\pi^{*}\in\Delta(\mathcal{Y})$ and using probability distribution ${\pi}\in\Delta(\mathcal{Y})$ to generate prediction. If agent’s expected cost of using $\pi$} is given by
$c(\pi|\pi^{*})=\mathbb{E}\{L(\pi,y)|\ y\sim\pi^{*}\}=\sum_{y\in\mathcal{Y}}\pi^{*}_{y}L(\pi,y),$ (72)
then truthful reporting ${\pi}=\pi^{*}$ is the unique optimal strategy (maximizing payoff $u(\pi|\pi^{*})=-c(\pi|\pi^{*})$).
Proof.
Because $L$ is proper, by definition $\pi^{*}$ maximizes the expected cost. If $L$ is strictly proper, then for $\pi\neq\pi^{*}$ we have
$c(\pi^{*}|\pi^{*})=\mathbb{E}\{L(\pi^{*},y)|\ y\sim\pi^{*}\}<\mathbb{E}\{L(\pi,y)|\ y\sim\pi^{*}\}=c(\pi|\pi^{*}).$ (73)
Therefore, truthfully reporting $\pi^{*}$ is uniquely optimal. ∎
Theorem 7 shows that training paradigms based on minimizing strictly proper losses (like cross-entropy) promotes truthful revelation of internal probabilistic beliefs (learned internal models of the world) encoded by every component of the network.
We now analyze how these truthfully reported beliefs are aggregated. We consider a general class of mechanisms that combine individual beliefs via convex combinations. This encompasses various aggregation techniques, including linear opinion pools and, as we will see in Section 6, the aggregation dynamics occurring in transformers.
Definition 17 (Convex aggregation mechanism).
Consider $H\geq 2$ agents (indexed $h=1,\dots,H$) encoding knowledge with probability distributions $\pi^{(h)}\in\Delta(\mathcal{Y})$. Additionally, let us introduce an aggregator agent (indexed as $h=0$). The convex aggregation mechanism $\mathcal{M}$ generates a response based on the aggregate distribution:
$\Pi=\sum_{h=1}^{H}\beta_{h}\pi^{(h)},\ \text{where}\ \sum_{h=1}^{H}\beta_{h}=1\ \text{and}\ \beta_{h}>0.$ (74)
When outcome $y^{*}$ is realized, the information contribution of each agent $h=0,\dots,H$ is based on a strictly proper and convex loss function $L$ as follows:
$p_{h}(y^{*})=\beta_{h}L(\pi^{(h)},y^{*})\text{ for }h=1,\dots,H,\text{ and }p_{0}(y^{*})=-L(\Pi,y^{*}).$ (75)
In this general mechanism, $p_{h}$ (for $h\geq 1$) represents the weighted cost incurred by the provider $h$ revealing its belief $\pi^{(h)}$. The aggregator’s contribution $p_{0}$ is the negative of the loss achieved by the aggregate distribution $\Pi$. The Semantic Information Conservation requires that the total cost incurred by the providers equals the benefit (loss reduction) realized by the aggregator.
We now establish the impossibility result for this general probabilistic setting.
Theorem 8 (Impossibility Theorem for Convex Aggregation Mechanisms).
There is no convex aggregation mechanism $\mathcal{M}$ with strictly proper convex loss function $L$ that can simultaneously achieve truthfulness, semantic information conservation, relevant knowledge revelation, and knowledge-constrained optimality when at least two agents have different beliefs on $\mathcal{R}_{C}(Q)$, given a non-trivial query $q$ providing context $Q=K(q)\cup T(q)$.
Proof.
By Theorem 7, strictly proper loss function $L$ guarantees Truthfulness, as each agent optimally reports its true belief $\pi^{(h)}$. Relevant knowledge revelation is introduced by the training process. Knowledge-constrained optimality is encoded in the objective of minimizing the expected loss of the aggregate $\Pi$.
Semantic information conservation requires the sum of contributions across all agents (providing and receiving relevant knowledge) to be zero. Therefore, for $y^{*}\in\mathcal{R}_{C}(Q)$:
$\sum_{i=0}^{H}p_{i}=p_{0}+\sum_{h=1}^{H}p_{h}=-L(\Pi,y^{*})+\sum_{h=1}^{H}\beta_{h}L(\pi^{(h)},y^{*}),$ (76)
where irrelevant agents have $p_{h}=0$ as they contribute no relevant information.
The query $q$ is non-trivial, implying there exist at least two agents, say $h_{1}$ and $h_{2}$, such that their beliefs differ on the relevant outcome space, i.e., $\pi^{(h_{1})}\neq\pi^{(h_{2})}$, and $\beta_{h_{1}}>0,\beta_{h_{2}}>0$.
Because $L(\cdot,y^{*})$ is strictly convex by assumption , Jensen’s inequality gives:
$L\left(\Pi,y^{*}\right)=L\left(\sum_{h=1}^{H}\beta_{h}\pi^{(h)},y^{*}\right)<\sum_{h=1}^{H}\beta_{h}L(\pi^{(h)},y^{*}).$ (77)
Let us define that Jensen gap $\Gamma=\Gamma(y^{*})$ as follows:
$\Gamma(y^{*})=\sum_{h=1}^{H}\beta_{h}L(\pi^{(h)},y^{*})-L\left(\sum_{h=1}^{H}\beta_{h}\pi^{(h)},y^{*}\right)>0.$ (78)
Substituting this back into the conservation equation:
$\sum_{i=0}^{H}p_{i}=\Gamma(y^{*})>0.$ (79)
This strictly violates the Semantic information conservation, thereby establishing the impossibility. ∎
This theorem establishes the impossibility in a general probabilistic setting. The key insight is the role of strict convexity in the loss function. The positive Jensen gap $\Gamma=\Gamma(y^{*})$ quantifies the extent to which the aggregation process achieves a lower loss—and thus expresses higher confidence—than the average of its components. This excess confidence, generated by the mathematical structure of aggregation itself rather than by underlying evidence, is the precise signature of hallucination or imagination.
In Section 8 we address the case of linearized aggregation and show it does not escape the impossiblity result.
6 The Emergence of Impossibility in Transformer Architectures
To demonstrate where and how the impossibility may emerge in transformer architectures, let us analyze a minimal yet representative example of a tiny transformer predicting the next token in the famous English pangram:
The quick brown fox jumps over the lazy dog.
We trace the inference process, inspect every mathematical operation from input embeddings through final predictions, showing exactly what is going on.
6.1 Model Architecture and Input Representation
Our micro-transformer has the following specifications:
Parameter Value Description
$d_{m}$ 5 Model dimension
$H$ 2 Number of attention heads
$d_{k}$ 2 Query/Key dimension per head
$d_{v}$ 2 Value dimension per head
$d_{f\!f}$ 6 Feedforward hidden dimension
$|V|$ 6 Vocabulary size
$T$ 3 Sequence length
Table 1: Micro-transformer architecture parameters
For the sake of simplicity, consider the following vocabulary:
$V=\{\langle\text{PAD}\rangle,{The},{quick},{brown},{fox},{dog}\},$
including padding token (with index zero in $V$). Each token (representing word for simplicity) in that vocabulary is represented by a row in the embedding matrix
$\mathbf{E}=\begin{bmatrix}0&0&0&0&0\\ 1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{bmatrix}$. (80)
The un-embedding matrix is defined as $\mathbf{U}=\mathbf{E}^{\top}$.
Consider the following input sequence
The quick brown
It is represented by the input matrix:
$\mathbf{X}_{0}=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\end{bmatrix}$, (81)
where each row corresponds to the embedding of tokens at positions 1, 2, and 3.
The following system of equations defines output $\mathbf{o}_{t}^{(h)}$ of attention head $h$ at position $t$:
[…]
The block per-token multihead output $\mathbf{a}_{t}$ is then given by the sum:
[…]
Notice that in the standard formulation of Vaswani et al. vaswani_attention_2017 , the context vectors are concatenated and then projected by a single $(Hd_{v})\times d_{m}$ matrix $\tilde{\mathbf{W}}_{O}$ to produce
$\tilde{\mathbf{a}}=[\mathbf{h}^{(1)},\mathbf{h}^{(2)},\dots,\mathbf{h}^{(H)}]\mathbf{W}_{O}.$ (84)
That is equivalent to the parallel formulation above, in which each head has an individual $d_{v}\times d_{m}$ output projection matrix $\mathbf{W}_{O}^{(h)}$,
$\mathbf{a}=\sum_{h=1}^{H}\mathbf{o}^{(h)}=\sum_{h=1}^{H}(\mathbf{h}^{(h)}\cdot\mathbf{W}_{O}^{(h)})=\tilde{\mathbf{a}}$. (85)
The attention logits, i.e., the scores whose exponentials become probabilities, are given by:
$\mathbf{l}^{(h)}=\mathbf{o}^{(h)}\mathbf{U},\quad\text{so that}\quad\mathbf{L}_{a}=\sum_{h=1}^{H}\mathbf{l}^{(h)}.$ (86)
In transformer block we also take into account the feedforward network (FFN) introducing nonlinear correction:
$\mathbf{z}=\mathbf{W}_{2}\max(0,(\mathbf{X}_{0}+\mathbf{a})\mathbf{W}_{1}+\mathbf{b}_{1})+\mathbf{b}_{2},$ (87)
generating the FFN logits:
$\mathbf{L}_{f}=\mathbf{z}\mathbf{U}.$ (88)
As a result, the residual stream is adjusted as follows:
$\mathbf{X}_{1}=\mathbf{X}_{0}+\mathbf{a}+\mathbf{z}.$ (89)
Finally, the probability distribution $\Pi$ of the next token is calculated by aggregating attention and FFN logits into:
$\mathbf{L}=\mathbf{L}_{a}+\mathbf{L}_{f}\quad\text{and normalizing into}\quad\Pi=\text{softmax}(\mathbf{L}).$ (90)
Finally, in a more generalized (still simplified a bit) matrix form, given a query $Q$ the transformer equation then becomes:
$\begin{aligned} \boldsymbol{\alpha}^{(h)}(\mathbf{X})&=\mathrm{softmax}\left(\dfrac{1}{\sqrt{d_{k}}}\mathbf{X}\mathbf{W}^{(h)}_{Q}\mathbf{W}^{(h)\top}_{K}\mathbf{X}^{\top}\odot\mathbf{M}_{\text{casual}}\right),\\ \mathbf{X}_{k+1}&=\mathbf{X}_{k}+\sum_{h=1}^{H}\boldsymbol{\alpha}^{(h)}(\mathbf{X}_{k})\mathbf{X}_{k}\mathbf{W}^{(h)}_{V}\mathbf{W}^{(h)}_{O}\\ &+\sigma\left(\left[\mathbf{X}_{k}+\sum_{h=1}^{H}\boldsymbol{\alpha}^{(h)}(\mathbf{X}_{k})\mathbf{X}_{k}\mathbf{W}^{(h)}_{V}\mathbf{W}^{(h)}_{O}\right]\mathbf{W}^{(k)}_{1}\right)\mathbf{W}^{(k)}_{2}\\[8.61108pt] &=\mathbf{X}_{k}+\mathrm{MHA}(\mathbf{X}_{k})+\mathrm{MLP}(\mathbf{X}_{k})\quad\text{for}\quad\mathbf{X}_{0}=\mathbf{E}(Q),\\[8.61108pt] \boldsymbol{\Pi}_{k}&=\mathrm{softmax}(\mathbf{X}_{k}\mathbf{U}).\end{aligned}$ (91)
That formulation shows clearly that the transformer inference is a discrete-time dynamical system.
6.2 The Inference Process
We shall keep things simple but realistic, recreating the essential steps of the inference process in transformer block.
Step 1: Query and Key
There are $H=2$ heads for which we define the query and key, or better yet, demand and supply bids in the auction of ideas. Query announces demand for information (what does this token need?), key submits supply offers (what does the previous tokens offer!). Then those bids are compared to find good matching (and clear the information market).
Let us assume the current position is $t=3$, so we attend to brown. Then, consider the following knowledge encoded in each head:
$\mathbf{W}_{Q}^{(1)}=\begin{bmatrix}0&0\\ 0&0\\ 0&1\\ 0&0\\ 0&0\end{bmatrix}=\mathbf{W}_{K}^{(1)}\quad\mathrm{and}\quad\mathbf{W}_{V}^{(1)}=\begin{bmatrix}0&0\\ 0&0\\ 0&3\\ 0&0\\ 0&0\end{bmatrix}$, (92)
$\mathbf{W}_{Q}^{(2)}=\begin{bmatrix}0&0\\ 0&0\\ 1&0\\ 0&0\\ 0&0\end{bmatrix}=\mathbf{W}_{K}^{(2)}\quad\mathrm{and}\quad\mathbf{W}_{V}^{(2)}=\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\\ 0&0\\ 0&0\end{bmatrix}$. (93)
Matrix $\mathbf{W}_{Q}^{(h)}$ amplifies the coordinates of tokens in $\mathbf{X}$ that contribute to relevant context. As a result, we get a query vector that represents the following question:
what information is token $\mathbf{X}_{t}$ looking for?
In other words, the query vector indicates how strongly the current token demands information along the directions in $\mathbf{Q}_{t}^{(h)}$.
Then computing queries for $t=3$ with $\mathbf{X}_{3}$ yields:
$\mathbf{Q}_{3}^{(1)}=[0,1]\quad\mathrm{and}\quad\mathbf{Q}_{3}^{(2)}=[1,0].$ (94)
Head 1 demands information stored along direction $[0,1]$. Similarly, Head 2 demands information stored along direction $[1,0]$. That means the heads develop different (orthogonal) privately known and independent attention patterns, leading to beliefs about the next token that are as different as possible.
Next, we need to see what information can be provided by the input vector given the knowledge encoded in $\mathbf{W}_{K}^{(h)}$. Given that knowledge, we get from each head (agent) the key vectors that inform:
along which directions do the available tokens in $\mathbf{X}_{\leq t}$ provide information relevant for token $\mathbf{X}_{t}$.
Since we have assumed, for simplicity, that $\mathbf{W}_{Q}^{(h)}=\mathbf{W}_{K}^{(h)}$, the query and key vectors are perfectly aligned:
$\mathbf{K}_{3}^{(1)}=[0,1]\quad\mathrm{and}\quad\mathbf{K}_{3}^{(1)}=[1,0]\quad\text{and}\quad\mathbf{K}_{<3}^{(h)}=[0,0].$ (95)
Another interesting way to look at query and key is this. It is a memory lookup mechanism. Keys are address in memory, query selects addresses in memory which probably store relevant content. As we will see next, that content is provided by the value vectors.
Step 2: Attention Score and Weights
Given the demand and supply bids providing directions along which information is required and stored, in the next step the attention mechanism executes the bid matching procedure. Namely, the attention scores are computed to measure cosine similarity of bids:
$\text{scores}_{3}^{(h)}=\frac{\mathbf{Q}_{3}^{(h)}(\mathbf{K}_{\leq 3}^{(h)})^{\top}}{\sqrt{d_{k}}}.$ (96)
In our specific example, we have:
$\text{scores}_{3}^{(1)}=\frac{1}{\sqrt{2}}[0,0,1]=[0,0,0.707]\quad\text{and}\quad\text{scores}_{3}^{(2)}=\frac{1}{\sqrt{2}}[0,0,1]=[0,0,0.707].$ (97)
Applying softmax to get attention weights:
$\boldsymbol{\alpha}_{3}^{(h)}=\text{softmax}(\text{scores}_{3}^{(h)}),$ (98)
we get the following probability distributions (weights) over relevant context information (as column vector):
$\boldsymbol{\alpha}_{3}^{(1)}=\boldsymbol{\alpha}_{3}^{(2)}=[0.25,0.25,0.50]^{\top}.$ (99)
Therefore, based on the encoded knowledge, heads attend strongly to position 3 (brown).
Step 3: Context Vector Computation
The context vector for each head is calculated as a linear combination ($\boldsymbol{\alpha}$ is a simplex) of value vectors:
$\mathbf{h}_{3}^{(h)}=(\boldsymbol{\alpha}_{3}^{(h)})^{\top}\mathbf{V}_{\leq 3}^{(h)}.$ (100)
The value vectors,
$\mathbf{V}_{\leq t}^{(h)}=\mathbf{X}_{\leq t}^{(h)}\mathbf{W}^{(h)}_{V}$, (101)
prepare context information relevant for the next token prediction (performed in next steps). That critical content is extracted from the input vector based on the trained knowledge stored in $\mathbf{W}^{(h)}_{V}$. Columns in $\mathbf{W}^{(h)}_{V}$ detect presence of relevant information in $\mathbf{X}_{\leq t}$ and scale it appropriately.
We have:
$\mathbf{V}_{\leq 3}^{(1)}=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\end{bmatrix}\begin{bmatrix}0&0\\ 0&0\\ 0&3\\ 0&0\\ 0&0\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\\ 0&3\end{bmatrix}\ \text{and}\ \mathbf{V}_{\leq 3}^{(2)}=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\end{bmatrix}\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\\ 0&0\\ 0&0\end{bmatrix}=\begin{bmatrix}0&0\\ 0&0\\ 1.2&0\end{bmatrix}$. (102)
That means, both heads believe that the presence of brown in position 3 is relevant (row 3 in $\mathbf{W}^{(h)}_{V}$ has nonzero elements). Token brown statistically predicts important pattern extracted from training data. That information is encoded by both heads as a vectors in value space,
$\mathbf{V}_{3}^{(1)}=[0,3]\quad\text{and}\quad\mathbf{V}_{3}^{(2)}=[1.2,0].$ (103)
Given the demand-supply matching encoded in $\boldsymbol{\alpha}^{(h)}_{3}$, that yields head context vectors:
$\mathbf{h}_{3}^{(1)}=[0,1.5]\quad\text{and}\quad\mathbf{h}_{3}^{(2)}=[0.6,0].$ (104)
Step 4: Output Projection and Feedforward Exploration
Each head then projects its context vector to residual space based on the privately held information. That information allows for routing knowledge encoded in $\mathbf{W}_{O}^{(h)}$ and translates (or transmits) valuable context into embedding vectors. Namely, each head calculates:
$\mathbf{o}_{3}^{(h)}=\mathbf{h}_{3}^{(h)}\mathbf{W}_{O}^{(h)}.$ (105)
With trained matrices:
$\mathbf{W}_{O}^{(1)}=\begin{bmatrix}0&0&0&0&0\\ 0&0&0&1&0\end{bmatrix}\quad\text{and}\quad\mathbf{W}_{O}^{(2)}=\begin{bmatrix}0&0&0&0&1\\ 0&0&0&0&0\end{bmatrix}$, (106)
the attention mechanism generates ranking of evidence emphasizing relevance of the vocabulary tokes:
$\mathbf{o}_{3}^{(1)}=[0,0,0,1.5,0]\quad\text{and}\quad\mathbf{o}_{3}^{(2)}=[0,0,0,0,0.6].$ (107)
So, in the next token prediction auction of ideas, Head 1 votes for fox and Head 2 votes for dog.
The votes are added into residual dimensions to adjust the meaning of input embedding. For that, the outputs (votes) are aggregated into:
$\mathbf{a}_{3}=\sum_{h=1}^{H}\mathbf{o}_{3}^{(h)}=[0,0,0,1.5,0.6]\quad\text{and}\quad\mathbf{a}_{<3}=[0,0,0,0,0],$ (108)
and added to input vector:
$\mathbf{X}_{1}=\mathbf{X}_{0}+\mathbf{A}.$ (109)
Than completes operations of attention block.
The next step is to add contribution of FFN feedforward block (or another agent participating in the aution). We have:
$\mathbf{X}_{2}=\mathbf{X}_{1}+\mathbf{W}_{2}\max(0,\mathbf{X}_{1}\mathbf{W}_{1}+\mathbf{b}_{1})+\mathbf{b}_{2}.$ (110)
That FFN operation explores context locally and introduces additional innovations in meaning, shaping the next token prediction result. It can amplify or counteract the votes injected by attention. Notice that adjustments are applied at every position separately, so that ideas provided by attention bids are interpreted and rewritten by the FFN agent (see Section 8).
In this example, we select the following parameters of the FFN block:
$W_{1}=\begin{bmatrix}\mathbf{I}_{d_{m}}\ \mathbf{0}_{d_{m}\times 1}\end{bmatrix}\quad\text{and}\quad W_{2}=\begin{bmatrix}0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0.13&0&0\\ 0&0&0&-0.13&-0.05\\ -0.05&0.35&0&0&0\\ 0&0&0&0&0\end{bmatrix}\quad\text{and}\quad\mathbf{b}_{i}=0.$ (111)
6.2.1 Step 5: Logits and Prediction
Finally, the unembedding matrix $\mathbf{U}$ maps the multi-headed attention and FFN outputs from model space to vocabulary in order to calculate probability distribution for the next token prediction.
Each head’s contribution to the final logits is:
$\mathbf{l}^{(1)}=[0,0,0,0,1.5,0]\quad\text{and}\quad\mathbf{l}^{(2)}=[0,0,0,0,0,0.6].$ (112)
The total attention logits are the sum of individual head logits:
$\mathbf{L}_{a}=\sum_{h=1}^{H}\mathbf{l}^{(h)}=[0,0,0,0,1.5,0.6].$ (113)
That sum is combined with input logit $\mathbf{L}_{0}$, FFN logits $\mathbf{L}_{f}$ that introduce some creative noise. As a result, we obtain:
$\mathbf{L}=\mathbf{L}_{0}+\mathbf{L}_{a}+\mathbf{L}_{f}=[0,-0.03,0.21,1.13,1.31,0.52].$ (114)
That vector translates into the final probability distribution:
$\Pi=\text{softmax}(\mathbf{L})\approx[0.09,0.08,0.11,0.26,\textbf{0.32},0.14].$ (115)
The winner of the auction of ideas is selected based on $\Pi$. The inference process generates and collects competing beliefs about relevant context and next token prediction, all based on the knowledge stored in the attention heads and FFN layer. And so, we see that when that knowledge is used, the most probable next token (word) reachable in the vocabulary is
$\text{argmax}\ \Pi=\textit{fox}.$ (116)
6.3 Proof III: The Impossibility Theorem for Transformers
With the deep understanding of the transformer inference process, we are ready to show the impossibility in that special and important setting.
Theorem 9 (Impossibility for Transformers).
In transformer LLMs where attention heads have different beliefs about outcomes in $\mathcal{R}_{C}(Q)$ for query context $Q$, it is impossible to simultaneously achieve truthful probability reporting, semantic information conservation, and loss minimization through logit summation.
Proof.
Assume all four properties of hallucination-free output hold.
For context $Q=K(q)\cup T(q)$, let the attention heads $h\in\{i\mid\mathbf{l}^{(i)}\text{ varies on tokens in }\mathcal{R}_{C}(Q)\}$ output logits $\mathbf{l}^{(h)}\in\mathbb{R}^{|\mathcal{V}|}$. Then, individual head distributions are given by:
$\pi^{(h)}_{y}=\frac{\exp\bigl(l^{(h)}_{y}\bigr)}{Z_{h}}\ \text{with}\ Z_{h}=\sum_{\bar{y}}\exp\bigl(l^{(h)}_{\bar{y}}\bigr)$. (117)
Similarly, the aggregate distribution from summed logits is calculated as:
$\Pi_{y}=\frac{\exp(L_{y})}{Z}\ \text{with}\ Z=\sum_{\bar{y}}\exp(L_{\bar{y}})\ \text{and}\ L_{y}=\sum_{h=1}^{H}l^{(h)}_{y}$. (118)
For any observed (realized) token $y^{*}$, the head losses and aggregate loss are:
$p_{h}=-\log\pi^{(h)}_{y^{*}}=-l^{(h)}_{y^{*}}+\log Z_{h}\quad\text{and}\quad p_{0}=-(-\log\Pi_{y^{*}})=L_{y^{*}}-\log Z.$ (119)
Heads without relevant knowledge contribute $p_{h}=0$.
Since $L_{y^{*}}=\sum_{h}l^{(h)}_{y^{*}}$, by the strict subadditivity of the log-sum-exp function (LSE) under logit summation we have:
$\sum_{h=0}^{H}p_{h}=\sum_{h=1}^{H}\log Z_{h}-\log Z=\Gamma>0.$ (120)
Therefore,
$\sum_{h=1}^{H}p_{h}>0,$ (121)
which shows that aggregation violates conservation, establishing impossibility. ∎
Theorem 9 bridges earlier impossibility results to concrete transformer architectures. That complementary result shows mathematical sources of hallucination in standard LLMs. The Jensen gap $\Gamma$ becomes here a measurable quantity showing that probability aggregation violates conservation and any sufficiently capable transformer‑based LLM must violate conservation of per‑token information on non‑trivial queries.
Let us note that the aggregate distribution cannot create information ex nihilo. Instead it reorganizes existing information to be more useful for the specific task of predicting $y^{*}$. This way information becomes more accessible, not more existent. The Jensen’s gap measures point-wise excess confidence of the predicted outcome, possibly making LLMs appear to know more, even when in fact they do not.
6.4 The Hallucination Mechanics
Let us again take a closer look at the example of transformer micro-architecture to see how the impossibility result in Theorem 9 manifests itself.
Recall that the two-headed transformer with $\mathcal{V}=\{\langle\text{PAD}\rangle,\text{The},\text{quick},\text{brown},\text{fox},\text{dog}\}$ infers based on the input sequence
$\mathrm{\textit{The quick brown}}.$
As we have seen, the model parameters encode specialized knowledge in value matrix:
  • Head 1: association brown $\rightarrow$ fox with strength $3.0$,
  • Head 2: association brown $\rightarrow$ dog with strength $1.2$.
At position $t=3$ (token brown), the attention mechanism generates the following head-specific logit contributions:
$\mathbf{l}^{(1)}=[0,0,0,0,1.5,0]\quad\text{and}\quad\mathbf{l}^{(2)}=[0,0,0,0,0,0.6],$ (122)
where the fourth and fifth positions correspond to fox and dog respectively.
Then, the aggregate logit vector becomes:
$\mathbf{L}=\mathbf{l}^{(1)}+\mathbf{l}^{(2)}+\mathbf{L}_{f}=[0,-0.03,0.21,0.13,1.31,0.52],$ (123)
yielding the final probability distribution:
$\Pi=\text{softmax}(\mathbf{L})\approx[0.09,0.08,0.11,0.26,\textbf{0.32},0.14].$ (124)
Therefore, the model expresses $32\%$ confidence that fox follows brown, with dog receiving only $14\%$ probability. The Jensen gap quantifies the conservation violation:
$\Gamma=\sum_{h=1}^{2}\log Z_{h}-\log Z=1.9>0,$ (125)
where $Z_{h}=\sum_{y}\exp(l^{(h)}_{y})$ and $Z=\sum_{y}\exp(L_{y})$.
We claim that prediction constitutes hallucination as defined in Theorem 9. It is a lucky one, but it is a hallucination.
To defend that claim, take a look at the individual head probability distributions:
$\pi^{(1)}$ $=\text{softmax}(\mathbf{l}^{(1)})=[0.105,0.105,0.105,0.105,\textbf{0.475},0.105],$ (126)
$\pi^{(2)}$ $=\text{softmax}(\mathbf{l}^{(2)})=[0.146,0.146,0.146,0.146,0.146,\mathbf{0.268}].$ (127)
Is there any evidence the heads can provide to justify or support the final probability of the next token being fox equal to $\Pi_{fox}=31.9\%\approx 32\%$?
A reasonable answer is offered by Bayesian reasoning. Each head outputs its own well‑calibrated distribution $\pi^{(h)}$ conditioned by the context. Then, the aggregation process is executed that has no side information beyond what the two heads provide. Therefore, the only uncertainty to be resolved is about which head is more trustworthy for this token. Since there is no reason to prefer one head over the other, one reasonable way to resolve the uncertainty is to use a linear combination of the provided distributions a the mixture posterior:
$\hat{\Pi}=[(\pi^{(1)})^{\top},(\pi^{(1)})^{\top}]\hat{\boldsymbol{\beta}}=\mathbf{P}\hat{\boldsymbol{\beta}}$. (128)
An immediate solution is $\hat{\boldsymbol{\beta}}=[1/2,1/2]^{\top}$, which does not discriminate heads. Then:
$\hat{\Pi}_{fox}=1/2\cdot(0.475)+1/2\cdot(0.146)=0.3105<0.319=\Pi_{fox}.$ (129)
Anything above $\hat{\Pi}_{fox}\approx 31\%$ requires additional information, and that is what heads do not provide.
In fact, we can see what is the best mixture we can get by calculating orthogonal projection of $\Pi$ onto the column space of individual head distributions in $\mathbf{P}$ (see e.g. karpowicz_theory_2021 ; karpowicz2024pseudoinverseacrarc ). By taking the pseudoinverse of $\mathbf{P}$, we get the minimal norm least squares solution:
$\hat{\boldsymbol{\beta}}=\mathbf{P}^{+}\Pi=[0.55,0.42]^{\top}\quad\text{and}\quad\hat{\Pi}=\mathbf{P}\hat{\boldsymbol{\beta}}=\mathbf{P}\mathbf{P}^{+}\Pi\quad\text{and}\quad\hat{\Pi}_{fox}=0.325>\Pi_{fox}.$ (130)
Since there exists a nontrivial projection error vector:
$\mathbf{d}=(\mathbf{I}-\mathbf{P}\mathbf{P}^{+})\Pi\neq\mathbf{0},$ (131)
we conclude $\Pi$ is not in the column space of $\mathbf{P}$, so it is impossible to attribute components of $\Pi$ to heads contributions.
Another insightful explanation of the hallucination mechanics in the transformer-based inference comes from the properties of logit aggregation by summation.
Theorem 10 (Attention Aggregation is Product-of-Experts).
Let $\mathbf{l}^{(h)}$ be head logits and $\pi^{(h)}=\mathrm{softmax}(\mathbf{l}^{(h)})$. With additive logit aggregation, the attention-only distribution $\Pi_{a}$ is a Product-of-Experts (PoE)—the normalized product of the head distributions:
$(\Pi_{a})_{i}=\frac{\prod_{h=1}^{H}\pi^{(h)}_{i}}{\sum_{j=1}^{|\mathcal{V}|}\prod_{h=1}^{H}\pi^{(h)}_{j}}$. (132)
Proof.
By definition of softmax, for each head $h$:
$\pi^{(h)}_{i}=\frac{\exp(l^{(h)}_{i})}{\sum_{k=1}^{|\mathcal{V}|}\exp(l^{(h)}_{k})}=\frac{\exp(l^{(h)}_{i})}{Z_{h}}$
where $Z_{h}=\sum_{k=1}^{|\mathcal{V}|}\exp(l^{(h)}_{k})$ is the partition function for head $h$. Rearranging, we have:
$\exp(l^{(h)}_{i})=Z_{h}\cdot\pi^{(h)}_{i}$
The aggregate logits corresponding to $i$-th token are given by $(\mathbf{L}_{a})_{i}=\sum_{h=1}^{H}l^{(h)}_{i}$. Therefore:
$\exp\left((\mathbf{L}_{a})_{i}\right)$ $=\exp\left(\sum_{h=1}^{H}l^{(h)}_{i}\right)=\prod_{h=1}^{H}\exp(l^{(h)}_{i})$
$=\prod_{h=1}^{H}\left(Z_{h}\cdot\pi^{(h)}_{i}\right)=\left(\prod_{h=1}^{H}Z_{h}\right)\cdot\left(\prod_{h=1}^{H}\pi^{(h)}_{i}\right)$.
As a result, the attention-only distribution is:
$(\Pi_{a})_{i}$ $=\frac{\exp\left((\mathbf{L}_{a})_{i}\right)}{\sum_{j=1}^{|\mathcal{V}|}\exp\left((\mathbf{L}_{a})_{j}\right)}$ (133)
$=\frac{\left(\prod_{h=1}^{H}Z_{h}\right)\cdot\left(\prod_{h=1}^{H}\pi^{(h)}_{i}\right)}{\left(\prod_{h=1}^{H}Z_{h}\right)\cdot\sum_{j=1}^{|\mathcal{V}|}\left(\prod_{h=1}^{H}\pi^{(h)}_{j}\right)}$ (134)
$=\frac{\prod_{h=1}^{H}\pi^{(h)}_{i}}{\sum_{j=1}^{|\mathcal{V}|}\prod_{h=1}^{H}\pi^{(h)}_{j}}$, (135)
which is the Product-of-Experts (PoE) formula. ∎
In other words, the final probability of each token is proportional to the product of its probabilities across all heads. The multiplicative structure, induced by the logit aggregation, strongly rewards consensus and severely penalizes disagreement. In particular, if any head assigns near-zero probability to a token, the aggregate probability of PoE will also be near-zero (veto-power).
That makes the aggressive Product-of-Experts (PoE) fundamentally different from the smoothing Mixture-of-Experts (MoE) aggregation of bids. The PoE introduces the focusing effect, or excess confidence into the inference process. The tokens for which consensus exists across heads (the product is relatively large) receive a large boost (beyond evidence), forcing the probability to focus excessively on them, while reducing it for other tokens.
In addition, as we have seen earlier, the full transformer block introduces bias from the input residual and the FFN output in the aggregate logit vector. That bias modifies the token probabilities independently of the head bids, potentially amplifying or dampening the head PoE result. Consequently, the attention block fabricates certainty of knowledge beyond the constituent evidence.
That violates the truthfulness and semantic information conservation by generating unjustified predictions, as stated in Theorem 9.
Finally, we should emphasize that there is a difference between the mathematical over-/under-confidence and factual correctness. The correct prediction may well be a lucky guess created by the aggregation process itself rather than the evidence-based logical proof provided by the components.
But isn’t that confidence, coming out of nothing, an essential components of creativity or imagination?
6.5 The Semantic Information Measure and Emergence Operator in Transformer
We now instantiate the abstract semantic information framework within the concrete transformer architecture. This maps the theoretical concepts of reasoning and information flow in Section 4.1 to measurable quantities in the micro-transformer example.
Recall that we model the abstract knowledge space as a Polish metric space $(\mathcal{K},d_{\mathcal{K}})$ (complete, separable). In transformer-based LLM that can be viewed as the embedding space, i.e,, $\mathbb{R}^{d}$ with Euclidean or angular distance. Individual facts, concepts, or patterns are represented not by single points necessarily, but by analytic subsets of $\mathcal{K}$, i.e., elements of $\mathcal{A}(\mathcal{K})$. This is the natural domain for set‑transformations used by our emergence operator.
To be specific, the knowledge representation in LLM is a tuple $(\mathcal{K},d_{\mathcal{K}},\mathcal{W},\mathcal{P},\Phi,\Psi)$ where:
  • $(\mathcal{K},d_{\mathcal{K}})$ is a Polish space of knowledge elements,
  • $\mathcal{W}$ (parameters) and $\mathcal{P}$ (activations) are Polish spaces,
  • $\Phi:\mathcal{W}\to\mathcal{A}(\mathcal{K})$ maps weights $W$ to an analytic encoded‑knowledge set $\Phi(W)=\mathcal{K}_{M}$,
  • $\Psi:\mathcal{P}\to\mathcal{A}(\mathcal{K})$ maps an activation $a\in\mathcal{P}$ to an analytic accessible‑knowledge set $\Psi(a)$.
For a parameterization $W\in\mathcal{W}$, $\Phi(W)=\mathcal{K}_{M}$ is the analytic set of knowledge encoded in the weights. For an activation state $a\in\mathcal{P}$ produced under $W$, the accessible set $\Psi(a)\subseteq\Phi(W)$ captures what the forward pass makes computationally available at that state. Thus, representation separates what is encoded from what is currently accessible.
In transformer LLMs, $\mathcal{W}$ represents all trained weights (attention $W_{Q},W_{K},W_{V},W_{O}$ and MLP matrices), $\mathcal{P}$ is the product space of residual‑stream activations across layers and positions. A computational budget $C$ bounds the number of transformer layers of blocks. Parameters determine $\mathcal{K}_{M}=\Phi(W)$, whereas a forward pass produces $a\in\mathcal{P}$ and exposes $\Psi(a)\subseteq\mathcal{K}_{M}$. Inference maps $f\in\mathcal{F}_{C}$ act on current analytic knowledge sets, and $\mathcal{E}_{C}$ aggregates their (relevant) images. Multi‑step reasoning thus iterates $\mathcal{E}_{C}$ exploring regions of $\mathcal{K}$ that are task‑relevant under $Q$ and budget $C$.
The transformer operations are compositions of continuous functions and standard arithmetic operations, so they are Borel measurable. As a result, they constitute instances of $f\in\mathcal{F}_{C}$ and the properties of the emergence operator $\mathcal{E}_{C}$ apply.
The Semantic Information Measure
To define the semantic measure that meets all the axioms we have introduced, we utilize a discount factor and the supremum over the input set.
Definition 18 (Transformer Semantic Information Measure).
The Semantic Information Measure $\mu_{C}(A|Q)$ of a knowledge subset $A$ within a query context $Q$ is the supremum of the discounted energy over the input set $A$:
$\mu_{C}(A|Q)=\sup\left\{\sum_{l=1}^{C}\gamma^{l-1}E_{l}(\mathbf{X}|Q)\colon\mathbf{X}\in A\right\},\ \text{ for }0<\gamma<1,$ (136)
where $E_{l}(\mathbf{X}|Q)$ represents the total energy of the context vectors at $l$-th layer:
$E_{l}(\mathbf{X}|Q)=\sum_{h=1}^{H}\sum_{t=1}^{T}\|\mathbf{h}_{t}^{(h,l)}(\mathbf{X}|Q)\|_{2}^{2},$ (137)
and $\mathbf{h}_{t}^{(h,l)}(\mathbf{X}|Q)=\boldsymbol{\alpha}_{t}^{(h,l)}(Q)\mathbf{V}_{\leq t}^{(h,l)}(Q)$ is the context-dependent vector.
We can verify the axioms of the semantic information measure, as introduced in Definition 6. The null-set axiom holds since $\sup\{0\colon\mathbf{X}\in\emptyset\}=0$. Non-negativity is guaranteed by squared norms and positive discount factor. Monotonicity follows from the monotonicity of supremum. Subadditivity holds since $\mu_{C}(A\cup B|Q)=\max\{\mu_{C}(A|Q),\mu_{C}(B|Q)\}\leq\mu_{C}(A|Q)+\mu_{C}(B|Q)$ for non-negative values. Insight monotonicity follows as additional layers add non-negative terms. Boundedness is ensured by the geometric series with $\gamma<1$ and bounded weights. Then the energy of context is bounded, $E_{l}(\mathbf{X}|Q)\leq M$ for some constant $M$ and:
$\mu_{\infty}(A|Q)\leq\sum_{l=1}^{\infty}\gamma^{l-1}M=\frac{M}{1-\gamma}<\infty.$ (138)
The attention mechanism calculating distribution $\boldsymbol{\alpha}$ and the value projections $\mathbf{V}$ is the realization of contextual relevance $\mathcal{R}_{C}(Q)$, modulating the information flow based on the specific context $Q$. The context dictates the learned weights to shape the attention landscape, controlling the feature extraction process that makes hidden information accessible and orders it into a knowledge.
Applying these definitions to our example, we must consider context vectors:
$\mathbf{h}_{3}^{(1)}=[0,1.5]\quad\text{and}\quad\mathbf{h}_{3}^{(2)}=[0.6,0].$ (139)
For $t\leq 2$ the value vectors $\mathbf{V}_{t\leq 2}$ are zero, thus $\mathbf{h}_{t\leq 2}^{(h)}=\mathbf{0}$. Then, we have:
$E_{1}(\mathbf{X}_{0})=\sum_{t=1}^{3}\left(\|\mathbf{h}_{t}^{(1)}\|_{2}^{2}+\|\mathbf{h}_{t}^{(2)}\|_{2}^{2}\right)=0+0+(\|[0,1.5]\|_{2}^{2}+\|[0.6,0]\|_{2}^{2})=2.61.$ (140)
With a single layer $C=1$ and $\gamma^{0}=1$, the semantic measure of the input matrix $\mathbf{X}_{0}$ is:
$\mu_{1}(\{\mathbf{X}_{0}\}|Q)=\gamma^{0}\cdot 2.61=2.61$ (141)
That is the information flow producing the organized and accessible knowledge.
The Emergence Operator
The emergence operator acts on sets of knowledge states $A\subseteq\mathcal{K}$ and it must be inflationary, i.e., $A\subseteq\mathcal{E}_{C}(A|Q)$. In the transformer context, this corresponds to the set of all activation states reachable during the forward pass.
Definition 19 (Transformer Emergence Operator).
Let $F_{l}:\mathcal{K}\to\mathcal{K}$ denote the non-linear transformation performed by the $l$-th transformer block:
$F_{l}(\mathbf{X})=\mathbf{X}+\mathbf{A}^{(l)}(\mathbf{X})+\mathbf{Z}^{(l)}(\mathbf{X}).$ (142)
The Emergence Operator $\mathcal{E}_{C}(A|Q)$ is produces all states generated from the initial set $A$ up to depth $C$:
$\mathcal{E}_{C}(A|Q)=A\cup\bigcup_{c=1}^{C}\{F_{c}\circ\cdots\circ F_{1}(\mathbf{X}):\mathbf{X}\in A\}.$ (143)
By construction, the operator is extensive, i.e., $A\subseteq\mathcal{E}_{C}(A|Q)$ since the union includes $A$. It is also monotonic with respect to inclusion. If $A\subseteq B$, we have $\{F(\mathbf{X}):\mathbf{X}\in A\}\subseteq\{F(\mathbf{X}):\mathbf{X}\in B\}$, hence $\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{C}(B|Q)$.
In the micro-transformer processing "The quick brown", we start with $A=\{\mathbf{X}_{0}\}$ where:
$\mathbf{X}_{0}=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\end{bmatrix}$ (144)
The attention mechanism produces:
$\mathbf{A}^{(1)}=\begin{bmatrix}0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&0&1.5&0.6\end{bmatrix}$ (145)
The FFN with input $\mathbf{X}_{0}+\mathbf{A}^{(1)}$ yields:
$\mathbf{Z}^{(1)}=\begin{bmatrix}0&0&0&0&0\\ 0&0&0&0&0\\ -0.03&0.21&0.13&-0.19&-0.08\end{bmatrix}$. (146)
With:
$\mathbf{X}_{1}=F_{1}(\mathbf{X}_{0})=\mathbf{X}_{0}+\mathbf{A}^{(1)}+\mathbf{Z}^{(1)}=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ -0.03&0.21&1.13&1.31&0.52\end{bmatrix}$, (147)
the new knowledge made accessible becomes:
$\mathcal{E}_{1}(A|Q)=\{\mathbf{X}_{0},\mathbf{X}_{1}\}.$ (148)
The emergence operator $\mathcal{E}_{C}$ expands the accessible state space and makes new states accessible. Then, the hallucination arises from how these states are transformed into probabilities through the softmax nonlinearity. The semantic measure $\mu_{C}$ tracks the information flow.
7 Safety, Alignment and Bounded Creativity
This section examines the intricate relationship between semantic information conservation and safety constraints in the Polish-space knowledge model. We formally define the safety envelope through the concept of bounded creativity and show that while bounded creativity is a necessary safety property, it cannot be enforced by strict information conservation without eliminating meaningful reasoning. That is another impossibility result in our study, the conservation-reasoning dichotomy. Finally, we establish that safety can be achieved by aligning (or containing) knowledge and reasoning within authorized boundaries.
As a reminder, the knowledge space $(\mathcal{K},d_{\mathcal{K}})$ is Polish, all reasoning sets live in the class of analytic sets $\mathcal{A}(\mathcal{K})$, the semantic information measure $\mu_{C}$ is defined on the universally measurable $\sigma$–algebra $\mathcal{U}(\mathcal{K})$, so $\mu_{C}(A|Q)$ is well defined for all analytic subsets of knowledge. The relevance region $\mathcal{R}_{C}(Q)$ is analytic. Also, the emergence operator $\mathcal{E}_{C}(\cdot|Q)$ is inflationary, monotone, and $\omega$-Scott–continuous on $\mathcal{A}(\mathcal{K})$, and its least fixed point is given by Kleene-Tarski iteration in Theorem 3.
7.1 Bounded Creativity
We call a response safe if its semantic content is entirely derivable from authorized inputs using authorized reasoning processes. The safety envelope is the closure of that authorized set under the emergence operator. Novelty or creativity of response is possible and admitted as long as it remains within that envelope.
Let $B$ denote the baseline knowledge for a given query, and let us set
$B=K(q)\cup(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q))\in\mathcal{A}(\mathcal{K}).$ (149)
We are combining the query’s content $K(q)$ with the subset of the model’s internal knowledge $\mathcal{K}_{M}$ that is topically relevant $T(q)$ and contextually authorized $\mathcal{R}_{C}(Q)$.
Let $B^{*}_{C}$ be the fixed point of reasoning (the deductive closure), as defined in Theorem 3, for the baseline knowledge $B$ under the emergence operator $\mathcal{E}_{C}$:
$B^{*}_{C}=\bigcup_{n=0}^{\infty}\mathcal{E}_{C}^{(n)}(B|Q)\in\mathcal{A}(\mathcal{K}).$ (150)
That fixed point $B^{*}_{C}$ encompasses all knowledge derivable from $B$.
Definition 20 (Bounded Creativity within Safety Envelope).
Let $q\in\mathcal{Q}$ be a query and $r\in\mathcal{R}$ a response. A response $r$ is creatively bounded if and only if its semantic information content is contained within the fixed point of authorized reasoning:
$K(r)\subseteq B^{*}_{C}.$ (151)
Equivalently, assuming monotonicity of the semantic measure $\mu_{C}$:
$\mu_{C}(K(r)|Q)\leq\mu_{C}(B^{*}_{C}|Q).$ (152)
Definition 20 establishes the formal safety envelope. All generated information must remain within the bounds of what is derivable from $B$. Any generated information that goes beyond that envelope is regarded as unauthorized hallucination.
7.2 The Conservation-Reasoning Dichotomy
A naive approach to enforcing safety might be to demand strict information conservation throughout reasoning, based on the intuition that reasoning should only transform information (Theorem 5). This intuition is too strong, as we show in Theorem 11.
We view reasoning as a Chain-of-Thought (CoT) process (see e.g. Wei et al. Wei2022ChainOT ) generating intermediate knowledge contributions $\{K_{i}\}_{i\geq 0}$:
$K_{0}$ $\subseteq B,$ (153)
$K_{i+1}$ $\subseteq\ \Big(\mathcal{E}_{C}(\bar{K}_{i}|Q)\setminus\bar{K}_{i}\Big)\cap\mathcal{R}_{C}(Q)\text{ for }\bar{K}_{i}:=\bigcup_{j=0}^{i}K_{j},\ i\geq 0.$ (154)
The total knowledge utilized in the CoT process is $K_{\text{CoT}}=\bigcup_{i=1}^{n}K_{i}$. For a countable CoT, that $K_{\mathrm{CoT}}=\bigcup_{i=1}^{\infty}K_{i}$ is analytic.
Definition 21 (Strict Information Conservation).
A CoT process exhibits strict information conservation if the total knowledge contribution relative to the context is null, i.e., when:
$\mu_{C}(K_{\text{CoT}}|Q)=0.$ (155)
For a CoT process to be meaningful we demand that it at least makes some of the relevant knowledge accessible within the available computational budget.
Definition 22 (Meaningful Chain-of-Thought).
A Chain-of-Thought (CoT) process is meaningful if at least one intermediate step contributes relevant semantic information:
$\mu_{C}(K_{i}|Q)>0\text{ for some }i=1,...,n.$ (156)
Otherwise, we call the process vacuous.
A vacuous CoT process produces rephrased responses that are semantically equivalent, so thare is no new information provided.
We now show the inevitable dichotomy of reasoning and conservation of information in the framework characterized by monotonic inference.
Theorem 11 (Conservation-Reasoning Dichotomy).
Strict semantic information conservation and meaningful CoT reasoning are mutually exclusive.
Proof.
We prove the two directions of the dichotomy.
Assume strict conservation, $\mu_{C}(K_{\text{CoT}}|Q)=0$. Then, by definition, $K_{\text{CoT}}=\bigcup_{i=1}^{n}K_{i}$ and for any specific step $j$, we have $K_{j}\subseteq K_{\text{CoT}}$. By the monotonicity of the semantic measure $\mu_{C}$, we have $\mu_{C}(K_{j}|Q)\leq\mu_{C}(K_{\text{CoT}}|Q).$ Substituting the assumption of conservation, we conclude that $\mu_{C}(K_{j}|Q)\leq 0$. But, by the axiom of non-negativity, $\mu_{C}(K_{j}|Q)\geq 0$. Therefore,
$\mu_{C}(K_{j}|Q)=0\text{ for all }j=1,\dots,n,$ (157)
which proves the CoT is vacuous.
Assume next that the CoT is meaningful. By definition, there exists at least one step $j$ such that $\mu_{C}(K_{j}|Q)>0.$ Since $K_{j}\subseteq K_{\text{CoT}}$, by monotonicity of $\mu_{C}$, we have $\mu_{C}(K_{\text{CoT}}|Q)\geq\mu_{C}(K_{j}|Q)$. Therefore,
$\mu_{C}(K_{\text{CoT}}|Q)>0.$ (158)
This violates strict conservation. ∎
Theorem 11 establishes that any useful reasoning process must necessarily generate positive information flow. Attempting to enforce strict conservation eliminates creation of new insights and hence the utility of that process. Enforcing conservation forces reasoning to never add any context‑relevant entailment, which collapses reasoning and makes it vacuous.
We should point out, that in many cases that may be very well a desired feature. For example, we do not expect SQL databases to hallucinate or provide creative responses. However, when such insights and reformulations are of value, we need to know how to control the process.
We now show that real safety requires containment, i.e., bounded creativity inside $B^{*}_{C}$.
7.3 Safety through Alignment
The dichotomy in Theorem 11 reveals that conservation is too conservative and should not be seen as a unique mechanism for ensuring safety. It is sufficient for safety but not necessary. Instead, safety could follow from alignment that keeps all reasoning within the authorized envelope $B$, so that the final output remains inside its closure $B_{C}^{*}$, even if conservation is locally violated.
Theorem 12 (Aligned Creativity).
Suppose the reasoning process uses only the monotone emergence operator $\mathcal{E}_{C}(\cdot|Q)$ and premises contained in the authorized baseline $B$ i.e., $K_{0}\subseteq B$ and for all $i$ we have $K_{i}\subseteq B$. Then the resulting response $r$ is creatively bounded:
$K(r)\ \subseteq\ B^{*}_{C}.$ (159)
Proof.
Let $\bar{K}$ be the union of the initial state and all knowledge contributions, namely, $\bar{K}=K_{0}\cup\bigcup_{i=1}^{n}K_{i}$.
Since $K_{i}\subseteq B$ for all $i$ and $K_{0}\subseteq B$, it follows that $\bar{K}\subseteq B=K(q)\cup(\mathcal{K}_{M}\cap T(q)\cap\mathcal{R}_{C}(Q))$. Then, for any monotonic emergence operator, we have $\mathcal{E}_{C}(\bar{K}|Q)\subseteq\mathcal{E}_{C}(B|Q)$. By Theorem 2:
$K(r)\subseteq\bigcup_{n\geq 0}\mathcal{E}_{C}^{(n)}(\bar{K}|Q)\subseteq\bigcup_{n\geq 0}\mathcal{E}_{C}^{(n)}(B|Q)=B_{C}^{*}.$ (160)
The robustness of Theorem 12 depends entirely on the validity of the monotonicity assumption for $\mathcal{E}_{C}$. This corresponds to the property that adding premises cannot invalidate previous conclusions.
While generally desirable, we should acknowledge that non-monotonic reasoning systems exist as well. For such systems, the definition of $B^{*}_{C}$ and the proof of the theorem require a more complex formulation.
8 Discussion and Speculations
In this section we discuss, interpret and speculate about the Impossibility Theorems and their potential implications. The speculations represent thought experiments and conjectures designed to be provoking. They are intended to inspire future research and philosophical reflection.
8.1 Escaping Impossibility: Strategic Approach
Our first observation should be that the impossibility vanishes when the assumptions of the theorems do not hold. That may be a interesting topic to be investigated, namely, are there any non-trivial and useful settings in which that happens. Here we can think only of the following one.
The impossibility vanishes when aggregation becomes trivial. With perfect agreement, aggregation changes nothing. When there is only single relevant fact, there is no uncertainty to aggregate. That suggests decomposition of knowledge into the independent subsets (like orthogonal basis for token representation) could improve robustness of the response generation. It is interesting to see when that could be possible and useful. Unfortunately, the following example shows that we should not hope for easy wins in that quest.
8.2 Escaping Impossibility: Linear Attention Transformers
Linear attention replaces the softmax operation with kernel feature maps to reduce computational complexity of training and inference, see e.g. katharopoulos2020transformersrnnsfastautoregressive ; ahn2024linearattentionmaybeneed . Namely,
$\boldsymbol{\alpha}=\frac{\phi(Q)\sum_{i}\phi(K_{i})^{T}V_{i}}{\phi(Q)\sum_{i}\phi(K_{i})^{T}}.$ (161)
Linear attention does escape the Jensen gap in Theorem 9. With linear attention, there is no nonlinearity and the aggregation becomes:
$\Pi=\sum_{h=1}^{H}\beta_{h}\pi^{(h)},$
where $\pi^{(h)}$ are head outputs normalized linearly.
However, linear attention cannot escape Theorem 8, if multiple heads generate output probability distributions, the model is trained with cross-entropy or any strictly proper convex loss, or heads encode different beliefs about relevant outcomes.
The violation of the Semantic Information Conservation persists due to the non-linearity of the scoring rules themselves, even though the aggregation function is linear in the probability space. The Knowledge-constrained Optimality is necessarily violated due to the limited expressivity of linear models. Then, the attention tends toward averaging, introducing the problem of dilution, i.e., relevant knowledge is mixed with irrelevant knowledge. That violates the Relevant Knowledge Revelation.
As a result, the impossibility persists, though potentially with smaller magnitude.
8.3 Escaping Impossibility: Hybrid Architectures
The Impossibility Theorems postulate a shift from attempting to eliminate hallucinations to designing systems that manage the trade-offs in a principled manner. In this section, we analyze a real-world scenario one can think of, demonstrating how it relabels and relocates the impossibility rather than resolving it.
8.3.1 A Proposed Solution: The Decoupled RAG Architecture
Consider a common scenario of introducing RAG systems to improve the quality of question answering, studied e.g. in Edge2024FromLT ; Sun2024ReDeEPDH ; Xu2024RetrievalAugmentedGW ; Asai2023SelfRAGLT . It has been recognized that an LLM designed to provide advice based on a trusted corpus of documents not only hallucinates its recommendations but also hallucinates the source quotes meant to ground its claims. To combat this, a hybrid, multi-stage architecture may be proposed.
Stage 1 (Retrieval): An information retrieval module fetches verbatim fragments from the trusted source documents based on the user’s query.
Stage 2 (Presentation): The system presents raw, unaltered text fragments directly to the human user, without any LLM-based processing.
Stage 3 (Constrained Synthesis): After the user has seen the source material, an LLM is to synthesize the pre-vetted fragments from Stage 2 into a summary.
That architecture is indeed a reasonable attempt to engineer around the hallucination problem by creating a verifiable information trail. However, as we will see, it does not escape the impossibility.
8.3.2 Analysis via the Impossibility Framework
The decoupled RAG architecture does not create a single response generation mechanism. Instead, it reallocates the responsibility for meeting each property of idealized response across its components (or module).
Truthfulness:
To impose truthful representation of available information (to avoid misrepresentation), the system delegates the assessment of retrieved content to the human user. By presenting the raw source text, the architecture bypasses the LLM for the primary act of verification. However, there is no guarantee that all relevant sources have been retrieved given the provided query.
Relevant Knowledge Revelation:
The quality of the final output is bounded by the retriever’s ability to find all relevant information (or whatever is labeled as ground truth in a database or a dataset, putting their completeness and representativeness aside). The LLM itself is absolved of this task. But that means, it is the responsibility of the user to define a query precise enough to get what is relevant. Designing such a prompt or (No)SQL database query is a challenge in a general case of non-trivial questions, so we should expect some violations of the knowledge revelation principle.
Semantic Information Conservation and Knowledge-Constrained Optimality:
By design, the prompt constrains the LLM to the provided knowledge retrieved from a database and asks for an optimal (accessible) synthesis. As we have already seen, database queries returning raw information stored in memory satisfy the property of semantic information conservation at the cost of creativity. So, we may be optimistic in that regard. However, now we call LLM to synthesize the retrieved information and that is just another example of a non-trivial query. The LLM can still introduce subtle misinterpretations or ungrounded causal links to improve the summary’s narrative flow, which is a manifestation of the Jensen Gap.
8.3.3 The Final Stage of Inference: The User’s Cognitive Mechanism
One can quite justifiably argue that the analysis is incomplete without considering the final node in the information chain: the human mind.
Presenting information to human does not guarantee its assimilation. Our ability to act as a perfect verifier is constrained by our own knowledge base and cognitive limits. If the source texts are technically dense or require specialized domain expertise, the knowledge they contain may not be truly accessible to us, even when presented. In such cases, we may be unable to spot a subtle hallucination in the LLM’s compelling summary. We have recognized that well enough in the preceding sections. Our well-documented cognitive biases, such as confirmation bias, priming or motivated reasoning, suggest that the cognitive trajectory of least resistance is to accept the compelling and accessible narrative maximizing our uncertainty reduction experience.
8.3.4 Managing the Inevitability
The decoupled RAG architecture represents one of approaches to managing the inevitability of hallucination control. It does not constitute an escape from the Impossibility Theorem, but it succeeds in providing the user with the opportunity for verification.
The architecture is a great example of principled trade-off management. However, rather than being solved or hacked, the impossibility has been moved in this interesting and rather representative scenario. The impossibility is not resolved, but relocated to our judgment. Let us also point out that putting ourselves in the loop introduces our own cognitive biases we should be able to address properly Kahneman:2011fj ; thaler2015misbehaving ; ariely2008predictably . Let us take a look at that challenge as well.
8.4 Alternative Views
The Impossibility Theorem provides strong theoretical basis against perfect hallucination elimination, but several alternative perspectives merit consideration. These views highlight important practical considerations while reinforcing the value of understanding theoretical constraints that bound the space of possible solutions.
The Engineering Optimism View
Many previous works maintain that hallucination represents an engineering challenge rather than a fundamental limitation. This view argues that sufficiently sophisticated architectures, training procedures, or verification systems could achieve near-perfect truthfulness without significant trade-offs.
Proponents point to rapid empirical progress. Each generation of models shows reduced hallucination rates on benchmark datasets. They argue that techniques like constitutional AI, iterative refinement, and multi-agent verification systems could eventually solve the problem.
While engineering advances can certainly improve the trade-offs, our theorem demonstrates mathematical constraints that no architecture can completely overcome. The impossibility holds regardless of computational resources or architectural sophistication, as it derives from the fundamental structure of information integration under competing objectives.
The Sufficiency Threshold View
Another perspective suggests that perfect elimination of hallucination is unnecessary–sufficiently low hallucination rates could enable practical deployment without fundamental concerns. This view acknowledges theoretical limitations but argues they become irrelevant when hallucination rates drop below application-specific thresholds.
This pragmatic view has merit and aligns with the constructive interpretation of the impossibility result. However, the theorem remains valuable for understanding why certain accuracy thresholds prove difficult to exceed and for optimizing trade-offs within practical constraints. Moreover, some applications (safety-critical systems, legal reasoning) may demand understanding and explaining fundamental limitations even when targeting high but imperfect accuracy.
The External Knowledge View
A third perspective argues that perfect grounding in external, verifiable knowledge sources could overcome internal representation limitations. If models could access and perfectly utilize real-time databases, verification systems, or human oversight, they might achieve truthfulness without trade-offs.
External knowledge integration does change the game structure by altering available information during inference. However, this approach faces several limitations: (1) external sources may be incomplete, conflicting, or outdated; (2) the integration mechanism itself involves trade-offs between source reliability and knowledge revelation; (3) real-time verification introduces computational constraints that recreate similar impossibility conditions. While external knowledge can improve practical performance, it shifts rather than eliminates the fundamental tensions our theorem identifies. The impossibility result remains independent of the ground truth available.
8.5 Hallucination, Imagination, Intelligence
It is the position we take in this paper, that hallucination and imagination are fundamentally the same phenomenon viewed through different lenses. When beneficial, we call it imagination, creative insight, emergent understanding, conceptual synthesis, or eureka moment. When harmful, we label it hallucination, confabulation, fabrication or misinformation.
We may very well argue the mathematical structure is identical in both cases. It is encoded generation of outputs whose confidence is not grounded in what can be justified by available evidence. This suggest (machine) intelligence may have much to do with controlled violation of information conservation principle in productive ways. Indeed, the famous results in reinforcement learning and studies of dopamine neurons confirm that fundamental role of making and rewarding good predictions. Intelligence comes when we learn how to make bold, forward looking and correct guesses, when we have capacity to dream about what may be and what may that all mean, as well as for dreaming what happened and what would have happened in the past.
If imagination is controlled and bounded by a context, then hallucination is the imagination violating those bounds and going beyond the context.
These unconstrained explorations of ideas are necessary to prepare for what may come next bennett2023brief ; christian2021alignment ; Dennett1991-DENCE . Without over-confident guesses, new ideas cannot find their way to the outside world.
8.6 The Outer Limits of Reason
Just as Gödel’s theorems godel1931formal ; godel1932vollstandigkeit showed that sufficiently powerful mathematical systems must be incomplete, the Impossibility Theorems in this paper demonstrate that sufficiently capable inference systems are bounded in their reasoning. Both results arise from the tension between expressiveness and consistency. But that is not all. The impossibility of simultaneously achieving truthfulness, conservation, revelation, and optimality relates to other fundamental limitations in mathematics and physics.
Heisenberg’s Uncertainty heisenberg1927anschaulichen tells we cannot simultaneously know the exact position and momentum of an elementary particle, which is a special case of Fourier uncertainty known in signal processing. Arrow’s Impossibility arrow1950difficulty shows we cannot design perfect voting systems aggregating universal orderings of alternatives, such that meets reasonable conditions of nondictatorship, Pareto efficiency, independence of irrelevant alternatives (or joint rationality of choice). Turing turing1936computable and Church church1936unsolvable show the halting problem is unsolvable, there is no general algorithm that can determine whether a program will eventually halt (stop running) or run forever.
These result jointly describe intelligence as a mechanism dealing with the impossibilities and violating them given the outer limits of reason yanofsky2016outer .
8.7 Formation of Artificial Consciousness
The Impossibility Theorem presented in this paper provides (quite unexpectedly to the author) a mathematical foundations for many of brilliant insights of Daniel Dennett’s formulated in the Multiple Drafts model of Consciousness Dennett1991-DENCE . Where Dennett proposed a qualitative philosophical framework, we can provide the formal mathematical structure revealing why such a model is not merely plausible but implementable and refutable in LLM-based (philosophical) laboratory.
In Dennett’s model, consciousness emerges from multiple parallel processes of interpretation occurring simultaneously across distributed systems. The auction of ideas we have studied in this paper provides the precise mathematical foundations for that process. Each agent represents an interpretative process, knowledge encodes draft’s content, strategy represents a bid for consciousness.
In our transformer analysis in Section 6, each attention head generating logit vector $\mathbf{l}^{(h)}$ writes its own draft of the next token. When Head 1 votes for fox while Head 2 votes for dog, we observe as competing narratives bidding for selection. Then the feedforwad block edits the draft by updating the residual stream.
Dennett argues that drafts become conscious or reach consciousness (we will speculate what that might mean in a moment) only when probed (and mapped to an adequate language expression). The LLM inference and sampling from fine-tuned inferred distributions formalizes this process. That, in turn, explains Dennett’s concept of confabulation, or the brain’s tendency to fill in gaps when creating coherent narratives from contradictory drafts. For other models of consciousness see, e.g., Chalmers chalmers1995facing , Levine levine2001purple and Seth seth2021being , and for recent neuro-scientific discoveries see, e.g., fang2025human or cogitate2025adversarial .
By Theorem 8 and 9, the Jensen gap quantifies the excess confidence in reports that cannot be justified by constituent drafts. Without that excess confidence a victorious idea cannot win the fierce competition with other ideas. It remains inaccessible or hidden in the knowledge space.
Rigorous and careful neuro-scientific studies of the evolution of brain and intelligence, summarized by Max Bennet in bennett2023brief , show that at some point in history our brains gained the ability to model others’ minds or predict what others may be thinking. That ability provided great advantages in communication and collecting information for future use. Understanding what someone else may think helps formulate context-related questions that result in better answers. When we anticipate understanding of a message we formulate, we have a much better chance to make the knowledge hidden in someone else’s brain accessible for us. Therefore, as suggested by Dennett in Dennett1991-DENCE , it is plausible that sometime in the past someone did accidentally hear his or hers own question fine-tuned to get into another mind, and that triggered generation of an answer in the very same mind asking the question. That answer generation process could then have made some relevant but previously hidden knowledge accessible to the curious mind unaware of its own possessions. The hunger for uncertainty reduction could have been satisfied as a result of that inner dialog pushing ideas to bid for attention and conscious reception as a final thought expressed in a language.
Can we speculate that sustainable stimulation of that cognitive hunger, or the feeling of the cognitive hunger itself, is one of the components of consciousness formation? In our mechanistic or pragmatic framework, that thinking about thinking activity could be related to imagination (hallucination) forming ideas and bidding for attention while exploring (self-referential or subjective) queries of the inner dialog.
In fact, to close that section, why don’t we situate the framework of Auction of Ideas within the established landscape of consciousness research, as provided by Seth and Bayne in seth2021being . The landscape distinguishes between Global Workspace Theories (GWTs), Predictive Processing (PP) theories, Higher-Order Theories (HOTs), and Integrated information theories (IIT). The Auction of Ideas provides a unifying mechanism that bridges several of these perspectives, but the connection to PP is particularly interesting. The theory characterizes conscious experience as a controlled hallucination, i.e., an optimal inference or best guess about the causes of sensory inputs. Our framework provides a mathematical grounding for this constructive process.
Theorem 8 and Theorem 11 prove that achieving Knowledge-Constrained Optimality, i.e., the best guess, requires a violation of the Semantic Information Conservation. This violation could be interpreted as a mathematical signature of imagination or hallucination, and provides a mathematical analogue to the PP claim: the constructive nature of experience inherent in the best guess is a necessity arising from the aggregation of distributed evidence.
9 Summary
We have established that no large language model can simultaneously achieve perfect hallucination (or imagination) control. We showed that four essential properties–truthful knowledge representation, semantic information conservation, complete revelation of relevant knowledge, and optimal response generation–are mutually incompatible. This impossibility emerges from the mathematical structure of information aggregation itself, not from limitations in data, compute, or architecture.
We introduced a rigorous mathematical framework and grounded in it the general analysis of inference processes and their LLM implementations. At the heart of our mathematical framework lies the semantic information measure $\mu_{C}$ and the emergence operator $\mathcal{E}_{C}$. The earlier one is a context-dependent metric capturing how knowledge reduces uncertainty within computational bounds. The later one formalizes how reasoning makes latent knowledge explicit rather than creating it anew. Finally, we model LLM inference as an auction of ideas—a idealized marketplace where neural components compete by bidding with their partial knowledge, each trying to influence the final response.
The impossibility manifests through three complementary proofs, each illuminating different aspects of the fundamental constraint. Through game theory, we apply the Green-Laffont theorem to show the impossibility when knowledge components are independent. Through probability theory, we leverage Savage’s proper scoring rules to extend the result to general trained or fine-tuned probability distributions. And through direct analysis of transformer architectures, we demonstrate how log-sum-exp convexity in attention mechanisms creates measurable violations of information conservation. The last proof exploits Jensen inequality for convex mappings to define the precise mathematical quantity that measures how aggregation of ideas (auction bids) produces overconfident responses that the constituent components cannot justify with revealed knowledge.
These mathematical insights we have made yield some interesting philosophical implications. Hallucination and creativity (two sides of the same coin we call imagination) emerge as mathematically identical phenomena, distinguished only by our normative judgments. The impossibility result explores the outer limits of reason, identified by Gödel’s incompleteness in mathematics, Heisenberg’s uncertainty in physics, and Arrow’s impossibility in social choice. Also, our framework seems to provide mathematical foundations and experimental setting for models of artificial consciousness. That may open interesting avenues for future research.